### v1

# FinTagging: An LLM-ready Benchmark for Extracting and Structuring Financial Information

Abstract
We introduce FinTagging, the first full-scope, table-aware XBRL benchmark designed to evaluate the structured information extraction and semantic alignment capabilities of large language models (LLMs) in the context of XBRL-based financial reporting. Unlike prior benchmarks that oversimplify XBRL tagging as flat multi-class classification and focus solely on narrative text, FinTagging decomposes the XBRL tagging problem into two subtasks: FinNI for financial entity extraction and FinCL for taxonomy-driven concept alignment. It requires models to jointly extract facts and align them with the full 10k+ US-GAAP taxonomy across both unstructured text and structured tables, enabling realistic, fine-grained evaluation. We assess a diverse set of LLMs under zero-shot settings, systematically analyzing their performance on both subtasks and overall tagging accuracy. Our results reveal that, while LLMs demonstrate strong generalization in information extraction, they struggle with fine-grained concept alignment, particularly in disambiguating closely related taxonomy entries. These findings highlight the limitations of existing LLMs in fully automating XBRL tagging and underscore the need for improved semantic reasoning and schema-aware modeling to meet the demands of accurate financial disclosure. Code is available at our GitHub repository [^1] https://github.com/The-FinAI/FinTagging and data is at our Hugging Face repository [^2] https://huggingface.co/collections/TheFinAI/fintagging-68270132372c6608ac069bef.
1 Introduction
Automated tagging is essential in financial reporting, converting structured content such as tables and text into machine-readable data by linking each number to its meaning. Globally, over 2 million companies publish financial reports disclosing earnings, expenses, and liabilities, which provide essential information for investors, regulators, and analysts. However, inconsistent terminology poses challenges for reliable interpretation. To address this, the eXtensible Business Reporting Language (XBRL) was introduced in 1999 as a global standard for machine-readable financial data richards2006introduction . By applying structured tags, XBRL transforms fragmented disclosures into a consistent format, enabling seamless data exchange, cross-entity comparisons, and large-scale analysis. Yet, as shown in Figure 1, accurately tagging over 2,000 numerical facts per report to more than 10,000 standardized concepts remains labor-intensive and error-prone. In 2023 alone, over 6,500 tagging errors were identified across 33,000 SEC filings [^3] https://xbrl.us/wp-content/uploads/2023/03/DQC-SECMeetingNotes-20240314.pdf.
Figure 1: An example of Financial Tagging.
Existing approaches to automated XBRL tagging rely on task-specific models and pretrained language models (PLMs) loukas2022finer ; sharma2023financial , but they require costly annotations, struggle to generalize to new filings, and perform poorly as taxonomies scale. In contrast, recent large language models (LLMs) show strong zero- and few-shot reasoning without task-specific supervision wei2022chain ; cao2025development ; wang2025ordrankbennovelrankingbenchmark ; peng2025plutusbenchmarkinglargelanguage , yet their application to XBRL tagging remains underexplored due to two major limitations (Table 1). First, task framing is oversimplified: prior benchmarks like FiNER loukas2022finer and FNXL sharma2023financial treat tagging as extreme multi-class classification over flat label sets, offering little context for disambiguating fine-grained financial facts. These datasets typically cover only 1k+ concepts, leaving most of the 10k+ US-GAAP taxonomy untested. As shown in Table 8, SOTA LLMs such as DeepSeek-V3 liu2024deepseek and GPT-4o hurst2024gpt achieve 0.0 precision, recall, and F1 under this setting. Second, structured data is ignored: existing datasets exclude tables, despite their central role in financial reporting. Many key facts appear in structured formats, as shown in recent financial QA benchmarks chen2021finqa ; zhu2021tat ; mathew2021docvqa , making this omission a key gap between benchmarks and real-world tagging needs.
Table 1: Detailed comparison of financial NLP benchmarks across task types, sources, and structural capabilities. “QA” means the question-answering task. “Num. Reasoning” indicates the numerical reasoning. “Struct. IE” denotes the structured information extraction.
Benchmark Scenario Data Source Task Modality #Entity Label #Taxonomy Label Num. Reasoning? Struct. IE? Concept Linking?
FinQA chen2021finqa Decision making SEC 10-K QA text/table 0 0
ConvFinQA chen2022convfinqa Decision making SEC 10-K QA text/table 0 0
TAT-QA zhu2021tat Financial analysis Chinese financial reports QA text/table 0 0
DocVQA mathew2021docvqa Enterprise automation Financial document images QA text/image 0 0
FiNER-ORD shah2023finer Financial tagging Chinese financial disclosures Classification text 49 0
FinRED sharma2022finred Knowledge graph construction English financial news Classification text 38 0
FiNER loukas2022finer XBRL tagging SEC 10-K Extreme classification text 0 139
FNXL sharma2023financial XBRL tagging SEC 10-K Extreme classification text 0 2,800
FinTagging (ours) Financial & XBRL tagging SEC 10-K IE + Alignment text/table 5 17,388
To address these issues, we present FinTagging, the first LLM-ready benchmark for full-scope, structure-aware XBRL tagging, which challenges models to perform end-to-end fact extraction and taxonomy alignment on corporate filings. As shown in Table 1, unlike prior work shah2023finer ; loukas2022finer ; sharma2023financial that frames tagging as flat multi-class classification and focuses solely on narrative text, FinTagging requires models to jointly extract financial facts and align them with the US-GAAP taxonomy across both unstructured text and structured tables. This is the first LLM-oriented formulation that scales to the full set of 10k+ taxonomy labels while avoiding the intractability of single-step classification, enabling more fine-grained evaluation of model performance. In collaboration with financial reporting experts, we decompose the task into two components: the first, structured numerical information extraction, assesses a model’s ability to identify key financial facts from both text and tables; the second, fine-grained concept linking, evaluates its ability to map each fact to the correct taxonomy concept among all semantically similar candidates. Unlike prior benchmarks limited to 1,000 frequent concepts, our setting covers the entire taxonomy, exposing the limitations of frequency-biased classification and demanding precise semantic reasoning bovee2005financial . We construct two evaluation sets: FinNI-eval  [^4] https://huggingface.co/datasets/TheFinAI/FinNI-eval for numerical fact extraction and FinCL-eval [^5] https://huggingface.co/datasets/TheFinAI/FinCL-eval for concept linking, both derived from real XBRL submissions and annotated with gold-standard mappings. We also introduce a unified evaluation framework that jointly measures extraction accuracy and concept alignment, providing a rigorous zero-shot assessment of LLMs’ capabilities for practical, table-aware XBRL tagging.
We evaluate ten state-of-the-art LLMs under a zero-shot setting on three fronts: (1) end-to-end macro-F1 over the full FinTagging benchmark, (2) subtask-specific performance on FinNI and FinCL, and (3) ablation of our unified extraction-and-alignment evaluation framework. DeepSeek-V3 liu2024deepseek and GPT-4o hurst2024gpt achieve the highest macro-F1 scores, indicating that our benchmark design enables strong handling of both frequent and rare financial tags, an improvement over traditional PLM baselines evaluated under token-classification settings. In subtask analyses, these LLMs excel at numerical information extraction (FinNI) but continue to struggle with precise concept linking (FinCL), underscoring the remaining challenge of fine-grained semantic alignment. Crucially, our ablation study shows that, without the joint evaluation framework, even top LLMs produce invalid tagging outputs, highlighting the framework’s essential role in realistic assessment of XBRL tagging readiness.
We conclude our main contributions as follows: (1) We introduce FinTagging, the first LLM-ready benchmark for full-scope, table-aware XBRL tagging with a two-stage pipeline—so that modern LLMs can be probed in true zero-shot setting. For each stage we release a new, professionally annotated evaluation set: FinNI-eval and FinCL-eval. (2) We conduct an extensive evaluation of state-of-the-art LLMs in zero-shot settings, systematically analyzing their performance across information extraction and semantic alignment subtasks, as well as overall tagging accuracy. (3) Our experimental results reveal a substantial performance gap between LLMs and the task requirements, particularly in fine-grained semantic alignment. This highlights the limitations of current LLMs in complex financial applications and underscores the need for continued advancements in task-specific adaptation.
2 Related Work
XBRL Tagging Benchmarks
Previous benchmarks have framed XBRL tagging as a large-scale classification problem, primarily focusing on entity recognition. FiNER loukas2022finer introduced a dataset of 1.1 million sentences from SEC filings, annotated with 139 XBRL entity types, addressing challenges like context-sensitive numeric entities and financial domain-specific expressions. Building on this, Sharma et al. sharma2023financial proposed the Financial Numeric Extreme Labelling (FNXL) dataset, expanding the label space to approximately 2,800 XBRL tags and employing extreme classification techniques to handle the increased scale. While these works emphasize the complexity of financial taxonomies, their flat extreme classification approach overlooks the reasoning and alignment needed for realistic XBRL tagging, limiting their suitability for evaluating large language models.
XBRL Tagging Methods
To improve XBRL tagging accuracy, prior work has incorporated structured knowledge and domain-specific modeling. Saini et al. saini2021galaxc proposed GalaXC, a graph neural network leveraging label hierarchies for improved classification. Ma et al. ma2022label enriched transformer models with taxonomy definitions to disambiguate semantically similar tags. Loukas et al. loukas2022finer introduced SECBERT and numeral-aware input transformations to enhance robustness. For extreme multi-label tagging, Sharma et al. sharma2023financial applied AttentionXML to focus on context-relevant segments. Wang wang2023standardizing aligned custom tags with standard taxonomies via semantic similarity using TF-IDF, Word2Vec, and FinBERT. More recently, Han et al. han2024xbrlagent developed XBRL-Agent, an LLM-based system for extracting insights from filings. These efforts highlight the shift toward taxonomy-aware and LLM-enhanced solutions for XBRL tagging.
Financial Evaluation Benchmarks
Broader financial NLP benchmarks address information extraction, reasoning, and document understanding tasks. FiNER-ORD shah2023finer and FinRED sharma2022finred focus on entity recognition and relation extraction, respectively. BizBench koncel2023bizbench assesses the quantitative-reasoning ability of LLMs for both business and finance. Pixiu xie2023pixiu evaluates LLMs across classification, QA, and summarization, while FinQA chen2021finqa and ConvFinQA chen2022convfinqa focus on numerical reasoning over text and tables. TAT-QA zhu2021tat and DocVQA mathew2021docvqa target table-text joint understanding. These benchmarks highlight the importance of integrating structured and unstructured information, yet they overlook the taxonomy-driven fact alignment and do not support the structured output required for XBRL tagging.
3 FinTagging
3.1 Task Formulation
[…]
We formalize the task as a mapping:
$f:(D,L,\mathcal{T})\mapsto\left\{(\texttt{Fact}_{i},\texttt{Type}_{i},\texttt{Tag}_{i})\right\}_{i=1}^{n}$ (1)
where each triplet corresponds to a financial value mention extracted from $D$ and semantically grounded in $\mathcal{T}$.
Inspired by information extraction and alignment works wu2019scalable ; wadden2019entity ; mukherjee2022ectsum ; wang2022conditional and after discussed with financial reporting specialists, we formulated FinTagging into two sub-tasks: Financial Numeric Identification (FinNI), a multi-modal numerical information extraction task, detects numerical value in $D$ and classifies each with its appropriate Type. Financial Concept Linking (FinCL), a numerical entity normalization task, then associates each identified value with its most appropriate Tag in $\mathcal{T}$ based on contextual and structural cues.
Financial Numeric Identification (FinNI).
[…]
We formalize this subtask as a mapping:
$f_{\text{FinNI}}:(D=(S,T),\ L)\mapsto\left\{(e_{i},l_{i})\right\}_{i=1}^{k}$ (2)
where $S$ and $T$ represent the textual and tabular components of the document, respectively, and $L$ is the set of predefined value data types. Each $e_{i}$ is a numerical entity extracted from either $S$ or $T$, and $l_{i}\in L$ denotes its assigned data type.
Financial Concept Linking (FinCL)
[…]
$f_{\text{FinCL}}:(e,l,C_{e},\mathcal{T})\mapsto\hat{c}$ (3)
where $e$ is a numerical entity identified in the document $D=(S,T)$, $l\in L$ is its predicted data type from FinNI, $C_{e}$ denotes the contextual information surrounding $e$ in $D$, and $\mathcal{T}=\{c_{1},c_{2},\dots,c_{n}\}$ is a financial taxonomy containing $n$ uniquely defined and semantically grounded concepts. The model is required to assign a concept $\hat{c}\in\mathcal{T}$ that best reflects the meaning of $e$ in its structural context.
3.2 Raw Data Collection
We collected 30 annual 10-K reports filed in 2024 by publicly listed companies from SEC  [^6] https://www.sec.gov/, as summarized in Table 2. There are a total of 81,325 facts, of which 69,451 are linked to standard taxonomy tags and 11,874 to SEC extension tags. Furthermore, using the BeautifulSoup tool to parse these reports, we identified 76,835 narrative sentences (approximately 16 million characters) and 5,450 financial tables. The companies’ information is shown in Appendix C. These reports follow the XBRL standard. We include all sections that may contain XBRL-tagged content to ensure comprehensive coverage.
Table 2: The statistic of collected financial reports.
Report type Period #Company #Sentence #Char #Table #Std Tag #Custom Tag #Total Tag
Information 10-k 2024-02-13 to 2025-02-13 30 76,835 16,595,283 5,450 69,451 11,874 81,325
3.3 FinTagging Data Annotation
To construct the FinTagging benchmark, we annotated a total of 3,354 sentences and 3,245 table sequences extracted from the narrative and tabular content of collected financial reports, as summarized in Table 3. Among them, 1,627 textual instances and 2,014 table sequences contain XBRL tagging information (positive), while the remaining 1,727 textual and 1,231 table instances do not (negative). On average, textual inputs contain 82.55 tokens, whereas table sequences are significantly longer, averaging 999.74 tokens. Based on this, we further construct two subtask-specific datasets, FinNI-eval and FinCL-eval, with the following steps. The statistics for these two datasets are shown in Table 4.
Table 3: The statistical information for the original dataset in our benchmark. Using the “cl100k_base” tokenizer to calculate tokens ($\pm$ standard deviation).
Structure Pos/Neg #Instance Avg. Tokens/S Avg. Entities/S Avg. Concepts/S Total Entities Unique Concepts
Sentence Positive 1,627 $82.55\pm 66.39$ $1.27\pm 1.84$ $1.27\pm 1.84$ 52,740 2,288
Negative 1,727
Table Positive 2,014 $999.74\pm 993.37$ $14.89\pm 23.50$ $14.89\pm 23.50$
Negative 1,231
For narrative sentences, we retain only textual segments that exceed 20 characters in length and contain numerical terms. For table cell sequences, we use specific labels such as <table>, <tr>, <td>, and <th> to preserve layout and structural information. Tables that contain only headers without any numerical values are considered invalid and thus excluded from our benchmark.
3.3.1 Curation of Numeric Entities and US-GAAP Concepts
Building on the initial filtering, we further refined the sentence and table set to retain only high-quality samples with gold-standard annotations of numerical entity types and their associated US-GAAP concepts. As shown in Table 3, we obtained 52,740 annotated numerical entities spanning 5 entity types, linked to a total of 2,288 unique US-GAAP concepts.
Specifically, we predefine the five of the most prevalent item types, including monetaryItemType, percentItemType, sharesItemType, perShareItemType, and integerItemType. We first excluded sentences and tables lacking US-GAAP labels. Then we conducted a statistical analysis of the entity types contained in the data, as summarized in Figure 2, to identify the most frequent types.
Figure 2: The statistic of numerical entity type.
From Figure 2, we identified 11 numerical entity types across the financial reports of 30 companies. The top five types, based on frequency and coverage, appear in all reports. Sentences or tables containing these types are labeled as positive instances; others are treated as negative.
Table 4: Statistics of evaluation datasets for FinNI-eval and FinCL-eval. Using the “cl100k_base” tokenizer to calculate tokens ($\pm$ standard deviation).
Dataset #Instance Avg Input Tokens Max Input Tokens Avg Output Tokens Max Output Tokens
FinNI-eval 6,599 $986.11\pm 835.52$ 13,205 $138.28\pm 302.92$ 5,465
FinCL-eval 52,572 $61.11\pm 44.60$ 750 $25.00\pm 213.77$ 37
3.3.2 FinNI-eval Dataset Construction
Based on the annotation procedure described in Section 3.3.1, we construct the FinNI-eval dataset, comprising 6,599 samples derived from the annotated sentences and tables. Specifically, the dataset comprises two components: the input block and the answer block. The input block includes a task instruction and the input content. The instruction defines the FinNI task by specifying the responsibilities of the LLMs, the definitions of entity types, and the rules to be followed. The answer block is formatted as a list aligned with the input block. The answer is a JSON list of all identified entity values and types or an empty JSON list. The specific prompt template is shown in the Appendix H.1.
FinNI-eval Dataset
Instruction: <FinNI Task Instruction>
Input: <Sentences or Table>
Answer: {"results": [{"value": <numeric entity>, "type": <entity type>}]}
        or {"results": []}
3.3.3 FinCL-eval Dataset Construction
Following Section 3.1, we further construct a FinCL-eval dataset that includes 52,572 query-answer pairs for numerical entity normalization. In addition, we build a US-GAAP taxonomy database containing 17,388 unique financial concepts. To create the FinCL-eval dataset, we utilize all positive instances identified in Section 3.3.1. Each query consists of a numerical entity, its corresponding entity type, and its surrounding context, while the answer is the associated US-GAAP concept.
FinCL-eval Dataset
Query: <entity> + <entity type> + <context>
Answer: <US-GAAP Concept>
3.4 Evaluation
3.4.1 Evaluation Framework
[…]
Figure 3: A unified evaluation framework on FinTagging benchmark.
FinNI Evaluation
We use pair-level metrics to evaluate the extraction performance on FinNI. The objective is to evaluate the LLM’s ability to distinguish between structured and unstructured contexts and to extract financial values based on its semantic understanding and domain-specific knowledge.
FinCL Evaluation
[…]
$f_{\text{rerank}}:(e,l,C_{e},\mathcal{T}_{h})\mapsto\hat{c}$ (4)
3.4.2 Evaluation Models
Our goal is to evaluate the foundational capabilities of SOTA LLMs on the FinTagging benchmark, to better understand their strengths and limitations in financial tagging. To this end, we evaluate models across three categories using our proposed evaluation framework, as detailed in Appendix E. These include: (1) General-purpose closed-source LLM: GPT-4o model hurst2024gpt ; (2) General-purpose open-source LLMs including DeepSeek-V3 liu2024deepseek , DeepSeek-R1-Distill-Qwen-32B deepseekai2025deepseekr1incentivizingreasoningcapability , Qwen2.5-0.5B-Instruct qwen2 , Qwen2.5-1.5B-Instruct qwen2 , Qwen2.5-14B-Instruct qwen2 , Llama-3.2-3B-Instruct grattafiori2024llama , Llama-3.1-8B-Instruct grattafiori2024llama , and gemma-2-27b-it gemma_2024 models; and (3) Domain-specific financial LLM: Fino1-8B qian2025fino1 model. In addition, we compare these LLMs with strong PLMs, including BERT-large DBLP:journals/corr/abs-1810-04805 , FinBERT araci2019finbert , and SECBERT loukas2022finer .
3.4.3 Evaluation setting
We use the LM Evaluation Harness eval-harness to build customized benchmark suites. Proprietary models such as GPT are accessed via the OpenAI API, while all other models, including DeepSeek and open-source LLMs, are evaluated through the TogetherAI API without local deployment. We standardize the input length to 2,048 tokens for FinNI and 4,096 for FinCL, with a generation limit of 1,024 tokens to support reasoning-intensive outputs. For fine-tuning strong PTMs such as BERT-large, FinBERT, and SECBERT, we curate 10 additional annual reports as training data (see Appendix F). All data follow the FiNER loukas2022finer and FNXL sharma2023financial formats. For retrieval in FinCL, we use ElasticSearch [^8] https://www.elastic.co/ with text-embedding-3-small as the embedding model.
4 Experiment and Result
4.1 Overall Results
Table 5 presents the overall performance on the FinTagging benchmark. It clearly demonstrates that under our framework, LLMs can effectively handle both frequent and rare financial tags, indicating their ability to mitigate long-tail label challenges and underscoring the advantage of our information extraction and alignment formulation over traditional token-level classification approaches.
Table 5: Overall Performance. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
Category Models Macro Micro
P R F1 P R F1
Closed-source LLM GPT-4o 0.0764 0.0576 0.0508 0.0947 0.0788 0.0860
Open-source LLMs DeepSeek-V3 0.0813 0.0696 0.0582 0.1058 0.1217 0.1132
DeepSeek-R1-Distill-Qwen-32B 0.0482 0.0288 0.0266 0.0692 0.0223 0.0337
Qwen2.5-14B-Instruct 0.0423 0.0256 0.0235 0.0197 0.0133 0.0159
gemma-2-27b-it 0.0430 0.0273 0.0254 0.0519 0.0453 0.0483
Llama-3.1-8B-Instruct 0.0287 0.0152 0.0137 0.0462 0.0154 0.0231
Llama-3.2-3B-Instruct 0.0182 0.0109 0.0083 0.0151 0.0102 0.0121
Qwen2.5-1.5B-Instruct 0.0180 0.0079 0.0069 0.0248 0.0060 0.0096
Qwen2.5-0.5B-Instruct 0.0014 0.0003 0.0004 0.0047 0.0001 0.0002
Financial LLM Fino1-8B 0.0299 0.0146 0.0140 0.0355 0.0133 0.0193
Fine-tuned PLMs BERT-large 0.0135 0.0200 0.0126 0.1397 0.1145 0.1259
FinBERT 0.0088 0.0143 0.0087 0.1293 0.0963 0.1104
SECBERT 0.0308 0.0483 0.0331 0.2144 0.2146 0.2145
From the macro-level perspective, which emphasizes balanced performance across both frequent and rare tags, DeepSeek-V3 and GPT-4o achieve the highest macro-F1 scores (0.0582 and 0.0508), outperforming all fine-tuned PLMs. This highlights the strong generalization of large LLMs and the effectiveness of our task design. DeepSeek-R1-Distill-Qwen-32B also achieves a solid macro-F1 (0.0266), suggesting that good architecture and pretraining can help smaller models perform well in zero-shot settings. From the micro-level perspective, which favors frequent labels, DeepSeek-V3 again performs strongly with a micro-F1 of 0.1132, ranking third overall despite no fine-tuning. GPT-4o also performs competitively with a score of 0.0860, outperforming most open-source and domain-specific models.
4.2 Subtask Results
4.2.1 FinNI subtask
Table 6 shows the performance of various LLMs on the FinNI subtask. Overall, larger models perform better at identifying numerical entities and generating structured outputs, even without financial domain training. DeepSeek-V3 achieves the highest precision and recall, outperforming all other models. In contrast, smaller models like Qwen2.5-1.5B and -0.5B score below 0.1 F1, highlighting their limitations in zero-shot settings. Interestingly, DeepSeek-R1-Distill-Qwen-32B shows strong precision, suggesting that model design and distillation can help mitigate the impact of smaller scale.
Table 6: Performance comparison of different models on the FinNI subtask. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
Category Model Precision Recall F1
Closed-source LLM GPT-4o 0.6105 0.5941 0.6022
Open-source LLMs Deepseek-V3 0.6329 0.8452 0.7238
DeepSeek-R1-Distill-Qwen-32B 0.5490 0.2238 0.3180
Qwen2.5-14B-Instruct 0.3632 0.0018 0.0035
gemma-2-27b-it 0.5319 0.5490 0.5403
Llama-3.1-8B-Instruct 0.3346 0.1746 0.2295
Llama-3.2-3B-Instruct 0.1887 0.1794 0.1839
Qwen2.5-1.5B-Instruct 0.1323 0.0636 0.0859
Qwen2.5-0.5B-Instruct 0.0116 0.0027 0.0043
Financial LLM Fino1-8B 0.3416 0.1481 0.2066
However, the domain-specific Fino1-8B model, fine-tuned on the financial reasoning QA task, does not perform competitively. This indicates that only task-specific training can effectively improve performance on FinNI; pretraining on financial corpora alone offers limited benefit if the task is not well aligned.
4.2.2 FinCL subtask
Table 7: Performance comparison of different models on the FinCL subtask. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
Category Model Accuracy
Closed-source LLM GPT-4o 0.1664
Open-source LLMs Deepseek-V3 0.1715
DeepSeek-R1-Distill-Qwen-32B 0.1013
Qwen2.5-14B-Instruct 0.1072
gemma-2-27b-it 0.1009
Llama-3.1-8B-Instruct 0.0807
Llama-3.2-3B-Instruct 0.0375
Qwen2.5-1.5B-Instruct 0.0419
Qwen2.5-0.5B-Instruct 0.0246
Financial LLM Fino1-8B 0.0704
Table 7 shows the accuracy of different models on the FinCL subtask. Overall performance is low, highlighting the challenge of fine-grained concept linking in finance. DeepSeek-V3 achieves the highest accuracy (0.1715), followed by GPT-4o, while all other models score below 0.11. Even large open-source models like Qwen2.5-14B and Gemma-2-27B underperform, and smaller models perform near random. This reflects the difficulty of handling complex taxonomies and subtle financial semantics.
4.3 Ablation analysis
[…]
Table 8: Performance comparison between w/wo our evaluation framework on the FinTagging benchmark dataset.
Evaluation Mode Model Precision Recall F1
FinTagging GPT-4o 0.1075 0.0972 0.1021
Deepseek-V3 0.1229 0.1519 0.1359
Fino1-8B 0.0283 0.0138 0.0185
Extreme Classification GPT-4o 0 0 0
Deepseek-V3 0 0 0
Fino1-8B 0 0 0
Table 8 shows that all LLMs fail completely with extreme classification, yielding zero precision, recall, and F1, confirming that a single-step choice among thousands of flat labels is not a valid LLM evaluation protocol. By decoupling extraction from concept linking and covering the full 10k+ taxonomy, FinTagging produces meaningful scores and thus offers a far more realistic test bed for future model improvements.
5 Conclusion
This paper presents FinTagging, a benchmark for evaluating large language models on XBRL tagging of real-world financial reports. The task is divided into two subtasks, financial numeric identification (FinNI) and concept linking (FinCL), to enable fine-grained evaluation of both information extraction and taxonomy alignment. Results show that while LLMs generalize well to long-tail entities and perform competitively in zero-shot settings, they struggle with accurate alignment to GAAP concepts. This reveals limitations in schema-aware reasoning and highlights the need for better semantic understanding. FinTagging offers a foundation for advancing research in XBRL tagging and regulatory reporting. The limitations of our work and directions for future research are discussed in Appendix A.
Appendix A Limitations
In this work, we propose FinTagging and conduct a comprehensive analysis of LLM performance on the task. However, several limitations remain. First, we have not yet evaluated a broader range of models, including recent releases such as GPT-4.1, LLaMA-4, and Qwen-3. We are currently conducting these experiments and will report the results in future work. Second, although we provide detailed empirical results, we did not perform statistical significance testing, as evaluating the complete benchmark, especially for models like GPT and DeepSeek, requires substantial time. This analysis is also in progress and will be included in future updates. Third, we collected only a small dataset for fine-tuning PLMs and did not construct task-specific training data for LLMs. Future work may explore building LLM-based financial tagging agents through targeted fine-tuning to further enhance performance.
Appendix B Literature Review
The XBRL provides a comprehensive taxonomy for financial reporting, encompassing thousands of detailed tags corresponding to concepts within financial statements. Applying NER to assign XBRL tags is an emerging yet challenging area.
B.1 XBRL Tagging benchmark
FiNER systematically benchmarked several neural architectures on the finer-139 dataset to address numeric-heavy XBRL tagging [14]. The initial experiments showed that standard BERT underperforms due to subword fragmentation, then the authors introduced pseudo-token strategies replacing numerals with [NUM] or [SHAPE] tokens to stabilize label assignment across fragmented numeric spans. These strategies, combined with domain-specific pretraining on SEC-BERT, significantly improved tagging performance, reaching 82.1 micro-F1 without the need for computationally expensive CRF layers. Their experiments demonstrated that subword-aware models with numeric-aware pseudo-tokens outperform both word-level BiLSTMs and vanilla BERT, particularly in numeric-heavy contexts, and avoid nonsensical label sequences. FNXL extended this benchmarking paradigm to a much larger label space of 2,794 US-GAAP tags, reframing the task as an extreme classification problem [23]. They compared the FiNER sequence labeling approach with a two-step pipeline that first identifies numeric spans and then assigns labels using AttentionXML. While FiNER achieved stronger micro-F1 (75.84), reflecting better performance on frequent tags, AttentionXML outperformed FiNER in macro-F1 (47.54), highlighting its strength in predicting infrequent, tail-end labels. FNXL further evaluated both models under a Hits@k setting, confirming that label recommendations from the AttentionXML pipeline could substantially reduce manual effort and maintain high inter-annotator agreement. Together, these benchmarks reveal the need for context-aware reasoning and label-ranking mechanisms in realistic XBRL tagging scenarios.
B.2 XBRL Tagging Methods
The previous studies also explored the approaches address scalability, semantic ambiguity, and reasoning gaps in XBRL tagging to improve the performance. Saini et al. [21] proposed GalaXC is a graph-based extreme classification framework that jointly learns over document-label graphs with per-label attention across multi-hop neighborhoods. By integrating label metadata and transitive label correlations, GalaXC outperformed leading deep classifiers by up to 18% in micro-F1 on standard benchmarks and achieved 25% gains in warm-start scenarios where partial labels are available. Moreover, Wang et al. [28] addressed the practical challenge of custom tag standardization through a semantic similarity pipeline that leverages TF-IDF, Word2Vec, and FinBERT embeddings. Although unsupervised, the method was tested across nearly 200,000 custom tags from SEC filings between 2009 and 2022, and showed strong alignment performance, with vector-based mappings identifying viable standard tag candidates for a substantial proportion of non-compliant elements—offering a low-cost, interpretable solution for downstream financial analysis. Shifting focus from classification to comprehension, XBRL-Agent evaluated the capabilities of large language models to reason over full XBRL reports [10]. The authors introduced two task types—domain taxonomy understanding and numeric reasoning and found that base LLMs often hallucinated or misinterpreted financial content. To overcome these issues, XBRL-Agent incorporated retrieval-augmented generation (RAG) and symbolic calculators within an LLM-agent framework. The enhanced system achieved a 17% accuracy gain on domain query tasks and a 42% boost on numeric reasoning queries compared to base LLMs, validating the utility of modular tool augmentation. These improvements enabled reliable multi-step reasoning over complex disclosures such as debt instruments and derivative gains, which are difficult to capture using span-level classifiers. Collectively, these works broaden the methodological landscape of XBRL tagging from graph-based label propagation and embedding-based normalization to LLM-driven report analysis and point to a hybrid future where structured priors and reasoning tools jointly support accurate, scalable financial information extraction.
B.3 Financial Evaluation Benchmarks
In parallel to XBRL-specific advances, the financial NLP community has developed comprehensive benchmarks to assess broader capabilities in information extraction, numerical reasoning, and document understanding. FiNER-ORD [22] introduced a high-quality, domain-specific NER dataset annotated over financial news, emphasizing general entity types like persons, organizations, and locations. While not numerically focused like FiNER-139, it highlights the lexical diversity of financial discourse and establishes a strong baseline for testing pretrained and zero-shot LLMs in real-world financial NER scenarios. FinQA [4] pushed toward explainable QA by pairing expert-written questions with annotated multi-step reasoning programs derived from earnings reports. ConvFinQA [5] extended this challenge to conversational contexts, simulating real-world question flows over sequential financial queries. TAT-QA [35] focused on hybrid tabular-text reasoning and required models to align cell values and document narratives, often involving aggregation, comparison, and unit-scale interpretation. Pixiu [33] introduced a broader evaluation framework by releasing FinMA, a financial LLM instruction-tuned across five tasks, and assessing it on a new benchmark covering sentiment classification, QA, summarization, NER, and stock prediction. BizBench [12] framed financial QA as program synthesis over realistic, multi-modal contexts, integrating reasoning, code generation, and domain knowledge into a single evaluation pyramid. While these benchmarks highlight the growing ability of models to integrate structured and unstructured financial data, they overlook taxonomy-driven fact alignment and do not support the structured output formats required for XBRL tagging.
Appendix C The Statistics of the Company
Table 9 lists the 30 publicly traded companies whose 2024 annual 10-K filings were used in our study. For each company, we report its Central Index Key (CIK), legal name, stock ticker symbol, and the fiscal year-end filing date. These reports were retrieved from the SEC EDGAR system and conform to the XBRL standard. The selected filings span a diverse set of industries and ensure a representative sample for analyzing structured and unstructured financial disclosures. This company list complements the dataset statistics summarized in Table 2.
Table 9: Company filing summary (2024) for benchmark data
CIK Company Name Ticker Filing Date
0001163739 NABORS INDUSTRIES LTD NBR 2024-12-31
0001418819 Iridium Communications Inc. IRDM 2024-12-31
0001993004 NorthWestern Energy Group, Inc. NWE 2024-12-31
0000024741 CORNING INC /NY GLW 2024-12-31
0000079282 BROWN & BROWN, INC BRO 2024-12-31
0001364479 HERC HOLDINGS INC HRI 2024-12-31
0000936340 DTE ENERGY CO DTM 2024-12-31
0000085535 ROYAL GOLD INC RGLD 2024-12-31
0000086312 TRAVELERS COMPANIES, INC. TRV 2024-12-31
0001968915 PHINIA INC. PHIN 2024-12-31
0000093751 STATE STREET CORP STT 2024-12-31
0000106640 WHIRLPOOL CORP WHR 2024-12-31
0000888491 OMEGA HEALTHCARE INVESTORS INC OHI 2024-12-31
0001616862 Axalta Coating Systems Ltd. AXTA 2024-12-31
0000922224 PPL Corp PPL 2024-12-31
0000066756 ALLETE INC ALE 2024-12-31
0001411207 Allison Transmission Holdings Inc ALSN 2024-12-31
0001478242 IQVIA HOLDINGS INC. IQV 2024-12-31
0001544400 NFiniTi inc. NFTN 2024-12-31
0000851205 COGNEX CORP CGNX 2024-12-31
0001318220 Waste Connections, Inc. WCN 2024-12-31
0000048898 HUBBELL INC HUBB 2024-12-31
0000937098 TRINET GROUP, INC. TNET 2024-12-31
0000910108 LXP Industrial Trust LXP 2024-12-31
0001932393 GE HealthCare Technologies Inc. GEHC 2024-12-31
0001770787 10x Genomics, Inc. TXG 2024-12-31
0001026214 FEDERAL HOME LOAN MORTGAGE CORP FMCC 2024-12-31
0000004904 AMERICAN ELECTRIC POWER CO INC AEP 2024-12-31
0001637459 Kraft Heinz Co KHC 2024-12-31
0001713445 Reddit, Inc. RDDT 2024-12-31
Appendix D Evaluation Metrics
To provide a fair evaluation of overall benchmark performance, we adopt a set of metrics, focusing primarily on macro-level and micro-level evaluation strategies inspired by the previous work [23]. Macro-level evaluation computes precision, recall, and F1 scores independently for each BIO-concept label derived from the US-GAAP taxonomy, and then averages them without weighting. This ensures that each concept, including rare or infrequent ones, contributes equally to the final score—making it especially suitable for domains with skewed label distributions. In contrast, micro-level evaluation aggregates token-level true positives, false positives, and false negatives across all labels before computing precision, recall, and F1. This approach emphasizes the model’s overall tagging accuracy by treating every token equally, and thus better reflects performance on frequent concepts. Together, these two metrics provide a balanced view of both per-concept performance and overall tagging quality.
[…]
$\text{Precision}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{P}|}$ (5)
$\text{Recall}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{G}|}$ (6)
[…]
where $(e_{i},l_{i})=(e^{\prime}_{j},l^{\prime}_{j})$ if and only if both the entity span $e$ and its assigned type $l$ exactly match.
For the FinCL subtask, Given a set of queries $\mathcal{Q}=\{q_{1},q_{2},\dots,q_{N}\}$, where each $q_{i}$ is associated with a ground-truth concept $c_{i}^{*}$ and a predicted concept $\hat{c}_{i}$, the accuracy is defined as:
$\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\delta(\hat{c}_{i},c_{i}^{*})$ (8)
where $\delta(\hat{c}_{i},c_{i}^{*})=1$ if $\hat{c}_{i}=c_{i}^{*}$, and 00 otherwise.
Appendix E Evaluation Models Details
Table 10 provides an overview of the models evaluated in this study, categorized by openness, domain specialization, and architectural foundation. The evaluation covers a diverse range of models:
  • Closed-source LLMs: We include GPT-4o [11], accessed via OpenAI’s API, as a representative of cutting-edge proprietary models with demonstrated performance across a variety of NLP tasks. Although model size details are undisclosed, GPT-4o serves as an upper-bound reference in our benchmark.
  • Open-source LLMs: This group encompasses recent, high-performing open models such as DeepSeek-V3 (685B) [13], DeepSeek-R1-Distill-Qwen (32B) [6], and multiple variants of Qwen2.5-Instruct  [34](ranging from 0.5B to 14B). We also include LLaMA-3.2 and 3.1 variants  [9] (3B and 8B), as well as Google’s Gemma-2-27B  [26], to ensure architectural diversity and scalability comparison. These models are primarily instruction-tuned and optimized for general-purpose NLP tasks.
  • Financial-specific LLMs: We evaluate Fino1-8B  [19], a domain-specialized model trained on financial corpora, designed to better capture the terminology and structure unique to financial disclosures. This category allows us to assess the benefits of domain adaptation in complex tagging and reasoning tasks.
  • Pretrained Language Models (PLMs): To establish strong baselines, we include non-generative encoder models: BERT-large  [7], FinBERT  [1], and SECBERT  [14]. These models have been widely used in prior financial NLP tasks and allow for a comparative analysis between generative LLMs and traditional pretrained models in terms of domain understanding and structured output capability.
Together, these models offer a comprehensive evaluation spectrum, from general-purpose to domain-specific, encoder-based to decoder-based, and open to closed source, facilitating an in-depth assessment of their performance across our proposed benchmark tasks.
Table 10: Model Categories and Corresponding Repositories
Model Size Source
Closed-source Large Language Models
GPT-4o - gpt-4o-2024-08-06
Open-source Large Language Models
DeepSeek-V3 685B deepseek-ai/DeepSeek-V3
DeepSeek-R1-Distill-Qwen 32B deepseek-ai/DeepSeek-R1-Distill-Qwen-32B
Qwen2.5-Instruct 0.5B, 1.5B, 14B Qwen/Qwen2.5-*B-Instruct
Llama-3.2-Instruct 3B meta-llama/Llama-3.2-3B-Instruct
Llama-3.1-Instruct 8B meta-llama/Llama-3.1-8B-Instruct
gemma-2-27b-it 27B google/gemma-2-27b-it
Financial-specific Large Language Models
Fino1 8B TheFinAI/Fino1-8B
Pretrained Language Models
BERT-large $\sim$340M google-bert/bert-large-uncased
FinBERT $\sim$110M ProsusAI/finbert
SECBERT $\sim$110M nlpaueb/sec-bert-base
Appendix F The details for the fine-tuning PTMs
F.1 Training data collection and processing
Similar to the collection process for the FinTagging benchmark data, we gathered an additional 10 annual 10-K financial reports filed with the SEC for the period from February 13, 2024, to February 13, 2025, as summarized in Table 11. These reports contain a total of 33,848 standard taxonomy-tagged facts. Using BeautifulSoup to parse these documents, we identified 22,847 narrative sentences (approximately 5.5 million characters) and 1,236 financial tables. The companies included in this dataset follow the XBRL standard, ensuring comprehensive coverage for training PTMs.
Table 11: Financial report statistics summary for raw training data
Item Information
Report type 10-K
Period 2024-02-13 to 2025-02-13
#Company 10
#Sentence 22,847
#Table 1,236
#Characters 5,539,198
#Standard Tags 33,848
After collection, we employed the same procedure to filter texts and tables, subsequently annotating numerical entities, entity types, and US-GAAP tags (concepts). Finally, as detailed in Table 12, we generated a total of 1,116 sentences and 953 tables as the training set for PTMs. Specifically, the sentence-level data consists of 558 positive and 558 negative instances, averaging approximately 84.24 tokens ($\pm$ 69.29), with 1.22 annotated entities and concepts per sentence. The table-level data comprises 594 positive and 359 negative instances, with a significantly higher average of 1,281.86 tokens ($\pm$ 6,438.37), and approximately 25 entities and concepts annotated per table. Overall, the annotated dataset includes 25,199 entities, covering 1,435 unique US-GAAP concepts.
Table 12: The statistical information for training data. Using the “cl100k_base” tokenizer to calculate tokens ($\pm$ standard deviation).
Structure Pos/Neg #Instance Avg. Tokens/S Avg. Entities/S Avg. Concepts/S Total Entities Unique Concepts
Sentence Positive 558 $84.24\pm 69.29$ $1.22\pm 1.78$ $1.22\pm 1.78$ 25,199 1,435
Negative 558
Table Positive 594 $1281.86\pm 6438.37$ $25.00\pm 213.77$ $25.00\pm 213.77$
Negative 359
However, to align with the extreme classification format used in previous XBRL tagging benchmarks, we directly adopt the US-GAAP tags as entity labels, annotating each token in sentences and tables using the BIO scheme. Specifically, B denotes the beginning of an entity phrase, I marks the continuation (inside) of an entity phrase, and O indicates tokens outside of any entity. As shown in Figure 4, "4.9" and "4.5" are single-token numerical entities labeled only with a B prefix (e.g., “B-us-gaap:AccountsReceivableNetNoncurrent”). To comprehensively cover all US-GAAP tags, we combine the entire set of 17,388 tags from the US-GAAP 2024 taxonomy with the BIO labeling scheme to construct an extreme classification label space, resulting in 34,777 unique entity labels ($2\times 17388+1$).
Figure 4: An example for training set annotation with BIO scheme.
After constructing the training set, we reconstruct the testing set from the original benchmark dataset. The training settings are detailed below.
F.2 Training settings
We fine-tune three pretrained models, BERT-large [7], FinBERT [1], and SECBERT [14], on our training set using the HuggingFace Transformers library. All models are trained with a batch size of 4, a learning rate of 3e-5, and for 20 epochs. Optimization is performed using AdamW without gradient accumulation or early stopping. Token classification heads are randomly initialized and trained jointly with the base encoder. Input sequences are tokenized with a maximum length of 512, and labels are aligned at the sub-token level following the BIO tagging scheme. Loss is computed only on the first sub-token of each word to avoid misalignment bias.
Training is conducted on two NVIDIA A5000 GPUs (24GB each) using data parallelism for 24 hours. All other hyperparameters follow the default settings in the HuggingFace Trainer API. Models are evaluated using the checkpoint from the final epoch. All experiments are run under a fixed random seed to ensure reproducibility.
Appendix G Retrieval Results in FinCL subtask
We investigate the impact of different context construction strategies for the queried entity at the retrieval stage. We consider two approaches: Fixed-Window Context (FWC) and Structure-Aware Context (SAC). In the FWC strategy, context is constructed by extracting a fixed window of 50 characters before and after the entity mention, regardless of whether the entity appears in a sentence or a table. In contrast, the SAC strategy builds context based on the structural location of the entity: if the entity appears in a sentence, the entire sentence is used; if the entity appears in a table, we linearize the entire row into a Markdown-style key-value format to serve as the context.
Table 13: The Acc@k performance of retrieval for different context construction strategies on the FinCL task. Here, FWC denotes the fixed window-based strategy, while SAC denotes the structure-aware approach that respects sentence or row boundaries.
Strategy Structure Acc@1 Acc@10 Acc@20 Acc@50 Acc@100 Acc@150 Acc@200
FWC Sentence 0.0677 0.2594 0.3554 0.4650 0.5515 0.6002 0.6404
Table 0.000 0.0030 0.0031 0.0036 0.0050 0.0064 0.0097
Overall 0.0055 0.0237 0.0316 0.0409 0.0492 0.0544 0.0608
SAC Sentence 0.0689 0.2643 0.3521 0.4711 0.5611 0.6096 0.6522
Table 0.0147 0.0867 0.1240 0.1855 0.2353 0.2632 0.2866
Overall 0.0191 0.1011 0.1424 0.2086 0.2617 0.2913 0.3163
From Table 13, we found the SAC strategy consistently outperforms FWC, particularly in the table context, highlighting the importance of aligning the context window with the underlying structural unit. For sentence-based entities, both strategies yield comparable results, with SAC achieving slightly higher Acc@1 (0.0689 vs. 0.0677) and showing marginal improvements across all cutoff values. This suggests that even for relatively unstructured text, preserving sentence boundaries may provide minor benefits over a fixed-length window.
In contrast, for table-based entities, the performance gap is substantial. SAC achieves significantly higher retrieval accuracy (e.g., Acc@100 of 0.2353 vs. 0.0050), indicating that row-level context is far more informative than arbitrary character windows when dealing with tabular structures. FWC performs poorly in this setting, likely due to the fragmented and semantically sparse nature of partial table text.
When aggregating results across both structures, SAC outperforms FWC by a wide margin at all retrieval depths (e.g., Acc@200 of 0.3163 vs. 0.0608). These results underscore the importance of structure-aware context construction, especially in scenarios where inputs span multiple formats such as sentences and tables.
Appendix H Prompt Templates
H.1 Template for FinNI subtask
Prompt Template You are a financial information extraction expert specializing in identifying financial numerical entities in XBRL reports. Your task is to extract all such numerical entities from the provided text or serialized <table></table> data and classify them into one of five categories: - "integerItemType": Counts of discrete items, such as the number of employees or total transactions. - "monetaryItemType": Financial amounts expressed in currency, such as revenue, profit, or total assets. - "perShareItemType": Per-share values, such as earnings per share (EPS) or book value per share. - "sharesItemType": Counts of shares, such as outstanding shares or ownership stakes. - "percentItemType": Ratios or percentages, such as tax rates, growth rates, or discount rates, usually expressed with a percentage symbol ("%"). Important Instructions: (1) Financial numerical entities are not limited to Arabic numerals (e.g., 10,000). They may also appear in word form (e.g., "ten million"), which must be correctly identified and converted into standard numerical format. (2) Not all numbers in the text should be extracted. Only those that belong to one of the five financial entity categories above should be included. Irrelevant numbers (such as phone numbers, dates, or general IDs) must be ignored. (3) If a number is followed by a magnitude term (e.g., Hundred, Thousand, Million, Billion), do not expand it into the full numerical value. * "Two hundred" -> Extract only "two", not "200". * "10.6 million" -> Extract only "10.6", not "10,600,000". (4) Standardize numerical formatting by removing currency symbols (e.g., "USD"), percentage signs ("%"), and commas (",") while preserving the numeric value. These elements must be removed to ensure consistency. (5) Output the extracted financial entities in JSON list format without explanations, structured as follows: {"result":[{"Fact": <Extracted Numerical Entity>, "Type": <Identified Entity Type>}]} Input: {text/table} Output:
H.2 Template for ablation
Prompt Template You are an XBRL tagging expert specializing in annotating financial numerical facts in XBRL reports. Your task is to (1) extract all such numerical entities from the provided text or serialized <table></table> data, (2) classify them into one of five categories, and (3) assign an appropriated US-GAAP tag to each entity. Categories: - "integerItemType": Counts of discrete items, such as the number of employees or total transactions. - "monetaryItemType": Financial amounts expressed in currency, such as revenue, profit, or total assets. - "perShareItemType": Per-share values, such as earnings per share (EPS) or book value per share. - "sharesItemType": Counts of shares, such as outstanding shares or ownership stakes. - "percentItemType": Ratios or percentages, such as tax rates, growth rates, or discount rates, usually expressed with a percentage symbol ("%"). US-GAAP tags: - A US-GAAP tag is a standardized semantic label used in XBRL filings to identify specific financial concepts defined by the U.S. Generally Accepted Accounting Principles (GAAP). Each tag represents a distinct accounting item and enables consistent, machine-readable financial reporting. - Examples: "us-gaap:AssetsCurrentAbstract", "us-gaap:AccruedInsuranceNoncurrent". Important Instructions: (1) Financial numerical entities are not limited to Arabic numerals (e.g., 10,000). They may also appear in word form (e.g., "ten million"), which must be correctly identified and converted into standard numerical format. (2) Not all numbers in the text should be extracted. Only those that belong to one of the five financial entity categories above should be included. Irrelevant numbers (such as phone numbers, dates, or general IDs) must be ignored. (3) If a number is followed by a magnitude term (e.g., Hundred, Thousand, Million, Billion), do not expand it into the full numerical value. * "Two hundred" -> Extract only "two", not "200". * "10.6 million" -> Extract only "10.6", not "10,600,000". (4) Standardize numerical formatting by removing currency symbols (e.g., "USD"), percentage signs ("%"), and commas (",") while preserving the numeric value. These elements must be removed to ensure consistency. (5) You should assign the most appropriate US-GAAP tag to each identified entity based on your internal understanding of the 2024 US-GAAP taxonomy. (6) Output the extracted financial entities in JSON list format without explanations, structured as follows: {"result":[{"Fact": <Extracted Numerical Entity>, "Type": <Identified Entity Type>, "Tag": <Assigned US-GAAP tag>}]} Input: {text/table} Output:
Appendix I Potential Societal Impacts of the Work
I.1 Potential Positive Societal Impacts
This work introduces FinTagging, a full-scope, structure-aware benchmark for evaluating large language models on XBRL-based financial reporting. By supporting realistic tagging scenarios grounded in US-GAAP taxonomy, this benchmark:
  • Promotes automation and transparency in financial disclosure, which can reduce manual errors and increase the reliability of corporate filings.
  • Enables more accurate, scalable tools for financial regulators, auditors, and analysts, potentially improving oversight and decision-making in global markets.
  • Encourages the development of LLMs with stronger schema-aware reasoning, a critical step toward trustworthy AI systems in high-stakes financial environments.
Through open access to annotated datasets and evaluation code, FinTagging also advances reproducible research and supports the broader community in benchmarking financial language models responsibly.
I.2 Potential Negative Societal Impacts
While FinTagging provides valuable tools for model evaluation, several risks should be noted:
  • Misuse and Misinterpretation: Benchmark performance may be misinterpreted as real-world readiness. Deployment of LLMs without sufficient regulatory safeguards could lead to incorrect tagging, misreporting, or financial misinformation.
  • Coverage Bias: The benchmark focuses on US-GAAP financial reports and may underrepresent other accounting standards or financial contexts, potentially limiting its applicability in non-U.S. jurisdictions.
  • Overreliance on Automation: Improved tagging accuracy may incentivize overautomation in financial reporting pipelines, increasing the risk of undetected errors if human oversight is reduced.
To mitigate these concerns, we recommend that future work includes broader financial contexts and integrates FinTagging into a framework with human-in-the-loop validation, fairness checks, and error auditing. Ongoing collaboration with domain experts and regulators is also essential.

### v2

@@
-FinTagging: An LLM-ready Benchmark for Extracting and Structuring Financial Information
-Abstract
-We introduce FinTagging, the first full-scope, table-aware XBRL benchmark designed to evaluate the structured information extraction and semantic alignment capabilities of large language models (LLMs) in the context of XBRL-based financial reporting.
-Unlike prior benchmarks that oversimplify XBRL tagging as flat multi-class classification and focus solely on narrative text, FinTagging decomposes the XBRL tagging problem into two subtasks: FinNI for financial entity extraction and FinCL for taxonomy-driven concept alignment.
-It requires models to jointly extract facts and align them with the full 10k+ US-GAAP taxonomy across both unstructured text and structured tables, enabling realistic, fine-grained evaluation.
-We assess a diverse set of LLMs under zero-shot settings, systematically analyzing their performance on both subtasks and overall tagging accuracy.
-Our results reveal that, while LLMs demonstrate strong generalization in information extraction, they struggle with fine-grained concept alignment, particularly in disambiguating closely related taxonomy entries.
-These findings highlight the limitations of existing LLMs in fully automating XBRL tagging and underscore the need for improved semantic reasoning and schema-aware modeling to meet the demands of accurate financial disclosure.
-Code is available at our GitHub repository [^1] https://github.com/The-FinAI/FinTagging and data is at our Hugging Face repository [^2] https://huggingface.co/collections/TheFinAI/fintagging-68270132372c6608ac069bef.
-1 Introduction
+
+Abstract.
+Accurately understanding numbers from financial reports is fundamental to how markets, regulators, algorithms, and normal people read the economy and the world, yet even with XBRL (eXtensible Business Reporting Language) designed to tag every figure with standardized accounting concepts, mapping thousands of facts to over 10,000 U.S.
+GAAP concepts remains costly, inconsistent, and error-prone.
+Existing benchmarks define tagging as flat, single-step, extreme classification over small subsets of US-GAAP concepts, overlooking both the taxonomy’s hierarchical semantics and the structured nature of real tagging, where each fact must be represented as a contextualized multi-field output.
+These simplifications prevent fair evaluation of large language models (LLMs) under realistic reporting conditions.
+To address these gaps, we introduce 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first comprehensive benchmark for structure-aware and full-scope XBRL tagging, designed to evaluate LLMs’ ability to extract and align financial facts through numerical reasoning and taxonomy alignment across text and tables.
+We define two subtasks: FinNI for numeric identification, which extracts numerical entities and their types from XBRL reports, and FinCL for concept linking, which maps each extracted entity to the corresponding concept in the full US-GAAP taxonomy.
+Together, these subtasks produce a structured representation of each financial fact.
+We evaluate diverse LLMs under zero-shot settings and analyze their performance across both subtasks and overall tagging accuracy.
+Results show that LLMs generalize well in numeric identification but struggle with fine-grained concept linking, revealing current limitations in structure-aware reasoning for accurate financial disclosure.
+All code and datasets are available on GitHub [^1] https://github.com/The-FinAI/FinTagging/ and Hugging Face [^2] https://huggingface.co/collections/TheFinAI/fintagging-68270132372c6608ac069bef.
+XBRL tagging, Benchmark, Large language model, Information extraction, Information retrieval, Reranking
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Test collectionsccs: Information systems Information extractionccs: Information systems Clustering and classification
+1. Introduction
 Automated tagging is essential in financial reporting, converting structured content such as tables and text into machine-readable data by linking each number to its meaning.
@@
 However, inconsistent terminology poses challenges for reliable interpretation.
-To address this, the eXtensible Business Reporting Language (XBRL) was introduced in 1999 as a global standard for machine-readable financial data richards2006introduction .
-By applying structured tags, XBRL transforms fragmented disclosures into a consistent format, enabling seamless data exchange, cross-entity comparisons, and large-scale analysis.
-Yet, as shown in Figure 1, accurately tagging over 2,000 numerical facts per report to more than 10,000 standardized concepts remains labor-intensive and error-prone.
+To address this, the eXtensible Business Reporting Language (XBRL) was introduced in 1999 as a global standard for machine-readable financial data (Richards et al., 2006).
+By applying structured tags, as shown in Figure 1, XBRL transforms fragmented disclosures into a consistent format, enabling seamless data exchange, cross-entity comparisons, and large-scale analysis.
+However, accurately tagging over 2,000 numerical facts per report to more than 10,000 standardized concepts remains labor-intensive and error-prone.
 In 2023 alone, over 6,500 tagging errors were identified across 33,000 SEC filings [^3] https://xbrl.us/wp-content/uploads/2023/03/DQC-SECMeetingNotes-20240314.pdf.
-Figure 1: An example of Financial Tagging.
-Existing approaches to automated XBRL tagging rely on task-specific models and pretrained language models (PLMs) loukas2022finer ; sharma2023financial , but they require costly annotations, struggle to generalize to new filings, and perform poorly as taxonomies scale.
-In contrast, recent large language models (LLMs) show strong zero- and few-shot reasoning without task-specific supervision wei2022chain ; cao2025development ; wang2025ordrankbennovelrankingbenchmark ; peng2025plutusbenchmarkinglargelanguage , yet their application to XBRL tagging remains underexplored due to two major limitations (Table 1).
-First, task framing is oversimplified: prior benchmarks like FiNER loukas2022finer and FNXL sharma2023financial treat tagging as extreme multi-class classification over flat label sets, offering little context for disambiguating fine-grained financial facts.
+Figure 1.
+An example of Financial Tagging in Realistic, where the orange numbers mark facts to be tagged, black dashed arrows show the structured outputs after tagging, and blue dashed arrows denote the US-GAAP taxonomy referenced during tagging.
+Existing approaches to automated XBRL tagging rely on task-specific models and pretrained language models (PLMs) (Loukas et al., 2022; Sharma et al., 2023), but they require costly annotations, struggle to generalize to new filings, and perform poorly as taxonomies scale.
+In contrast, recent large language models (LLMs) show strong zero- and few-shot reasoning without task-specific supervision (Wei et al., 2022; Cao et al., 2025; Wang et al., 2025; Peng et al., 2025), yet their application to XBRL tagging remains underexplored due to two major limitations (Table 1).
+First, task framing is oversimplified: prior benchmarks like FiNER (Loukas et al., 2022) and FNXL (Sharma et al., 2023) treat tagging as extreme multi-class classification over flat label sets, offering little context for disambiguating fine-grained financial facts.
 These datasets typically cover only 1k+ concepts, leaving most of the 10k+ US-GAAP taxonomy untested.
-As shown in Table 8, SOTA LLMs such as DeepSeek-V3 liu2024deepseek and GPT-4o hurst2024gpt achieve 0.0 precision, recall, and F1 under this setting.
+As shown in Table 8, SOTA LLMs such as DeepSeek-V3 (Liu et al., 2024) and GPT-4o (Hurst et al., 2024) achieve 0.0 precision, recall, and F1 under this setting.
 Second, structured data is ignored: existing datasets exclude tables, despite their central role in financial reporting.
-Many key facts appear in structured formats, as shown in recent financial QA benchmarks chen2021finqa ; zhu2021tat ; mathew2021docvqa , making this omission a key gap between benchmarks and real-world tagging needs.
-Table 1: Detailed comparison of financial NLP benchmarks across task types, sources, and structural capabilities.
-“QA” means the question-answering task.
-“Num.
-Reasoning” indicates the numerical reasoning.
-“Struct.
-IE” denotes the structured information extraction.
-Benchmark Scenario Data Source Task Modality #Entity Label #Taxonomy Label Num. Reasoning? Struct. IE? Concept Linking?
-FinQA chen2021finqa Decision making SEC 10-K QA text/table 0 0
-ConvFinQA chen2022convfinqa Decision making SEC 10-K QA text/table 0 0
-TAT-QA zhu2021tat Financial analysis Chinese financial reports QA text/table 0 0
-DocVQA mathew2021docvqa Enterprise automation Financial document images QA text/image 0 0
-FiNER-ORD shah2023finer Financial tagging Chinese financial disclosures Classification text 49 0
-FinRED sharma2022finred Knowledge graph construction English financial news Classification text 38 0
-FiNER loukas2022finer XBRL tagging SEC 10-K Extreme classification text 0 139
-FNXL sharma2023financial XBRL tagging SEC 10-K Extreme classification text 0 2,800
-FinTagging (ours) Financial & XBRL tagging SEC 10-K IE + Alignment text/table 5 17,388
-To address these issues, we present FinTagging, the first LLM-ready benchmark for full-scope, structure-aware XBRL tagging, which challenges models to perform end-to-end fact extraction and taxonomy alignment on corporate filings.
-As shown in Table 1, unlike prior work shah2023finer ; loukas2022finer ; sharma2023financial that frames tagging as flat multi-class classification and focuses solely on narrative text, FinTagging requires models to jointly extract financial facts and align them with the US-GAAP taxonomy across both unstructured text and structured tables.
+Many key facts appear in structured formats, as shown in recent financial QA benchmarks (Chen et al., 2021; Zhu et al., 2021; Mathew et al., 2021), making this omission a key gap between benchmarks and real-world tagging needs.
+Table 1.
+Detailed comparison of financial NLP benchmarks across task types, sources, and structural capabilities.
+“Num.Reasoning” indicates the numerical reasoning.
+“Struct.IE” denotes the structured information extraction.
+Benchmark Scenario Data Source Task Modality
+#Entity
+Label
+#Taxonomy
+Label
+Num.
+Reasoning?
+Struct.
+IE?
+Concept
+Linking?
+FinQA (Chen et al., 2021) Decision making SEC 10-K QA text/table 0 0
+ConvFinQA (Chen et al., 2022) Decision making SEC 10-K QA text/table 0 0
+TAT-QA (Zhu et al., 2021) Financial analysis Chinese financial reports QA text/table 0 0
+DocVQA (Mathew et al., 2021) Enterprise automation Financial document images QA text/image 0 0
+FiNER-ORD (Shah et al., 2023) Financial tagging Chinese financial disclosures Classification text 49 0
+FinRED (Sharma et al., 2022) Knowledge graph construction English financial news Classification text 38 0
+FiNER (Loukas et al., 2022) XBRL tagging SEC 10-K Extreme classification text 0 139 / 17,388
+FNXL (Sharma et al., 2023) XBRL tagging SEC 10-K Extreme classification text 0 2,800 / 17,388
+FinTagging (ours) Financial & XBRL tagging SEC 10-K IE + Alignment text/table 5 17,388 / 17,388
+To address these issues, we present 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first LLM-ready benchmark for full-scope, structure-aware XBRL tagging, which challenges models to perform end-to-end fact extraction and taxonomy alignment on corporate filings.
+As shown in Table 1, unlike prior work (Shah et al., 2023; Loukas et al., 2022; Sharma et al., 2023) that frames tagging as flat multi-class classification and focuses solely on narrative text, 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: requires models to jointly extract financial facts and align them with the US-GAAP taxonomy across both unstructured text and structured tables.
 This is the first LLM-oriented formulation that scales to the full set of 10k+ taxonomy labels while avoiding the intractability of single-step classification, enabling more fine-grained evaluation of model performance.
 In collaboration with financial reporting experts, we decompose the task into two components: the first, structured numerical information extraction, assesses a model’s ability to identify key financial facts from both text and tables; the second, fine-grained concept linking, evaluates its ability to map each fact to the correct taxonomy concept among all semantically similar candidates.
-Unlike prior benchmarks limited to 1,000 frequent concepts, our setting covers the entire taxonomy, exposing the limitations of frequency-biased classification and demanding precise semantic reasoning bovee2005financial .
-We construct two evaluation sets: FinNI-eval  [^4] https://huggingface.co/datasets/TheFinAI/FinNI-eval for numerical fact extraction and FinCL-eval [^5] https://huggingface.co/datasets/TheFinAI/FinCL-eval for concept linking, both derived from real XBRL submissions and annotated with gold-standard mappings.
+Unlike prior benchmarks limited to 1,000 frequent concepts, our setting covers the entire taxonomy, exposing the limitations of frequency-biased classification and demanding precise semantic reasoning (Bovee et al., 2005).
+We construct two evaluation sets: FinNI-eval [^4] https://huggingface.co/datasets/TheFinAI/FinNI-eval-v1 for numerical fact extraction and FinCL-eval [^5] https://huggingface.co/datasets/TheFinAI/FinCL-eval-v1 for concept linking, both derived from real XBRL submissions and annotated with gold-standard mappings.
 We also introduce a unified evaluation framework that jointly measures extraction accuracy and concept alignment, providing a rigorous zero-shot assessment of LLMs’ capabilities for practical, table-aware XBRL tagging.
-We evaluate ten state-of-the-art LLMs under a zero-shot setting on three fronts: (1) end-to-end macro-F1 over the full FinTagging benchmark, (2) subtask-specific performance on FinNI and FinCL, and (3) ablation of our unified extraction-and-alignment evaluation framework.
-DeepSeek-V3 liu2024deepseek and GPT-4o hurst2024gpt achieve the highest macro-F1 scores, indicating that our benchmark design enables strong handling of both frequent and rare financial tags, an improvement over traditional PLM baselines evaluated under token-classification settings.
+We evaluate 13 state-of-the-art LLMs under a zero-shot setting on three fronts: (1) end-to-end macro-F1 over the full 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark, (2) subtask-specific performance on FinNI and FinCL, and (3) ablation of our unified extraction-and-alignment evaluation framework.
+DeepSeek-V3 (Liu et al., 2024) and GPT-4o (Hurst et al., 2024) achieve the highest macro-F1 scores, indicating that our benchmark design enables strong handling of both frequent and rare financial tags, an improvement over traditional PLM baselines evaluated under token-classification settings.
 In subtask analyses, these LLMs excel at numerical information extraction (FinNI) but continue to struggle with precise concept linking (FinCL), underscoring the remaining challenge of fine-grained semantic alignment.
 Crucially, our ablation study shows that, without the joint evaluation framework, even top LLMs produce invalid tagging outputs, highlighting the framework’s essential role in realistic assessment of XBRL tagging readiness.
-We conclude our main contributions as follows: (1) We introduce FinTagging, the first LLM-ready benchmark for full-scope, table-aware XBRL tagging with a two-stage pipeline—so that modern LLMs can be probed in true zero-shot setting.
-For each stage we release a new, professionally annotated evaluation set: FinNI-eval and FinCL-eval.
+We conclude our main contributions as follows: (1) We introduce 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first LLM-ready benchmark for full-scope, table-aware XBRL tagging with a two-stage pipeline, so that modern LLMs can be probed in a true zero-shot setting.
+For each stage, we release a new, professionally annotated evaluation set: FinNI-eval and FinCL-eval.
 (2) We conduct an extensive evaluation of state-of-the-art LLMs in zero-shot settings, systematically analyzing their performance across information extraction and semantic alignment subtasks, as well as overall tagging accuracy.
@@
 This highlights the limitations of current LLMs in complex financial applications and underscores the need for continued advancements in task-specific adaptation.
-2 Related Work
-XBRL Tagging Benchmarks
-Previous benchmarks have framed XBRL tagging as a large-scale classification problem, primarily focusing on entity recognition.
-FiNER loukas2022finer introduced a dataset of 1.1 million sentences from SEC filings, annotated with 139 XBRL entity types, addressing challenges like context-sensitive numeric entities and financial domain-specific expressions.
+2. Related Work
+XBRL Tagging Benchmarks: Previous benchmarks have framed XBRL tagging as a large-scale classification problem, primarily focusing on entity recognition.
+FiNER (Loukas et al., 2022) introduced a dataset of 1.1 million sentences from SEC filings, annotated with 139 XBRL entity types, addressing challenges like context-sensitive numeric entities and financial domain-specific expressions.
 Building on this, Sharma et al.
-sharma2023financial proposed the Financial Numeric Extreme Labelling (FNXL) dataset, expanding the label space to approximately 2,800 XBRL tags and employing extreme classification techniques to handle the increased scale.
+(Sharma et al., 2023) proposed the Financial Numeric Extreme Labelling (FNXL) dataset, expanding the label space to approximately 2,800 XBRL tags and employing extreme classification techniques to handle the increased scale.
 While these works emphasize the complexity of financial taxonomies, their flat extreme classification approach overlooks the reasoning and alignment needed for realistic XBRL tagging, limiting their suitability for evaluating large language models.
-XBRL Tagging Methods
-To improve XBRL tagging accuracy, prior work has incorporated structured knowledge and domain-specific modeling.
+XBRL Tagging Methods: To improve XBRL tagging accuracy, prior work has incorporated structured knowledge and domain-specific modeling.
 Saini et al.
-saini2021galaxc proposed GalaXC, a graph neural network leveraging label hierarchies for improved classification.
+(Saini et al., 2021) proposed GalaXC, a graph neural network leveraging label hierarchies for improved classification.
 Ma et al.
-ma2022label enriched transformer models with taxonomy definitions to disambiguate semantically similar tags.
+(Ma et al., 2022) enriched transformer models with taxonomy definitions to disambiguate semantically similar tags.
 Loukas et al.
-loukas2022finer introduced SECBERT and numeral-aware input transformations to enhance robustness.
+(Loukas et al., 2022) introduced SECBERT and numeral-aware input transformations to enhance robustness.
 For extreme multi-label tagging, Sharma et al.
-sharma2023financial applied AttentionXML to focus on context-relevant segments.
-Wang wang2023standardizing aligned custom tags with standard taxonomies via semantic similarity using TF-IDF, Word2Vec, and FinBERT.
+(Sharma et al., 2023) applied AttentionXML to focus on context-relevant segments.
+Wang (Wang, 2023) aligned custom tags with standard taxonomies via semantic similarity using TF-IDF, Word2Vec, and FinBERT.
 More recently, Han et al.
-han2024xbrlagent developed XBRL-Agent, an LLM-based system for extracting insights from filings.
+(Han et al., 2024) developed XBRL-Agent, an LLM-based system for extracting insights from filings.
 These efforts highlight the shift toward taxonomy-aware and LLM-enhanced solutions for XBRL tagging.
-Financial Evaluation Benchmarks
-Broader financial NLP benchmarks address information extraction, reasoning, and document understanding tasks.
-FiNER-ORD shah2023finer and FinRED sharma2022finred focus on entity recognition and relation extraction, respectively.
-BizBench koncel2023bizbench assesses the quantitative-reasoning ability of LLMs for both business and finance.
-Pixiu xie2023pixiu evaluates LLMs across classification, QA, and summarization, while FinQA chen2021finqa and ConvFinQA chen2022convfinqa focus on numerical reasoning over text and tables.
-TAT-QA zhu2021tat and DocVQA mathew2021docvqa target table-text joint understanding.
+Financial Evaluation Benchmarks: Broader financial NLP benchmarks address information extraction, reasoning, and document understanding tasks.
+FiNER-ORD (Shah et al., 2023) and FinRED (Sharma et al., 2022) focus on entity recognition and relation extraction, respectively.
+BizBench (Koncel-Kedziorski et al., 2023) assesses the quantitative-reasoning ability of LLMs for both business and finance.
+Pixiu (Xie et al., 2023) evaluates LLMs across classification, QA, and summarization, while FinQA (Chen et al., 2021) and ConvFinQA (Chen et al., 2022) focus on numerical reasoning over text and tables.
+TAT-QA (Zhu et al., 2021) and DocVQA (Mathew et al., 2021) target table-text joint understanding.
 These benchmarks highlight the importance of integrating structured and unstructured information, yet they overlook the taxonomy-driven fact alignment and do not support the structured output required for XBRL tagging.
-3 FinTagging
-3.1 Task Formulation
-[…]
+3.
+0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:
+3.1. Task Formulation
+Formally, given a financial document $D$ consisting of multiple textual contents and structured tables, a set of predefined value data types $L$ specified by the XBRL specification, and a taxonomy database $\mathcal{T}$ containing a set of financial semantic concepts, the XBRL tagging task is to identify all relevant numerical values in $D$ and annotate each with a structured triplet $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$.
+Here, Tag denotes the linked concept in $\mathcal{T}$ (e.g.
+“us-gaap:CashAndDueFrom Banks”, etc.), Fact represents the extracted numerical value as it appears in context (e.g., “2.5”, “10,000”, “two”, etc.), and Type specifies the corresponding data type in $L$ (including monetaryItemType, percentItemType, sharesItemType, perShareItem, and integerItemType).
+Each numerical fact is assigned a single, most specific tag from the taxonomy, without overlapping.
+Specifically, monetaryItemType refers to a financial amount, such as revenue or expenses, typically reported in currency units (e.g., USD).
+percentItemType represents a rate or ratio, expressed as a decimal (e.g., 25%).
+sharesItemType indicates the number of shares, such as stocks held or issued.
+perShareItemType captures values reported on a per-share basis, like earnings per share.
+Lastly, integerItemType is used for whole number counts, such as the number of employees or transactions.
 We formalize the task as a mapping:
-$f:(D,L,\mathcal{T})\mapsto\left\{(\texttt{Fact}_{i},\texttt{Type}_{i},\texttt{Tag}_{i})\right\}_{i=1}^{n}$ (1)
+(1) $f:(D,L,\mathcal{T})\mapsto\left\{(\texttt{Fact}_{i},\texttt{Type}_{i},\texttt{Tag}_{i})\right\}_{i=1}^{n}$
 where each triplet corresponds to a financial value mention extracted from $D$ and semantically grounded in $\mathcal{T}$.
-Inspired by information extraction and alignment works wu2019scalable ; wadden2019entity ; mukherjee2022ectsum ; wang2022conditional and after discussed with financial reporting specialists, we formulated FinTagging into two sub-tasks: Financial Numeric Identification (FinNI), a multi-modal numerical information extraction task, detects numerical value in $D$ and classifies each with its appropriate Type.
+Inspired by information extraction and alignment works (Wu et al., 2019; Wadden et al., 2019; Mukherjee et al., 2022; Wang et al., 2022) and after discussed with financial reporting specialists, we formulated 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: into two sub-tasks: Financial Numeric Identification (FinNI), a multi-modal numerical information extraction task, detects numerical value in $D$ and classifies each with its appropriate Type.
 Financial Concept Linking (FinCL), a numerical entity normalization task, then associates each identified value with its most appropriate Tag in $\mathcal{T}$ based on contextual and structural cues.
-Financial Numeric Identification (FinNI).
-[…]
+Financial Numeric Identification (FinNI)
+The first subtask of 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: focuses on identifying numerical values in a financial document and assigning each a coarse-grained value data type.
+This corresponds to detecting the Fact and Type components of each triplet $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$ defined in the overall task.
 We formalize this subtask as a mapping:
-$f_{\text{FinNI}}:(D=(S,T),\ L)\mapsto\left\{(e_{i},l_{i})\right\}_{i=1}^{k}$ (2)
+(2) $f_{\text{FinNI}}:(D=(S,T),\ L)\mapsto\left\{(e_{i},l_{i})\right\}_{i=1}^{k}$
 where $S$ and $T$ represent the textual and tabular components of the document, respectively, and $L$ is the set of predefined value data types.
@@
 Financial Concept Linking (FinCL)
-[…]
-$f_{\text{FinCL}}:(e,l,C_{e},\mathcal{T})\mapsto\hat{c}$ (3)
+Building on the output of the FinNI subtask, the goal of FinCL is to semantically ground each identified numerical entity $e$ by linking it to a concept $\hat{c}$ in a predefined financial taxonomy, which is the Tag component of each triplet $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$ defined in the overall task.
+Formally, we define the mapping as:
+(3) $f_{\text{FinCL}}:(e,l,C_{e},\mathcal{T})\mapsto\hat{c}$
 where $e$ is a numerical entity identified in the document $D=(S,T)$, $l\in L$ is its predicted data type from FinNI, $C_{e}$ denotes the contextual information surrounding $e$ in $D$, and $\mathcal{T}=\{c_{1},c_{2},\dots,c_{n}\}$ is a financial taxonomy containing $n$ uniquely defined and semantically grounded concepts.
 The model is required to assign a concept $\hat{c}\in\mathcal{T}$ that best reflects the meaning of $e$ in its structural context.
-3.2 Raw Data Collection
-We collected 30 annual 10-K reports filed in 2024 by publicly listed companies from SEC  [^6] https://www.sec.gov/, as summarized in Table 2.
-There are a total of 81,325 facts, of which 69,451 are linked to standard taxonomy tags and 11,874 to SEC extension tags.
-Furthermore, using the BeautifulSoup tool to parse these reports, we identified 76,835 narrative sentences (approximately 16 million characters) and 5,450 financial tables.
-The companies’ information is shown in Appendix C.
-These reports follow the XBRL standard.
-We include all sections that may contain XBRL-tagged content to ensure comprehensive coverage.
-Table 2: The statistic of collected financial reports.
-Report type Period #Company #Sentence #Char #Table #Std Tag #Custom Tag #Total Tag
-Information 10-k 2024-02-13 to 2025-02-13 30 76,835 16,595,283 5,450 69,451 11,874 81,325
-3.3 FinTagging Data Annotation
-To construct the FinTagging benchmark, we annotated a total of 3,354 sentences and 3,245 table sequences extracted from the narrative and tabular content of collected financial reports, as summarized in Table 3.
-Among them, 1,627 textual instances and 2,014 table sequences contain XBRL tagging information (positive), while the remaining 1,727 textual and 1,231 table instances do not (negative).
-On average, textual inputs contain 82.55 tokens, whereas table sequences are significantly longer, averaging 999.74 tokens.
-Based on this, we further construct two subtask-specific datasets, FinNI-eval and FinCL-eval, with the following steps.
-The statistics for these two datasets are shown in Table 4.
-Table 3: The statistical information for the original dataset in our benchmark. Using the “cl100k_base” tokenizer to calculate tokens ($\pm$ standard deviation).
-Structure Pos/Neg #Instance Avg. Tokens/S Avg. Entities/S Avg. Concepts/S Total Entities Unique Concepts
-Sentence Positive 1,627 $82.55\pm 66.39$ $1.27\pm 1.84$ $1.27\pm 1.84$ 52,740 2,288
-Negative 1,727
-Table Positive 2,014 $999.74\pm 993.37$ $14.89\pm 23.50$ $14.89\pm 23.50$
-Negative 1,231
-For narrative sentences, we retain only textual segments that exceed 20 characters in length and contain numerical terms.
-For table cell sequences, we use specific labels such as <table>, <tr>, <td>, and <th> to preserve layout and structural information.
-Tables that contain only headers without any numerical values are considered invalid and thus excluded from our benchmark.
-3.3.1 Curation of Numeric Entities and US-GAAP Concepts
-Building on the initial filtering, we further refined the sentence and table set to retain only high-quality samples with gold-standard annotations of numerical entity types and their associated US-GAAP concepts.
-As shown in Table 3, we obtained 52,740 annotated numerical entities spanning 5 entity types, linked to a total of 2,288 unique US-GAAP concepts.
-Specifically, we predefine the five of the most prevalent item types, including monetaryItemType, percentItemType, sharesItemType, perShareItemType, and integerItemType.
-We first excluded sentences and tables lacking US-GAAP labels.
-Then we conducted a statistical analysis of the entity types contained in the data, as summarized in Figure 2, to identify the most frequent types.
-Figure 2: The statistic of numerical entity type.
-From Figure 2, we identified 11 numerical entity types across the financial reports of 30 companies.
-The top five types, based on frequency and coverage, appear in all reports.
-Sentences or tables containing these types are labeled as positive instances; others are treated as negative.
-Table 4: Statistics of evaluation datasets for FinNI-eval and FinCL-eval. Using the “cl100k_base” tokenizer to calculate tokens ($\pm$ standard deviation).
-Dataset #Instance Avg Input Tokens Max Input Tokens Avg Output Tokens Max Output Tokens
-FinNI-eval 6,599 $986.11\pm 835.52$ 13,205 $138.28\pm 302.92$ 5,465
-FinCL-eval 52,572 $61.11\pm 44.60$ 750 $25.00\pm 213.77$ 37
-3.3.2 FinNI-eval Dataset Construction
-Based on the annotation procedure described in Section 3.3.1, we construct the FinNI-eval dataset, comprising 6,599 samples derived from the annotated sentences and tables.
+3.2. Raw Data Collection
+We compiled 142 annual 10-K reports filed in 2023 and 2024 by publicly listed companies from the SEC [^6] https://www.sec.gov/, as shown in Table 2.
+To ensure diversity, the collection covers all 11 major industry sectors, with about 95% of companies distributed across 31 U.S.
+states and the remaining 5% based outside the United States.
+Using BeautifulSoup, we parsed these reports and extracted 319,893 narrative sentences (roughly 76 million characters) along with 21,576 financial tables.
+Company-level details are provided in Appendix D.2.
+Each report follows the SEC’s XBRL filing standard, meaning that the underlying instance, schema, and linkbase documents already contain machine-readable tags (ground-truth tags) that link textual and tabular values to US-GAAP taxonomy concepts.
+We included all sections that may reference or embed XBRL-tagged content to ensure comprehensive coverage.
+For the justification of our benchmark granularity, please see the Appendix E.
+Table 2. Statistics of collected financial reports.
+Item Value
+Report type 10-K
+Period 2023-02-13 to 2025-02-13
+#Company 142
+#Covered sector 11
+#Covered jurisdiction 31 States + outside US
+#Sentence 319,893
+#Char 75,748,949
+#Table 21,576
+3.3. Data Curation
+To construct the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark and ensure both efficiency and fairness in evaluation, we designed a two-stage filtering and annotation pipeline.
+This pipeline systematically processes the raw 10-K filings to retain representative, high-quality instances that collectively cover all 142 companies, five valid entity types, and every numerical US-GAAP concept observed in the corpus.
+The overall workflow is illustrated in Algorithms 1 and 2.
+Stage 1: Filtering Raw Instances (Algorithm 1)
+The first stage aims to remove noisy or redundant samples while ensuring comprehensive coverage of both entity types and US-GAAP concepts.
+Each narrative sentence or table instance is automatically examined by a rule-based parser that checks if it contains numeric entities tagged with valid US-GAAP concepts.
+The parser performs schema-level consistency checks based on the taxonomy definitions to ensure tagging correctness.
+Instances without valid entities are classified as NEG (negative).
+Those containing valid entities are treated as positive and further divided into two categories: (1) POS_SELECT, which introduces at least one new US-GAAP concept not yet covered; and (2) POS_CANDIDATE, which contains only previously covered concepts.
+This design ensures that each concept appears at least once in the dataset while minimizing redundancy across companies and reports.
+Figure 2. Statistics of numerical entity types.
+To determine which entity types should be considered valid, we first profiled all XBRL numeric item types that appeared in the raw corpus and summarized the results in Figure 2.
+For each type, we calculated both its cross-company coverage (the number of companies in which it occurs) and its overall frequency (the proportion of all annotated numeric entities belonging to that type).
+The figure shows a clear long-tailed distribution: a total of eleven numeric item types were identified, but only a few dominate in both coverage and frequency.
+The five types that consistently appear across nearly all companies and collectively account for the majority of entities are monetaryItemType, percentItemType, sharesItemType, perShareItemType, and integerItemType.
+These five types form the label space for all subsequent stages.
+Under this definition, Algorithm 1 treats an instance as a potential positive only if it contains at least one US-GAAP–tagged entity belonging to this label space; otherwise it is assigned to NEG.
+To maintain structural fidelity, additional validity rules are applied.
+For narrative texts, we retain only segments longer than 20 characters that contain numerical tokens.
+For tables, we preserve the <table>, <tr>, <td>, and <th> tags to ensure layout integrity, while tables with headers only (i.e., without numeric values) are excluded.
+Input: 10-K reports; US-GAAP taxonomy; 5 valid entity types
+Output: POS_SELECT, POS_CANDIDATE, NEG
+Initialize: an empty set of covered concepts; empty lists for POS_SELECT, POS_CANDIDATE, and NEG;
+foreach instance (texts or tables) in 10-K reports do
+  Check if the instance contains numeric entities that are both: (i) tagged with a US-GAAP concept, and (ii) of a valid entity type;
+ if no such entities are found then
+     Add the instance to NEG;
+ else
+     Identify all US-GAAP concepts in this instance;
+     if the instance introduces at least one new concept not yet covered then
+        Add the instance to POS_SELECT and update the covered concepts set;
+    else
+        Add the instance to POS_CANDIDATE;
+return {POS_SELECT, POS_CANDIDATE, NEG};
+Algorithm 1 Screening of Text and Table Instances
+Stage 2: Company-Capped Sampling and Minimal Cover (Algorithm 2)
+The second stage refines the positive pool by enforcing both per-company balance and concept coverage completeness.
+For each modality (text or table), we first perform company-capped seeding: from each company’s POS_SELECT, up to $K=10$ positive instances are retained as the seed set, while the remaining positives are moved into a candidate pool.
+We then mark all US-GAAP concepts covered by the seed and identify those still uncovered.
+Finally, a greedy minimal-cover procedure is executed: candidates are iteratively selected according to the number of uncovered concepts they can cover, with ties broken by shorter text length, fewer entities, or earlier order.
+Each selected candidate is added to the final set $P^{*}$, and the process continues until all concepts are covered or the candidate pool is exhausted.
+The resulting positive set $P^{*}$ thus maintains full US-GAAP concept coverage, balanced company representation, and reduced redundancy, forming a compact yet coverage-complete dataset.
+Input: POS_SELECT grouped by company (run separately for text and table) from algorithm 1;
+   All concepts observed in the POS_SELECT;
+   Cap $K=10$ (maximum positives per company for seeding)
+Output: Final positive instances $P^{*}$ and leftover instances
+Seed selection. For each company, keep up to $K$ positives as the seed set. Put the rest into a new candidate pool.
+Check coverage. Mark all concepts already covered by the seed. Identify which concepts are still uncovered.
+Greedy covering. While uncovered concepts remain:
+  1. (1)
+    For each candidate, count how many uncovered concepts it can cover.
+  2. (2)
+    Pick the candidate with the highest count (ties broken by shorter text, fewer entities, or earlier order).
+  3. (3)
+    Add the candidate to the final positives, update the covered concepts, and remove it from the pool.
+Finalize. The final positives $P^{*}$ are the union of the seed and the greedy additions. The leftover candidates form the candidate set.
+Algorithm 2 Company-Capped Seeding + Minimal Cover (for Text and Table)
+Final Dataset and Validation
+After two-stage filtering, we obtained 3,084 textual sentences and 3,067 table sequences from financial filings (Table 3).
+Among them, 2,054 text and 2,045 table instances are positives containing valid XBRL tagging information, while 1,030 text and 1,022 table instances are negatives.
+Textual inputs contain on average $96.25$ tokens, whereas table sequences are much longer ($1,204.72$ tokens on average).
+Overall, the dataset includes 62,668 annotated entities linked to 3,953 unique US-GAAP concepts across the five valid entity types.
+Based on this foundation, we further derive two subtask-specific datasets: FinNI-eval, targeting numerical entity identification, and FinCL-eval, targeting concept linking.
+Both are the first datasets in the financial domain to emphasize structured information extraction and concept-level alignment.
+Their detailed statistics are reported in Table 4.
+To verify annotation reliability, two domain experts independently reviewed 100 randomly sampled entries.
+The results indicate a 96% raw agreement and a Cohen’s Kappa of 81.13, confirming the robustness and reproducibility of our semi-automated annotation pipeline.
+The validation guideline is provided in Appendix F, and a complete unfiltered version of the benchmark is included in Appendix D.1.
+Table 3. Statistical information for the original dataset in our benchmark. Tokens are calculated using the “cl100k_base” tokenizer ($\pm$ standard deviation).
+Item Sentence Table
+Positive instances 2,054 2,045
+Negative instances 1,030 1,022
+Avg. Tokens/S $96.25\pm 72.75$ $1,204.72\pm 1,162.41$
+Avg. Entities/S $1.88\pm 2.11$ $18.54\pm 29.14$
+Avg. Concepts/S $1.88\pm 2.11$ $18.54\pm 29.14$
+Total Entities 62,668
+Entity Types 5
+Unique Concepts 3,953
+Table 4. Statistics of evaluation datasets for FinNI-eval and FinCL-eval. Tokens are calculated using the “cl100k_base” tokenizer ($\pm$ standard deviation).
+Item FinNI-eval FinCL-eval
+#Instance 6,151 62,668
+Avg. Input Tokens $1101.50\pm 992.00$ $60.43\pm 48.25$
+Max Input Tokens 18,831 750
+Avg. Output Tokens $175.24\pm 375.90$ $12.82\pm 5.09$
+Max Output Tokens 7,813 41
+3.4. FinNI-eval Dataset Construction
+Based on the annotation procedure described in Section 3.3, we construct the FinNI-eval dataset, comprising 6,151 instances derived from the annotated sentences and tables.
 Specifically, the dataset comprises two components: the input block and the answer block.
@@
 The answer is a JSON list of all identified entity values and types or an empty JSON list.
-The specific prompt template is shown in the Appendix H.1.
+The specific prompt template is shown in Figure 9 in the Appendix.
 FinNI-eval Dataset
@@
         or {"results": []}
-3.3.3 FinCL-eval Dataset Construction
-Following Section 3.1, we further construct a FinCL-eval dataset that includes 52,572 query-answer pairs for numerical entity normalization.
-In addition, we build a US-GAAP taxonomy database containing 17,388 unique financial concepts.
-To create the FinCL-eval dataset, we utilize all positive instances identified in Section 3.3.1.
+3.5. FinCL-eval Dataset Construction
+Following the formulation of the FinCL subtask in Section 3.1, we further construct a FinCL-eval dataset that includes 62,668 query-answer pairs for numerical entity normalization.
+In addition, we built a US-GAAP taxonomy database containing 17,688 unique financial concepts.
+To create the FinCL-eval dataset, we utilize all positive instances identified in Section 3.3.
 Each query consists of a numerical entity, its corresponding entity type, and its surrounding context, while the answer is the associated US-GAAP concept.
@@
 Answer: <US-GAAP Concept>
-3.4 Evaluation
-3.4.1 Evaluation Framework
-[…]
-Figure 3: A unified evaluation framework on FinTagging benchmark.
-FinNI Evaluation
-We use pair-level metrics to evaluate the extraction performance on FinNI.
+3.6. Evaluation
+3.6.1. Evaluation Framework
+We propose a unified evaluation framework for the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark to assess LLM performance.
+As shown in Figure 3, given a financial report $D$, the framework reformulates tagging into two subtasks, FinNI and FinCL, to generate structured triplets $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$.
+This design jointly evaluates an LLM’s zero-shot ability in fact extraction and concept alignment.
+For fair comparison with token-classification baselines, we report macro and micro Precision, Recall, and F1 (Sokolova and Lapalme, 2009).
+FinNI is evaluated using pair-level metrics, while FinCL uses accuracy.
+Full metric definitions are provided in Appendix G.
+Figure 3.
+A unified evaluation framework on 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark.
+Note: The Sentence is drawn from financial report $D$, and the Query provides an example that incorporates the entity 4.9.
+The black arrows show the overall framework flow, green arrows denote the FinNI process, red arrows indicate the FinCL process, red bidirectional arrows represent retrieval interactions with the taxonomy, and red dashed arrows mark the recomposed input between the query and retrieved candidates.
+FinNI Evaluation: We use pair-level metrics to evaluate the extraction performance on FinNI.
 The objective is to evaluate the LLM’s ability to distinguish between structured and unstructured contexts and to extract financial values based on its semantic understanding and domain-specific knowledge.
-FinCL Evaluation
-[…]
-$f_{\text{rerank}}:(e,l,C_{e},\mathcal{T}_{h})\mapsto\hat{c}$ (4)
-3.4.2 Evaluation Models
-Our goal is to evaluate the foundational capabilities of SOTA LLMs on the FinTagging benchmark, to better understand their strengths and limitations in financial tagging.
-To this end, we evaluate models across three categories using our proposed evaluation framework, as detailed in Appendix E.
-These include: (1) General-purpose closed-source LLM: GPT-4o model hurst2024gpt ; (2) General-purpose open-source LLMs including DeepSeek-V3 liu2024deepseek , DeepSeek-R1-Distill-Qwen-32B deepseekai2025deepseekr1incentivizingreasoningcapability , Qwen2.5-0.5B-Instruct qwen2 , Qwen2.5-1.5B-Instruct qwen2 , Qwen2.5-14B-Instruct qwen2 , Llama-3.2-3B-Instruct grattafiori2024llama , Llama-3.1-8B-Instruct grattafiori2024llama , and gemma-2-27b-it gemma_2024 models; and (3) Domain-specific financial LLM: Fino1-8B qian2025fino1 model.
-In addition, we compare these LLMs with strong PLMs, including BERT-large DBLP:journals/corr/abs-1810-04805 , FinBERT araci2019finbert , and SECBERT loukas2022finer .
-3.4.3 Evaluation setting
-We use the LM Evaluation Harness eval-harness to build customized benchmark suites.
-Proprietary models such as GPT are accessed via the OpenAI API, while all other models, including DeepSeek and open-source LLMs, are evaluated through the TogetherAI API without local deployment.
+FinCL Evaluation: To assess whether LLMs can perform fine-grained XBRL entity normalization, we reformulate the task as a retrieval-reranking problem, where LLMs are used to disambiguate and select the most appropriate taxonomy concept (reranking stage) from a reduced candidate set (retrieval stage), avoiding the impracticality of direct multi-thousand-way classification (prompt template used for reranking in the FinCL subtask is shown in Figure 10 in Appendix).
+To this end, we first generate embeddings for each taxonomy concept using text-embedding-3-small [^7] https://platform.openai.com/docs/models/text-embedding-3-small, and retrieve the top-$h$ candidate tags from the US-GAAP taxonomy $\mathcal{T}$ based on semantic similarity.
+Therefore, the LLMs are required to rerank the retrieved candidates by leveraging deeper contextual understanding to select the best-matched tag $\hat{c}\in\mathcal{T}_{h}$, formalized as:
+(4) $f_{\text{rerank}}:(e,l,C_{e},\mathcal{T}_{h})\mapsto\hat{c}$
+3.6.2. Evaluation Models
+Our goal is to evaluate the foundational capabilities of SOTA LLMs on the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark, to better understand their strengths and limitations in financial tagging.
+To this end, we evaluate models across three categories using our proposed evaluation framework, as detailed in Table 13.
+These include: (1) one general-purpose closed-source LLM: GPT-4o model (Hurst et al., 2024); (2) eleven general-purpose open-source LLMs including DeepSeek-V3 (Liu et al., 2024), DeepSeek-R1-Distill-Qwen-32B (DeepSeek-AI, 2025), Qwen3-32B (Team, 2025), Qwen3-14B (Team, 2025), Qwen3-1.7B (Team, 2025), Qwen3-0.6B (Team, 2025), Llama-4-Scout-17B-16E-Instruct (Meta, 2025), Llama-3.3-70B-Instruct (Grattafiori et al., 2024), Llama-3.1-8B-Instruct (Grattafiori et al., 2024), Llama-3.2-3B-Instruct (Grattafiori et al., 2024), and gemma-2-27b-it (Team, 2024) models; and (3) one domain-specific financial LLM: Fino1-8B (Qian et al., 2025) model.
+In addition, we compare these LLMs with strong PLMs, including BERT-large (Devlin et al., 2018), FinBERT (Araci, 2019), and SECBERT (Loukas et al., 2022).
+3.6.3. Evaluation setting
+We use the LM Evaluation Harness (Gao et al., 2024) to build customized benchmark suites.
+Proprietary models are accessed via APIs, while open-source models are evaluated locally on a 4×H100 GPU cluster (80GB each).
 We standardize the input length to 2,048 tokens for FinNI and 4,096 for FinCL, with a generation limit of 1,024 tokens to support reasoning-intensive outputs.
-For fine-tuning strong PTMs such as BERT-large, FinBERT, and SECBERT, we curate 10 additional annual reports as training data (see Appendix F).
-All data follow the FiNER loukas2022finer and FNXL sharma2023financial formats.
+For fine-tuning strong PTMs such as BERT-large, FinBERT, and SECBERT, we curate 10 additional annual reports as training data (see Appendix I).
+All data follow the FiNER (Loukas et al., 2022) and FNXL (Sharma et al., 2023) formats.
 For retrieval in FinCL, we use ElasticSearch [^8] https://www.elastic.co/ with text-embedding-3-small as the embedding model.
-4 Experiment and Result
-4.1 Overall Results
-Table 5 presents the overall performance on the FinTagging benchmark.
+For evaluation, we spent a total of 500 GPU hours and about \$1,500 for GPT (from OpenAI) and DeepSeek (from TogetherAI) API.
+4. Experiment and Result
+4.1. Overall Results
+Table 5 presents the overall performance on the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark.
 It clearly demonstrates that under our framework, LLMs can effectively handle both frequent and rare financial tags, indicating their ability to mitigate long-tail label challenges and underscoring the advantage of our information extraction and alignment formulation over traditional token-level classification approaches.
-Table 5: Overall Performance. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
+Table 5. Overall Performance. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
 Category Models Macro Micro
 P R F1 P R F1
-Closed-source LLM GPT-4o 0.0764 0.0576 0.0508 0.0947 0.0788 0.0860
-Open-source LLMs DeepSeek-V3 0.0813 0.0696 0.0582 0.1058 0.1217 0.1132
-DeepSeek-R1-Distill-Qwen-32B 0.0482 0.0288 0.0266 0.0692 0.0223 0.0337
-Qwen2.5-14B-Instruct 0.0423 0.0256 0.0235 0.0197 0.0133 0.0159
-gemma-2-27b-it 0.0430 0.0273 0.0254 0.0519 0.0453 0.0483
-Llama-3.1-8B-Instruct 0.0287 0.0152 0.0137 0.0462 0.0154 0.0231
-Llama-3.2-3B-Instruct 0.0182 0.0109 0.0083 0.0151 0.0102 0.0121
-Qwen2.5-1.5B-Instruct 0.0180 0.0079 0.0069 0.0248 0.0060 0.0096
-Qwen2.5-0.5B-Instruct 0.0014 0.0003 0.0004 0.0047 0.0001 0.0002
-Financial LLM Fino1-8B 0.0299 0.0146 0.0140 0.0355 0.0133 0.0193
-Fine-tuned PLMs BERT-large 0.0135 0.0200 0.0126 0.1397 0.1145 0.1259
-FinBERT 0.0088 0.0143 0.0087 0.1293 0.0963 0.1104
-SECBERT 0.0308 0.0483 0.0331 0.2144 0.2146 0.2145
-From the macro-level perspective, which emphasizes balanced performance across both frequent and rare tags, DeepSeek-V3 and GPT-4o achieve the highest macro-F1 scores (0.0582 and 0.0508), outperforming all fine-tuned PLMs.
+Closed-source LLM GPT-4o 0.0865 0.0626 0.0569 0.1052 0.0731 0.0863
+Open-source LLMs DeepSeek-V3 0.0949 0.0778 0.0659 0.1074 0.1264 0.1162
+Llama-4-Scout-17B-16E-Instruct 0.0683 0.0414 0.0400 0.1045 0.0526 0.0699
+Llama-3.3-70B-Instruct 0.0544 0.0279 0.0288 0.0665 0.0382 0.0485
+DeepSeek-R1-Distill-Qwen-32B 0.0532 0.0283 0.0285 0.0814 0.0214 0.0339
+Qwen3-32B 0.0639 0.0314 0.0324 0.1127 0.0230 0.0382
+gemma-2-27b-it 0.0471 0.0291 0.0276 0.0533 0.0390 0.0451
+Qwen3-14B 0.0591 0.0274 0.0288 0.1069 0.0182 0.0311
+Llama-3.1-8B-Instruct 0.0345 0.0178 0.0169 0.0575 0.0166 0.0258
+Llama-3.2-3B-Instruct 0.0194 0.0118 0.0100 0.0178 0.0084 0.0114
+Qwen3-1.7B 0.0207 0.0067 0.0080 0.1063 0.0031 0.0060
+Qwen3-0.6B 0.0026 0.0008 0.0010 0.0562 0.0002 0.0005
+Financial LLM Fino1-8B 0.0344 0.0143 0.0151 0.0419 0.0128 0.0197
+Fine-tuned PLMs BERT-large 0.0252 0.0266 0.0205 0.1518 0.1283 0.1391
+FinBERT 0.0046 0.0064 0.0042 0.0872 0.0526 0.0656
+SECBERT 0.0231 0.0295 0.0203 0.1870 0.1697 0.1779
+From the macro-level perspective, which emphasizes balanced performance across both frequent and rare tags, DeepSeek-V3, GPT-4o, and Llama-4-Scout-17B-16E-Instruct achieve the top 3 macro-F1 scores (0.0659, 0.0569, and 0.0400), outperforming all fine-tuned PLMs.
 This highlights the strong generalization of large LLMs and the effectiveness of our task design.
-DeepSeek-R1-Distill-Qwen-32B also achieves a solid macro-F1 (0.0266), suggesting that good architecture and pretraining can help smaller models perform well in zero-shot settings.
-From the micro-level perspective, which favors frequent labels, DeepSeek-V3 again performs strongly with a micro-F1 of 0.1132, ranking third overall despite no fine-tuning.
-GPT-4o also performs competitively with a score of 0.0860, outperforming most open-source and domain-specific models.
-4.2 Subtask Results
-4.2.1 FinNI subtask
-Table 6 shows the performance of various LLMs on the FinNI subtask.
-Overall, larger models perform better at identifying numerical entities and generating structured outputs, even without financial domain training.
-DeepSeek-V3 achieves the highest precision and recall, outperforming all other models.
-In contrast, smaller models like Qwen2.5-1.5B and -0.5B score below 0.1 F1, highlighting their limitations in zero-shot settings.
-Interestingly, DeepSeek-R1-Distill-Qwen-32B shows strong precision, suggesting that model design and distillation can help mitigate the impact of smaller scale.
-Table 6: Performance comparison of different models on the FinNI subtask.
+Qwen3-32B also achieves a solid macro-F1 (0.0324), suggesting that good architecture and pretraining can help smaller models perform well in zero-shot settings.
+From the micro-level perspective, which favors frequent labels, DeepSeek-V3 again performs strongly with a micro-F1 of 0.1162, ranking third overall despite no fine-tuning.
+GPT-4o also performs competitively with a score of 0.0863, outperforming most open-source and domain-specific models.
+4.2. Subtask Results
+4.2.1. FinNI subtask
+Table 6 presents the results of different models on the FinNI subtask.
+Overall, larger models demonstrate stronger capabilities in identifying numerical entities and producing structured outputs, even without financial domain adaptation.
+Among them, DeepSeek-V3 achieves the best performance with both the highest recall (0.8430) and F1 (0.6932), while also maintaining competitive precision.
+In comparison, GPT-4o, the closed-source baseline, performs relatively well with balanced scores, securing the second-best F1 (0.5397).
+Interestingly, Qwen3-32B and Qwen3-14B yield very high precision (0.6991 and 0.6912, respectively), but their recall is extremely low, leading to weak F1 scores, which highlights that precision alone is insufficient in this task.
+Table 6.
+Performance comparison of different models on the FinNI subtask.
 Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
 Category Model Precision Recall F1
-Closed-source LLM GPT-4o 0.6105 0.5941 0.6022
-Open-source LLMs Deepseek-V3 0.6329 0.8452 0.7238
-DeepSeek-R1-Distill-Qwen-32B 0.5490 0.2238 0.3180
-Qwen2.5-14B-Instruct 0.3632 0.0018 0.0035
-gemma-2-27b-it 0.5319 0.5490 0.5403
-Llama-3.1-8B-Instruct 0.3346 0.1746 0.2295
-Llama-3.2-3B-Instruct 0.1887 0.1794 0.1839
-Qwen2.5-1.5B-Instruct 0.1323 0.0636 0.0859
-Qwen2.5-0.5B-Instruct 0.0116 0.0027 0.0043
-Financial LLM Fino1-8B 0.3416 0.1481 0.2066
-However, the domain-specific Fino1-8B model, fine-tuned on the financial reasoning QA task, does not perform competitively.
-This indicates that only task-specific training can effectively improve performance on FinNI; pretraining on financial corpora alone offers limited benefit if the task is not well aligned.
-4.2.2 FinCL subtask
-Table 7: Performance comparison of different models on the FinCL subtask.
+Closed-source LLM GPT-4o 0.5893 0.4977 0.5397
+Open-source LLMs Deepseek-V3 0.5886 0.8430 0.6932
+Llama-4-Scout-17B-16E-Instruct 0.4668 0.3164 0.3771
+Llama-3.3-70B-Instruct 0.4826 0.3301 0.3920
+DeepSeek-R1-Distill-Qwen-32B 0.5676 0.1942 0.2894
+Qwen3-32B 0.6991 0.1804 0.2868
+gemma-2-27b-it 0.5060 0.4526 0.4778
+Qwen3-14B 0.6912 0.1487 0.2448
+Llama-3.1-8B-Instruct 0.3874 0.1761 0.2421
+Llama-3.2-3B-Instruct 0.1856 0.1203 0.1460
+Qwen3-1.7B 0.7362 0.0281 0.0541
+Qwen3-0.6B 0.2803 0.0019 0.0038
+Financial LLM Fino1-8B 0.3431 0.1293 0.1878
+Smaller open-source models such as Qwen3-1.7B and Qwen3-0.6B show severe limitations, with F1 below 0.1, underscoring the difficulty of the FinNI subtask in zero-shot settings.
+In contrast, models like gemma-2-27B-it achieve relatively balanced results, ranking third in both recall and F1.
+Finally, the domain-specific Fino1-8B, despite being fine-tuned on financial reasoning QA, underperforms compared to general-purpose LLMs.
+This suggests that domain pretraining on financial text alone provides limited benefit unless the training task aligns closely with FinNI’s requirements.
+4.2.2. FinCL subtask
+Table 7.
+Performance comparison of different models on the FinCL subtask.
 Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
 Category Model Accuracy
-Closed-source LLM GPT-4o 0.1664
-Open-source LLMs Deepseek-V3 0.1715
-DeepSeek-R1-Distill-Qwen-32B 0.1013
-Qwen2.5-14B-Instruct 0.1072
-gemma-2-27b-it 0.1009
-Llama-3.1-8B-Instruct 0.0807
-Llama-3.2-3B-Instruct 0.0375
-Qwen2.5-1.5B-Instruct 0.0419
-Qwen2.5-0.5B-Instruct 0.0246
-Financial LLM Fino1-8B 0.0704
-Table 7 shows the accuracy of different models on the FinCL subtask.
-Overall performance is low, highlighting the challenge of fine-grained concept linking in finance.
-DeepSeek-V3 achieves the highest accuracy (0.1715), followed by GPT-4o, while all other models score below 0.11.
-Even large open-source models like Qwen2.5-14B and Gemma-2-27B underperform, and smaller models perform near random.
-This reflects the difficulty of handling complex taxonomies and subtle financial semantics.
-4.3 Ablation analysis
-[…]
-Table 8: Performance comparison between w/wo our evaluation framework on the FinTagging benchmark dataset.
+Closed-source LLM GPT-4o 0.1829
+Open-source LLMs Deepseek-V3 0.1889
+Llama-4-Scout-17B-16E-Instruct 0.1649
+Llama-3.3-70B-Instruct 0.1318
+DeepSeek-R1-Distill-Qwen-32B 0.1141
+Qwen3-32B 0.1277
+gemma-2-27b-it 0.1099
+Qwen3-14B 0.1144
+Llama-3.1-8B-Instruct 0.0913
+Llama-3.2-3B-Instruct 0.0415
+Qwen3-1.7B 0.0735
+Qwen3-0.6B 0.0414
+Financial LLM Fino1-8B 0.0807
+Table 7 reports the accuracy of different models on the FinCL subtask, which is performed on candidate sets obtained from the best top-200 retrieval results, as shown in Table 16 (Appendix J).
+Although the retrieval stage provides a foundation, the low Acc@k values, especially for table-based entities, directly constrain the FinCL performance.
+Within this reranking formulation, DeepSeek-V3 achieves the best accuracy at 0.1889, closely followed by GPT-4o at 0.1829 and Llama-4-Scout-17B-16E-Instruct at 0.1649.
+Other large open-source models perform notably worse, with Llama-3.3-70B at 0.1318.
+The medium-scale models, such as Qwen3-32B (0.1277), DeepSeek-R1-Distill-Qwen-32B (0.1141), gemma-2-27B-it (0.1099), and Qwen3-14B (0.1144), show similar levels.
+While smaller models, including Llama-3.2-3B and Qwen3-0.6B, perform near random with 0.0415 and 0.0414.
+The domain-specific Fino1-8B also lags at 0.0807, suggesting that domain pretraining without explicit task alignment offers little benefit.
+As a result, overall accuracy remains low, reflecting both the inherent difficulty of handling complex taxonomies and subtle financial semantics, and the cascading effect of retrieval errors that further lowers FinCL performance.
+4.3. Ablation analysis
+We further compare our benchmark against extreme multi-class classification settings.
+The prompt template is shown in Figure 11 in the Appendix, and we select the best-performing model from each category (closed-source, open-source, and financial LLMs).
+The evaluation uses triplet-level ($\{\texttt{Tag},\texttt{Fact},\texttt{Type}\}$) Precision, Recall, and F1.
+Table 8.
+Performance comparison between with/without our evaluation framework on the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark dataset.
 Evaluation Mode Model Precision Recall F1
-FinTagging GPT-4o 0.1075 0.0972 0.1021
-Deepseek-V3 0.1229 0.1519 0.1359
-Fino1-8B 0.0283 0.0138 0.0185
+FinTagging GPT-4o 0.1166 0.0915 0.1026
+Deepseek-V3 0.1187 0.1581 0.1356
+Fino1-8B 0.0336 0.0140 0.0197
 Extreme Classification GPT-4o 0 0 0
@@
 Fino1-8B 0 0 0
-Table 8 shows that all LLMs fail completely with extreme classification, yielding zero precision, recall, and F1, confirming that a single-step choice among thousands of flat labels is not a valid LLM evaluation protocol.
-By decoupling extraction from concept linking and covering the full 10k+ taxonomy, FinTagging produces meaningful scores and thus offers a far more realistic test bed for future model improvements.
-5 Conclusion
-This paper presents FinTagging, a benchmark for evaluating large language models on XBRL tagging of real-world financial reports.
+Table 8 shows that extreme classification is not a valid way to evaluate LLMs on XBRL tagging.
+This single-step formulation requires presenting the full space of thousands of US-GAAP concepts and directly assigning tags to raw text, without first identifying which spans are true numerical entities.
+In practice, this is infeasible: the model cannot realistically process the entire label space and is forced to rely only on its internal knowledge, which leads to meaningless zero-shot predictions.
+As the results indicate, all tested models, including GPT-4o and DeepSeek-V3, collapse to zero precision, recall, and F1 under this setting.
+In contrast, our 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark reformulates the task into a two-stage process that aligns with how tagging is actually performed.
+The FinNI first detects and types coarse-grained numerical entities from text and tables, ensuring that only valid values enter the linking step.
+The FinCL then narrows the taxonomy search space by retrieving a candidate set of relevant tags and requiring the LLM to select the best match using contextual reasoning.
+With this design, models achieve non-trivial performance: for example, DeepSeek-V3 reaches an F1 of 0.1356 and GPT-4o 0.1026.
+These results demonstrate that FinTagging provides a more faithful evaluation of LLMs’ zero-shot tagging capabilities, surfacing meaningful signals where extreme classification fails.
+4.4.
+Error propagation in 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: pipeline
+Table 9 breaks down model accuracy across three evaluation settings: full pipeline, FinCL, and reranking-only in FinCL with guaranteed gold concepts in the top 200 candidates.
+The results reveal clear evidence of cascading errors.
+End-to-end pipeline accuracy is consistently the lowest, since errors from entity extraction (FinNI) and retrieval both propagate into concept linking.
+When gold entities are provided, accuracy improves substantially (e.g., GPT-4o from 0.1052 to 0.1829; DeepSeek-V3 from 0.1074 to 0.1889), confirming that entity identification is a key bottleneck in this context.
+With gold candidates guaranteed in the pool (rerank-only), performance further improves (e.g., GPT-4o 0.2023; DeepSeek-V3 0.2144), showing that narrowing the label space alleviates but does not eliminate difficulty, as semantic ambiguity among fine-grained US-GAAP concepts remains challenging.
+Table 9. Accuracy comparison across the full pipeline, FinCL subtask, and rerank-only settings.
+Model Pipeline FinCL Rerank-only
+GPT-4o 0.1052 0.1829 0.2023
+Deepseek-V3 0.1074 0.1889 0.2144
+Llama-4-Scout-17B-16E-Instruct 0.1045 0.1649 0.1792
+Llama-3.3-70B-Instruct 0.0665 0.1318 0.1482
+DeepSeek-R1-Distill-Qwen-32B 0.0814 0.1141 0.1258
+Qwen3-32B 0.1127 0.1277 0.1402
+gemma-2-27b-it 0.0533 0.1099 0.1210
+Qwen3-14B 0.1069 0.1144 0.1321
+Llama-3.1-8B-Instruct 0.0575 0.0913 0.1076
+Llama-3.2-3B-Instruct 0.0178 0.0415 0.0566
+Qwen3-1.7B 0.0598 0.0735 0.0880
+Qwen3-0.6B 0.0362 0.0414 0.0533
+Fino1-8B 0.0419 0.0807 0.0908
+4.5. Error cases
+To better illustrate the semantic ambiguity challenge, we present a representative error case from our best-performing model, DeepSeek-V3. Consider the following input paragraph:
+“Cash equivalents include term deposits with banks, money market funds, and all highly liquid investments with original maturities of three months or less…… At December 28, 2024, we had restricted cash of \$31 million recorded in other current assets and restricted cash of \$121 million recorded in other non-current assets.”
+This paragraph contains 2 “monetary entities”: 31 and 121.
+The gold concepts and predicted concepts that are assigned to both entities are shown in Table 10.
+For case 1, the model’s choice was likely influenced by the phrase “recorded in current assets”.
+Yet the paragraph describes the overall composition of restricted cash, and the correct concept reflects this aggregate view rather than a single classification.
+For case 2, although both concepts denote non-current restricted cash, the model selected a narrower variant, showing the difficulty of distinguishing between highly similar concepts with subtle differences in scope.
+Table 10. Error cases from DeepSeek-V3 rerank.
+Case 1: \$31m
+Gold Concept us-gaap:RestrictedCashAndCashEquivalentsAtCarryingValue
+Gold Concept Rank 18 / 200 (not in Top-5)
+Model Prediction us-gaap:RestrictedCashAndCashEquivalentsCurrent
+Case 2: \$121m
+Gold Concept us-gaap:RestrictedCashAndCashEquivalentsNoncurrent
+Gold Concept Rank 2 / 200 (in Top-5)
+Model Prediction us-gaap:CashCashEquivalentsRestrictedCashAndRestrictedCashEquivalents
+5. Conclusion
+This paper presents 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, a benchmark for evaluating large language models on XBRL tagging of real-world financial reports.
 The task is divided into two subtasks, financial numeric identification (FinNI) and concept linking (FinCL), to enable fine-grained evaluation of both information extraction and taxonomy alignment.
 Results show that while LLMs generalize well to long-tail entities and perform competitively in zero-shot settings, they struggle with accurate alignment to GAAP concepts.
-This reveals limitations in schema-aware reasoning and highlights the need for better semantic understanding.
-FinTagging offers a foundation for advancing research in XBRL tagging and regulatory reporting.
+This reveals limitations in structure-aware reasoning and highlights the need for better semantic understanding.
+0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: offers a foundation for advancing research in XBRL tagging and regulatory reporting.
 The limitations of our work and directions for future research are discussed in Appendix A.
 Appendix A Limitations
-In this work, we propose FinTagging and conduct a comprehensive analysis of LLM performance on the task.
+In this work, we propose 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: and conduct a comprehensive analysis of LLM performance on the task.
 However, several limitations remain.
-First, we have not yet evaluated a broader range of models, including recent releases such as GPT-4.1, LLaMA-4, and Qwen-3.
+First, we have not yet evaluated the latest models, such as GPT-5, gpt-oss, and Qwen3-235B.
 We are currently conducting these experiments and will report the results in future work.
-Second, although we provide detailed empirical results, we did not perform statistical significance testing, as evaluating the complete benchmark, especially for models like GPT and DeepSeek, requires substantial time.
-This analysis is also in progress and will be included in future updates.
-Third, we collected only a small dataset for fine-tuning PLMs and did not construct task-specific training data for LLMs.
+Second, we collected only a small dataset for fine-tuning PLMs and did not construct task-specific training data for LLMs.
 Future work may explore building LLM-based financial tagging agents through targeted fine-tuning to further enhance performance.
-Appendix B Literature Review
+Appendix B Significance analysis
+To further examine the comparative performance across models, we conducted a pairwise significance analysis, with the results summarized in Figure 4.
+The upper-triangular matrix in SubFigure 4(a) presents the pairwise significance outcomes, where a value of 1 indicates that the row model’s performance is significantly different from the column model.
+In contrast, 0 indicates that the difference is not statistically significant.
+The diagonal elements are set to zero, since self-comparisons are not meaningful.
+(a) Model-vs-Model significance matrix.
+(b) Per-model count of significant pairwise differences.
+Figure 4.
+Visualization of significance analysis across models.
+(a) shows the pairwise-bootstrap significance matrix, while (b) summarizes how many significant differences each model has relative to others.
+As shown in SubFigure 4(b), the number of significant pairwise differences varies across models, reflecting heterogeneous performance behaviors on the evaluated benchmark.
+The models DeepSeek-R1-Distill-Qwen-32B and Qwen3-32B exhibit the largest numbers of significant differences, followed by GPT-4o, Llama-4, and Qwen3-14B, suggesting that their performance diverges more noticeably from other models.
+These results may indicate that models within the Qwen family display stronger variability across comparisons, potentially due to differences in training objectives or architectural configurations.
+In contrast, Deepseek-V3, Llama-3.3-70B-Instruct, gemma-2-27b-it, and Fino1-8B demonstrate relatively fewer significant differences, implying that their performance remains more stable and comparable to other high-performing models.
+Overall, these findings suggest that larger or instruction-tuned models tend to produce more consistent outcomes, while models with smaller sizes or domain-oriented tuning show greater variation and clearer statistical separations under significance testing.
+Overall, this analysis highlights not only which models outperform or underperform others but also which models behave similarly when subjected to statistical evaluation.
+The pairwise significance matrix, therefore, provides an interpretable view of model robustness and performance distinctiveness beyond average accuracy metrics.
+Appendix C Literature Review
 The XBRL provides a comprehensive taxonomy for financial reporting, encompassing thousands of detailed tags corresponding to concepts within financial statements.
 Applying NER to assign XBRL tags is an emerging yet challenging area.
-B.1 XBRL Tagging benchmark
-FiNER systematically benchmarked several neural architectures on the finer-139 dataset to address numeric-heavy XBRL tagging [14].
+C.1. XBRL Tagging benchmark
+FiNER systematically benchmarked several neural architectures on the finer-139 dataset to address numeric-heavy XBRL tagging (Loukas et al., 2022).
 The initial experiments showed that standard BERT underperforms due to subword fragmentation, then the authors introduced pseudo-token strategies replacing numerals with [NUM] or [SHAPE] tokens to stabilize label assignment across fragmented numeric spans.
@@
 Their experiments demonstrated that subword-aware models with numeric-aware pseudo-tokens outperform both word-level BiLSTMs and vanilla BERT, particularly in numeric-heavy contexts, and avoid nonsensical label sequences.
-FNXL extended this benchmarking paradigm to a much larger label space of 2,794 US-GAAP tags, reframing the task as an extreme classification problem [23].
+FNXL extended this benchmarking paradigm to a much larger label space of 2,794 US-GAAP tags, reframing the task as an extreme classification problem (Sharma et al., 2023).
 They compared the FiNER sequence labeling approach with a two-step pipeline that first identifies numeric spans and then assigns labels using AttentionXML.
@@
 Together, these benchmarks reveal the need for context-aware reasoning and label-ranking mechanisms in realistic XBRL tagging scenarios.
-B.2 XBRL Tagging Methods
+C.2. XBRL Tagging Methods
 The previous studies also explored the approaches address scalability, semantic ambiguity, and reasoning gaps in XBRL tagging to improve the performance.
 Saini et al.
-[21] proposed GalaXC is a graph-based extreme classification framework that jointly learns over document-label graphs with per-label attention across multi-hop neighborhoods.
+(Saini et al., 2021) proposed GalaXC is a graph-based extreme classification framework that jointly learns over document-label graphs with per-label attention across multi-hop neighborhoods.
 By integrating label metadata and transitive label correlations, GalaXC outperformed leading deep classifiers by up to 18% in micro-F1 on standard benchmarks and achieved 25% gains in warm-start scenarios where partial labels are available.
 Moreover, Wang et al.
-[28] addressed the practical challenge of custom tag standardization through a semantic similarity pipeline that leverages TF-IDF, Word2Vec, and FinBERT embeddings.
+(Wang, 2023) addressed the practical challenge of custom tag standardization through a semantic similarity pipeline that leverages TF-IDF, Word2Vec, and FinBERT embeddings.
 Although unsupervised, the method was tested across nearly 200,000 custom tags from SEC filings between 2009 and 2022, and showed strong alignment performance, with vector-based mappings identifying viable standard tag candidates for a substantial proportion of non-compliant elements—offering a low-cost, interpretable solution for downstream financial analysis.
-Shifting focus from classification to comprehension, XBRL-Agent evaluated the capabilities of large language models to reason over full XBRL reports [10].
+Shifting focus from classification to comprehension, XBRL-Agent evaluated the capabilities of large language models to reason over full XBRL reports (Han et al., 2024).
 The authors introduced two task types—domain taxonomy understanding and numeric reasoning and found that base LLMs often hallucinated or misinterpreted financial content.
@@
 Collectively, these works broaden the methodological landscape of XBRL tagging from graph-based label propagation and embedding-based normalization to LLM-driven report analysis and point to a hybrid future where structured priors and reasoning tools jointly support accurate, scalable financial information extraction.
-B.3 Financial Evaluation Benchmarks
+C.3. Financial Evaluation Benchmarks
 In parallel to XBRL-specific advances, the financial NLP community has developed comprehensive benchmarks to assess broader capabilities in information extraction, numerical reasoning, and document understanding.
-FiNER-ORD [22] introduced a high-quality, domain-specific NER dataset annotated over financial news, emphasizing general entity types like persons, organizations, and locations.
+FiNER-ORD (Shah et al., 2023) introduced a high-quality, domain-specific NER dataset annotated over financial news, emphasizing general entity types like persons, organizations, and locations.
 While not numerically focused like FiNER-139, it highlights the lexical diversity of financial discourse and establishes a strong baseline for testing pretrained and zero-shot LLMs in real-world financial NER scenarios.
-FinQA [4] pushed toward explainable QA by pairing expert-written questions with annotated multi-step reasoning programs derived from earnings reports.
-ConvFinQA [5] extended this challenge to conversational contexts, simulating real-world question flows over sequential financial queries.
-TAT-QA [35] focused on hybrid tabular-text reasoning and required models to align cell values and document narratives, often involving aggregation, comparison, and unit-scale interpretation.
-Pixiu [33] introduced a broader evaluation framework by releasing FinMA, a financial LLM instruction-tuned across five tasks, and assessing it on a new benchmark covering sentiment classification, QA, summarization, NER, and stock prediction.
-BizBench [12] framed financial QA as program synthesis over realistic, multi-modal contexts, integrating reasoning, code generation, and domain knowledge into a single evaluation pyramid.
+FinQA (Chen et al., 2021) pushed toward explainable QA by pairing expert-written questions with annotated multi-step reasoning programs derived from earnings reports.
+ConvFinQA (Chen et al., 2022) extended this challenge to conversational contexts, simulating real-world question flows over sequential financial queries.
+TAT-QA (Zhu et al., 2021) focused on hybrid tabular-text reasoning and required models to align cell values and document narratives, often involving aggregation, comparison, and unit-scale interpretation.
+Pixiu (Xie et al., 2023) introduced a broader evaluation framework by releasing FinMA, a financial LLM instruction-tuned across five tasks, and assessing it on a new benchmark covering sentiment classification, QA, summarization, NER, and stock prediction.
+BizBench (Koncel-Kedziorski et al., 2023) framed financial QA as program synthesis over realistic, multi-modal contexts, integrating reasoning, code generation, and domain knowledge into a single evaluation pyramid.
 While these benchmarks highlight the growing ability of models to integrate structured and unstructured financial data, they overlook taxonomy-driven fact alignment and do not support the structured output formats required for XBRL tagging.
-Appendix C The Statistics of the Company
-Table 9 lists the 30 publicly traded companies whose 2024 annual 10-K filings were used in our study.
-For each company, we report its Central Index Key (CIK), legal name, stock ticker symbol, and the fiscal year-end filing date.
-These reports were retrieved from the SEC EDGAR system and conform to the XBRL standard.
-The selected filings span a diverse set of industries and ensure a representative sample for analyzing structured and unstructured financial disclosures.
-This company list complements the dataset statistics summarized in Table 2.
-Table 9: Company filing summary (2024) for benchmark data
-CIK Company Name Ticker Filing Date
-0001163739 NABORS INDUSTRIES LTD NBR 2024-12-31
-0001418819 Iridium Communications Inc. IRDM 2024-12-31
-0001993004 NorthWestern Energy Group, Inc. NWE 2024-12-31
-0000024741 CORNING INC /NY GLW 2024-12-31
-0000079282 BROWN & BROWN, INC BRO 2024-12-31
-0001364479 HERC HOLDINGS INC HRI 2024-12-31
-0000936340 DTE ENERGY CO DTM 2024-12-31
-0000085535 ROYAL GOLD INC RGLD 2024-12-31
-0000086312 TRAVELERS COMPANIES, INC. TRV 2024-12-31
-0001968915 PHINIA INC. PHIN 2024-12-31
-0000093751 STATE STREET CORP STT 2024-12-31
-0000106640 WHIRLPOOL CORP WHR 2024-12-31
-0000888491 OMEGA HEALTHCARE INVESTORS INC OHI 2024-12-31
-0001616862 Axalta Coating Systems Ltd. AXTA 2024-12-31
-0000922224 PPL Corp PPL 2024-12-31
-0000066756 ALLETE INC ALE 2024-12-31
-0001411207 Allison Transmission Holdings Inc ALSN 2024-12-31
-0001478242 IQVIA HOLDINGS INC. IQV 2024-12-31
-0001544400 NFiniTi inc. NFTN 2024-12-31
-0000851205 COGNEX CORP CGNX 2024-12-31
-0001318220 Waste Connections, Inc. WCN 2024-12-31
-0000048898 HUBBELL INC HUBB 2024-12-31
-0000937098 TRINET GROUP, INC. TNET 2024-12-31
-0000910108 LXP Industrial Trust LXP 2024-12-31
-0001932393 GE HealthCare Technologies Inc. GEHC 2024-12-31
-0001770787 10x Genomics, Inc. TXG 2024-12-31
-0001026214 FEDERAL HOME LOAN MORTGAGE CORP FMCC 2024-12-31
-0000004904 AMERICAN ELECTRIC POWER CO INC AEP 2024-12-31
-0001637459 Kraft Heinz Co KHC 2024-12-31
-0001713445 Reddit, Inc. RDDT 2024-12-31
-Appendix D Evaluation Metrics
-To provide a fair evaluation of overall benchmark performance, we adopt a set of metrics, focusing primarily on macro-level and micro-level evaluation strategies inspired by the previous work [23].
+Appendix D Data and Ticker Information
+D.1. Overall Data Annotation
+Different from the filtered dataset described in Section 3.3, we construct an overall annotation dataset based directly on the unfiltered raw data.
+As summarized in Table 11, the dataset consists of both sentence-level and table-level instances drawn from the narrative and tabular content of financial reports.
+In total, we annotated 15,986 sentence instances and 12,801 table instances, including 7,768 positive and 8,218 negative sentences, as well as 8,709 positive and 4,092 negative tables.
+On average, textual sentences contain 80.91 tokens, while table sequences are considerably longer, averaging 1,212.42 tokens.
+Across all data, we identified 261,457 numerical entities linked to 3,953 unique US-GAAP concepts.
+Table 11. Statistical information for the original dataset in our benchmark. Tokens are calculated using the “cl100k_base” tokenizer ($\pm$ standard deviation).
+Item Sentence Table
+Positive instances 7,768 8,709
+Negative instances 8,218 4,092
+Avg. Tokens/S $80.91\pm 63.62$ $1212.42\pm 1421.76$
+Avg. Entities/S $1.24\pm 1.82$ $18.87\pm 37.16$
+Avg. Concepts/S $1.24\pm 1.82$ $18.87\pm 37.16$
+Total Entities 261,457
+Unique Concepts 3,953
+D.2. The Statistics of the Tickers
+Figure 5, Figure 6, and Figure 7 illustrate the distribution of 142 tickers across industry sectors, market capitalization categories, and geographic regions, highlighting the diversity of our collected financial reports.
+Considering that practical XBRL tagging practices often vary by industry, company size, and legal jurisdiction, we curated a diverse set of companies covering all 11 major industry sectors.
+The sample maintains a balanced distribution across firm sizes, following market capitalization categories: micro-cap (¡\$300M), small-cap (\$300M–\$2B), mid-cap (\$2B–\$10B), large-cap (\$10B–\$200B), and mega-cap (¿\$200B).
+In addition, we incorporated firms from over 30 states and international jurisdictions to capture regional differences in reporting and tagging conventions.
+This diversity ensures that the benchmark reflects realistic heterogeneity observed in financial disclosures across sectors, scales, and regulatory environments.
+Figure 5. Distribution of Tickers by Industry Sector.
+Figure 6. Distribution of Tickers by Market Cap Size.
+Figure 7. Distribution of Tickers by Geography.
+Appendix E Benchmark Granularity Justification
+Each 10-K report follows the SEC’s XBRL filing standard, where all narrative and tabular disclosures are linked through machine-readable tags defined in the US-GAAP taxonomy.
+While the XBRL framework provides document-level structure via instance, schema, and linkbase files, the tagging of financial values and concepts occurs at a much finer granularity—typically within individual sentences or table rows that contain numeric facts and their contextual descriptions.
+For this reason, our benchmark adopts the sentence and table sequence as the fundamental input unit rather than entire reports. This design choice is motivated by three considerations:
+  • Locality of Semantic Scope: Financial facts and their US-GAAP tags are usually expressed within a single sentence or a bounded table region.
+Modeling at this level aligns the input scope with the tagging scope, reducing noise from unrelated content within long filings.
+  • Computational Efficiency: Full 10-K filings often exceed hundreds of pages and contain heterogeneous sections (e.g., risk factors, management discussion, and financial statements).
+Processing the entire report as one input sequence would exceed model context limits and obscure local numerical relations.
+Segment-level inputs enable fine-grained supervision and efficient evaluation.
+  • Faithful Tag Alignment: Because each numerical entity is explicitly linked to its taxonomy tag in the XBRL instance document, using sentence or table segments ensures a one-to-one correspondence between input content and ground-truth annotations.
+This granularity preserves the original tagging fidelity of the filings.
+Overall, this segmentation strategy allows our benchmark to remain consistent with how XBRL tagging operates in practice (on localized factual statements), while maintaining scalability and interpretability across large collections of filings.
+Appendix F Validation Guideline
+F.1. Task Definition
+Given a context and an extracted triplet consisting of:
+  • an entity,
+  • its entity type,
+  • a US-GAAP concept,
+Your task is to determine whether the triplet is correct.
+Make a binary decision:
+  • 1 for correct,
+  • 0 for incorrect.
+F.2. Validation Procedure
+Follow the rules below, the definitions of entity type are shown in Table 12:
+  1. (1)
+    Entity Check Determine whether the extracted entity is a numerical value.
+    • If yes, proceed to Step 2.
+    • If no, set is_correct = 0.
+  2. (2)
+    Entity Type Validation Verify whether the identified entity type is correct based on the context and the definitions in Table 12.
+    • If yes, proceed to Step 3.
+    • If no, set is_correct = 0.
+  3. (3)
+    US-GAAP Concept Validation Assess whether the assigned US-GAAP concept is appropriate based on the taxonomy tag definitions.
+    • If yes, set is_correct = 1.
+    • If no, set is_correct = 0.
+Table 12. Entity type definitions used for validation.
+Entity Type Definition
+monetaryItemType Financial amounts expressed in currency, such as revenue, profit, or total assets.
+integerItemType Counts of discrete items, such as the number of employees or total transactions.
+perShareItemType Per-share values, such as earnings per share (EPS) or book value per share.
+sharesItemType Counts of shares, such as outstanding shares or ownership stakes.
+percentItemType Ratios or percentages, such as tax rates, growth rates, or discount rates, usually expressed with a percentage symbol (%).
+F.3. Important Instructions
+  1. (1)
+    Non-Arabic Formats Financial numerical entities may appear in word form (e.g., ten million) and must be correctly identified and converted into standard numerical format.
+  2. (2)
+    Magnitude Terms If a number is followed by a magnitude term (e.g., hundred, million, billion), do not expand it into the full numerical value:
+    • Two hundred $\rightarrow$ extract only two, not 200.
+    • 10.6 million $\rightarrow$ extract only 10.6, not 10,600,000.
+  3. (3)
+    Standardization of Format Remove formatting symbols while preserving the numerical value:
+    • Remove currency symbols (e.g., USD).
+    • Remove percentage signs (e.g., %).
+    • Remove commas (e.g., 1,000 $\rightarrow$ 1000).
+Appendix G Evaluation Metrics
+To provide a fair evaluation of overall benchmark performance, we adopt a set of metrics, focusing primarily on macro-level and micro-level evaluation strategies inspired by the previous work (Sharma et al., 2023).
 Macro-level evaluation computes precision, recall, and F1 scores independently for each BIO-concept label derived from the US-GAAP taxonomy, and then averages them without weighting.
-This ensures that each concept, including rare or infrequent ones, contributes equally to the final score—making it especially suitable for domains with skewed label distributions.
+This ensures that each concept, including rare or infrequent ones, contributes equally to the final score, making it especially suitable for domains with skewed label distributions.
 In contrast, micro-level evaluation aggregates token-level true positives, false positives, and false negatives across all labels before computing precision, recall, and F1.
-This approach emphasizes the model’s overall tagging accuracy by treating every token equally, and thus better reflects performance on frequent concepts.
+This approach emphasizes the model’s overall tagging accuracy by treating every token equally and thus better reflects performance on frequent concepts.
 Together, these two metrics provide a balanced view of both per-concept performance and overall tagging quality.
-[…]
-$\text{Precision}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{P}|}$ (5)
-$\text{Recall}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{G}|}$ (6)
-[…]
+For the FinNI subtask, the objective is to extract correct (entity, type) pairs, that is $(\texttt{Fact},\texttt{Type})$, from the financial document.
+Let $\mathcal{G}={(e_{i},l_{i})}$ denote the set of ground-truth (entity, type) pairs, and $\mathcal{P}={(e^{\prime}_{i},l^{\prime}_{i})}$ denote the set of predicted (entity, type) pairs.
+We evaluate the performance based on the following metrics:
+(5) $\text{Precision}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{P}|}$
+(6) $\text{Recall}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{G}|}$
+(7) $\text{F1}=\frac{2\times\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$
 where $(e_{i},l_{i})=(e^{\prime}_{j},l^{\prime}_{j})$ if and only if both the entity span $e$ and its assigned type $l$ exactly match.
 For the FinCL subtask, Given a set of queries $\mathcal{Q}=\{q_{1},q_{2},\dots,q_{N}\}$, where each $q_{i}$ is associated with a ground-truth concept $c_{i}^{*}$ and a predicted concept $\hat{c}_{i}$, the accuracy is defined as:
-$\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\delta(\hat{c}_{i},c_{i}^{*})$ (8)
-where $\delta(\hat{c}_{i},c_{i}^{*})=1$ if $\hat{c}_{i}=c_{i}^{*}$, and 00 otherwise.
-Appendix E Evaluation Models Details
+(8) $\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\delta(\hat{c}_{i},c_{i}^{*})$
+where $\delta(\hat{c}_{i},c_{i}^{*})=1$ if $\hat{c}_{i}=c_{i}^{*}$, and 0 otherwise.
+Appendix H Evaluation Models Details
 Table 13 provides an overview of the models evaluated in this study, categorized by openness, domain specialization, and architectural foundation. The evaluation covers a diverse range of models:
-  • Closed-source LLMs: We include GPT-4o [11], accessed via OpenAI’s API, as a representative of cutting-edge proprietary models with demonstrated performance across a variety of NLP tasks.
+  • Closed-source LLMs: We include GPT-4o (Hurst et al., 2024), accessed via OpenAI’s API, as a representative of cutting-edge proprietary models with demonstrated performance across a variety of NLP tasks.
 Although model size details are undisclosed, GPT-4o serves as an upper-bound reference in our benchmark.
-  • Open-source LLMs: This group encompasses recent, high-performing open models such as DeepSeek-V3 (685B) [13], DeepSeek-R1-Distill-Qwen (32B) [6], and multiple variants of Qwen2.5-Instruct  [34](ranging from 0.5B to 14B).
-We also include LLaMA-3.2 and 3.1 variants  [9] (3B and 8B), as well as Google’s Gemma-2-27B  [26], to ensure architectural diversity and scalability comparison.
+  • Open-source LLMs: This group encompasses recent, high-performing open models such as DeepSeek-V3 (685B) (Liu et al., 2024), DeepSeek-R1-Distill-Qwen (32B) (DeepSeek-AI, 2025), and multiple variants of Qwen3  (Team, 2025)(ranging from 0.6B to 32B).
+We also include Llama-3.3, Llama-3.2, and 3.1 variants  (Grattafiori et al., 2024) (70B, 3B, and 8B), Llama-4-Scout-17B-16E-Instruct (Meta, 2025)(109B), as well as Google’s Gemma-2-27B  (Team, 2024), to ensure architectural diversity and scalability comparison.
 These models are primarily instruction-tuned and optimized for general-purpose NLP tasks.
-  • Financial-specific LLMs: We evaluate Fino1-8B  [19], a domain-specialized model trained on financial corpora, designed to better capture the terminology and structure unique to financial disclosures.
+  • Financial-specific LLMs: We evaluate Fino1-8B  (Qian et al., 2025), a domain-specialized model trained on financial corpora, designed to better capture the terminology and structure unique to financial disclosures.
 This category allows us to assess the benefits of domain adaptation in complex tagging and reasoning tasks.
-  • Pretrained Language Models (PLMs): To establish strong baselines, we include non-generative encoder models: BERT-large  [7], FinBERT  [1], and SECBERT  [14].
+  • Pretrained Language Models (PLMs): To establish strong baselines, we include non-generative encoder models: BERT-large  (Devlin et al., 2018), FinBERT  (Araci, 2019), and SECBERT  (Loukas et al., 2022).
 These models have been widely used in prior financial NLP tasks and allow for a comparative analysis between generative LLMs and traditional pretrained models in terms of domain understanding and structured output capability.
 Together, these models offer a comprehensive evaluation spectrum, from general-purpose to domain-specific, encoder-based to decoder-based, and open to closed source, facilitating an in-depth assessment of their performance across our proposed benchmark tasks.
-Table 10: Model Categories and Corresponding Repositories
+Table 13. Model categories and corresponding repositories.
 Model Size Source
@@
 DeepSeek-V3 685B deepseek-ai/DeepSeek-V3
+Llama-4-Scout-17B-16E-Instruct 109B meta-llama/Llama-4-Scout-17B-16E-Instruct
+Llama-3.3-70B-Instruct 70B meta-llama/Llama-3.3-70B-Instruct
 DeepSeek-R1-Distill-Qwen 32B deepseek-ai/DeepSeek-R1-Distill-Qwen-32B
-Qwen2.5-Instruct 0.5B, 1.5B, 14B Qwen/Qwen2.5-*B-Instruct
-Llama-3.2-Instruct 3B meta-llama/Llama-3.2-3B-Instruct
-Llama-3.1-Instruct 8B meta-llama/Llama-3.1-8B-Instruct
+Qwen3-32B 32B Qwen/Qwen3-32B
 gemma-2-27b-it 27B google/gemma-2-27b-it
+Qwen3-14B 14B Qwen/Qwen3-14B
+Llama-3.1-8B-Instruct 8B meta-llama/Llama-3.1-8B-Instruct
+Llama-3.2-3B-Instruct 3B meta-llama/Llama-3.2-3B-Instruct
+Qwen3-1.7B 1.7B Qwen/Qwen3-1.7B
+Qwen3-0.6B 0.6B Qwen/Qwen3-0.6B
 Financial-specific Large Language Models
@@
 SECBERT $\sim$110M nlpaueb/sec-bert-base
-Appendix F The details for the fine-tuning PTMs
-F.1 Training data collection and processing
-Similar to the collection process for the FinTagging benchmark data, we gathered an additional 10 annual 10-K financial reports filed with the SEC for the period from February 13, 2024, to February 13, 2025, as summarized in Table 11.
+Appendix I The details for the fine-tuning PTMs
+I.1. Training data collection and processing
+Similar to the collection process for the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark data, we gathered an additional 10 annual 10-K financial reports filed with the SEC for the period from February 13, 2024, to February 13, 2025, as summarized in Table 14.
 These reports contain a total of 33,848 standard taxonomy-tagged facts.
@@
 The companies included in this dataset follow the XBRL standard, ensuring comprehensive coverage for training PTMs.
-Table 11: Financial report statistics summary for raw training data
+Table 14. Financial report statistics summary for raw training data
 Item Information
@@
 Overall, the annotated dataset includes 25,199 entities, covering 1,435 unique US-GAAP concepts.
-Table 12: The statistical information for training data. Using the “cl100k_base” tokenizer to calculate tokens ($\pm$ standard deviation).
+Table 15. Statistics of training data (tokens calculated with “cl100k_base” tokenizer, $\pm$ standard deviation).
 Structure Pos/Neg #Instance Avg. Tokens/S Avg. Entities/S Avg. Concepts/S Total Entities Unique Concepts
-Sentence Positive 558 $84.24\pm 69.29$ $1.22\pm 1.78$ $1.22\pm 1.78$ 25,199 1,435
+Sentence Positive 558 $84.24\pm 69.29$ $2\pm 1.78 1.22±1$$2\pm 1.78$ 25,199 1,435
 Negative 558
@@
 Specifically, B denotes the beginning of an entity phrase, I marks the continuation (inside) of an entity phrase, and O indicates tokens outside of any entity.
-As shown in Figure 4, "4.9" and "4.5" are single-token numerical entities labeled only with a B prefix (e.g., “B-us-gaap:AccountsReceivableNetNoncurrent”).
+As shown in Figure 8, ”4.9” and ”4.5” are single-token numerical entities labeled only with a B prefix (e.g., “B-us-gaap:AccountsReceivableNetNoncurrent”).
 To comprehensively cover all US-GAAP tags, we combine the entire set of 17,388 tags from the US-GAAP 2024 taxonomy with the BIO labeling scheme to construct an extreme classification label space, resulting in 34,777 unique entity labels ($2\times 17388+1$).
-Figure 4: An example for training set annotation with BIO scheme.
+Figure 8. An example for training set annotation with BIO scheme.
 After constructing the training set, we reconstruct the testing set from the original benchmark dataset. The training settings are detailed below.
-F.2 Training settings
-We fine-tune three pretrained models, BERT-large [7], FinBERT [1], and SECBERT [14], on our training set using the HuggingFace Transformers library.
+I.2. Training settings
+We fine-tune three pretrained models, BERT-large (Devlin et al., 2018), FinBERT (Araci, 2019), and SECBERT (Loukas et al., 2022), on our training set using the HuggingFace Transformers library.
 All models are trained with a batch size of 4, a learning rate of 3e-5, and for 20 epochs.
@@
 All experiments are run under a fixed random seed to ensure reproducibility.
-Appendix G Retrieval Results in FinCL subtask
+Appendix J Retrieval Results in FinCL subtask
 We investigate the impact of different context construction strategies for the queried entity at the retrieval stage.
@@
 In contrast, the SAC strategy builds context based on the structural location of the entity: if the entity appears in a sentence, the entire sentence is used; if the entity appears in a table, we linearize the entire row into a Markdown-style key-value format to serve as the context.
-Table 13: The Acc@k performance of retrieval for different context construction strategies on the FinCL task.
-Here, FWC denotes the fixed window-based strategy, while SAC denotes the structure-aware approach that respects sentence or row boundaries.
+Table 16. Acc@k retrieval performance on the FinCL task. FWC: fixed window-based; SAC: structure-aware.
 Strategy Structure Acc@1 Acc@10 Acc@20 Acc@50 Acc@100 Acc@150 Acc@200
-FWC Sentence 0.0677 0.2594 0.3554 0.4650 0.5515 0.6002 0.6404
-Table 0.000 0.0030 0.0031 0.0036 0.0050 0.0064 0.0097
+FWC Sentence 0.0658 0.2614 0.3562 0.4650 0.5618 0.6114 0.6502
+Table 0.0000 0.0029 0.0038 0.0042 0.0050 0.0065 0.0089
 Overall 0.0055 0.0237 0.0316 0.0409 0.0492 0.0544 0.0608
-SAC Sentence 0.0689 0.2643 0.3521 0.4711 0.5611 0.6096 0.6522
-Table 0.0147 0.0867 0.1240 0.1855 0.2353 0.2632 0.2866
-Overall 0.0191 0.1011 0.1424 0.2086 0.2617 0.2913 0.3163
-From Table 13, we found the SAC strategy consistently outperforms FWC, particularly in the table context, highlighting the importance of aligning the context window with the underlying structural unit.
-For sentence-based entities, both strategies yield comparable results, with SAC achieving slightly higher Acc@1 (0.0689 vs.
-0.0677) and showing marginal improvements across all cutoff values.
-This suggests that even for relatively unstructured text, preserving sentence boundaries may provide minor benefits over a fixed-length window.
+SAC Sentence 0.0696 0.2742 0.3631 0.4798 0.5701 0.6125 0.6535
+Table 0.0159 0.0872 0.1245 0.1938 0.2452 0.2723 0.2959
+Overall 0.0202 0.1152 0.1534 0.2188 0.2727 0.3012 0.3274
+From Table 16, we observe that the SAC strategy consistently outperforms FWC, particularly in the table context, highlighting the importance of aligning the context window with the underlying structural unit.
+For sentence-based entities, both strategies yield comparable results, with SAC achieving slightly higher Acc@1 (0.0696 vs.
+0.0658) and showing marginal improvements across all cutoff values (e.g., Acc@100 of 0.5701 vs.
+0.5618).
+This suggests that even for relatively unstructured text, preserving sentence boundaries provides minor benefits over fixed-length windows.
 In contrast, for table-based entities, the performance gap is substantial.
-SAC achieves significantly higher retrieval accuracy (e.g., Acc@100 of 0.2353 vs.
+SAC achieves significantly higher retrieval accuracy (e.g., Acc@100 of 0.2452 vs.
 0.0050), indicating that row-level context is far more informative than arbitrary character windows when dealing with tabular structures.
 FWC performs poorly in this setting, likely due to the fragmented and semantically sparse nature of partial table text.
-When aggregating results across both structures, SAC outperforms FWC by a wide margin at all retrieval depths (e.g., Acc@200 of 0.3163 vs.
+When aggregating results across both structures, SAC outperforms FWC by a wide margin at all retrieval depths (e.g., Acc@200 of 0.3274 vs.
 0.0608).
-These results underscore the importance of structure-aware context construction, especially in scenarios where inputs span multiple formats such as sentences and tables.
-Appendix H Prompt Templates
-H.1 Template for FinNI subtask
-Prompt Template You are a financial information extraction expert specializing in identifying financial numerical entities in XBRL reports.
+These findings underscore the importance of structure-aware context construction, especially in scenarios where inputs span multiple formats such as sentences and tables.
+Prompt Template for FinNI Subtask You are a financial information extraction expert specializing in identifying financial numerical entities in XBRL reports.
 Your task is to extract all such numerical entities from the provided text or serialized <table></table> data and classify them into one of five categories: - "integerItemType": Counts of discrete items, such as the number of employees or total transactions.
@@
 (5) Output the extracted financial entities in JSON list format without explanations, structured as follows: {"result":[{"Fact": <Extracted Numerical Entity>, "Type": <Identified Entity Type>}]} Input: {text/table} Output:
-H.2 Template for ablation
-Prompt Template You are an XBRL tagging expert specializing in annotating financial numerical facts in XBRL reports.
-Your task is to (1) extract all such numerical entities from the provided text or serialized <table></table> data, (2) classify them into one of five categories, and (3) assign an appropriated US-GAAP tag to each entity.
+Figure 9. Prompt template used for the FinNI subtask.
+Prompt Template for FinCL Subtask (Reranking) You are a financial tagging assistant trained in US-GAAP taxonomy.
+Given a query consisting of an entity, its type, its surrounding context, and the source format (either text or table), your task is to select the single most appropriate US-GAAP tag from a list of 200 candidate tags.
+Make your decision by carefully analyzing the meaning and context of the entity and matching it with the semantics of the tags.
+Only output one tag, the best match.
+Do not explain or list multiple tags.
+The output is a JSON format: {"result": <the best matched tag>}.
+Input Query: <entity> + <entity type> + <context> Candidate Tags: {Top 200 US-GAAP tags} Answer:
+Figure 10. Prompt template used for the Reranking statge in the FinCL subtask.
+Prompt Template for Ablation You are an XBRL tagging expert specializing in annotating financial numerical facts in XBRL reports.
+Your task is to (1) extract all such numerical entities from the provided text or serialized <table></table> data, (2) classify them into one of five categories, and (3) assign an appropriate US-GAAP tag to each entity.
 Categories: - "integerItemType": Counts of discrete items, such as the number of employees or total transactions.
@@
 (6) Output the extracted financial entities in JSON list format without explanations, structured as follows: {"result":[{"Fact": <Extracted Numerical Entity>, "Type": <Identified Entity Type>, "Tag": <Assigned US-GAAP tag>}]} Input: {text/table} Output:
-Appendix I Potential Societal Impacts of the Work
-I.1 Potential Positive Societal Impacts
-This work introduces FinTagging, a full-scope, structure-aware benchmark for evaluating large language models on XBRL-based financial reporting.
-By supporting realistic tagging scenarios grounded in US-GAAP taxonomy, this benchmark:
-  • Promotes automation and transparency in financial disclosure, which can reduce manual errors and increase the reliability of corporate filings.
-  • Enables more accurate, scalable tools for financial regulators, auditors, and analysts, potentially improving oversight and decision-making in global markets.
-  • Encourages the development of LLMs with stronger schema-aware reasoning, a critical step toward trustworthy AI systems in high-stakes financial environments.
-Through open access to annotated datasets and evaluation code, FinTagging also advances reproducible research and supports the broader community in benchmarking financial language models responsibly.
-I.2 Potential Negative Societal Impacts
-While FinTagging provides valuable tools for model evaluation, several risks should be noted:
-  • Misuse and Misinterpretation: Benchmark performance may be misinterpreted as real-world readiness.
-Deployment of LLMs without sufficient regulatory safeguards could lead to incorrect tagging, misreporting, or financial misinformation.
-  • Coverage Bias: The benchmark focuses on US-GAAP financial reports and may underrepresent other accounting standards or financial contexts, potentially limiting its applicability in non-U.S.
-jurisdictions.
-  • Overreliance on Automation: Improved tagging accuracy may incentivize overautomation in financial reporting pipelines, increasing the risk of undetected errors if human oversight is reduced.
-To mitigate these concerns, we recommend that future work includes broader financial contexts and integrates FinTagging into a framework with human-in-the-loop validation, fairness checks, and error auditing.
-Ongoing collaboration with domain experts and regulators is also essential.
+Figure 11. Prompt template used for ablation study.

### v3

@@
 
-Abstract.
-Accurately understanding numbers from financial reports is fundamental to how markets, regulators, algorithms, and normal people read the economy and the world, yet even with XBRL (eXtensible Business Reporting Language) designed to tag every figure with standardized accounting concepts, mapping thousands of facts to over 10,000 U.S.
-GAAP concepts remains costly, inconsistent, and error-prone.
-Existing benchmarks define tagging as flat, single-step, extreme classification over small subsets of US-GAAP concepts, overlooking both the taxonomy’s hierarchical semantics and the structured nature of real tagging, where each fact must be represented as a contextualized multi-field output.
-These simplifications prevent fair evaluation of large language models (LLMs) under realistic reporting conditions.
-To address these gaps, we introduce 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first comprehensive benchmark for structure-aware and full-scope XBRL tagging, designed to evaluate LLMs’ ability to extract and align financial facts through numerical reasoning and taxonomy alignment across text and tables.
-We define two subtasks: FinNI for numeric identification, which extracts numerical entities and their types from XBRL reports, and FinCL for concept linking, which maps each extracted entity to the corresponding concept in the full US-GAAP taxonomy.
-Together, these subtasks produce a structured representation of each financial fact.
-We evaluate diverse LLMs under zero-shot settings and analyze their performance across both subtasks and overall tagging accuracy.
-Results show that LLMs generalize well in numeric identification but struggle with fine-grained concept linking, revealing current limitations in structure-aware reasoning for accurate financial disclosure.
-All code and datasets are available on GitHub [^1] https://github.com/The-FinAI/FinTagging/ and Hugging Face [^2] https://huggingface.co/collections/TheFinAI/fintagging-68270132372c6608ac069bef.
-XBRL tagging, Benchmark, Large language model, Information extraction, Information retrieval, Reranking
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Test collectionsccs: Information systems Information extractionccs: Information systems Clustering and classification
-1. Introduction
-Automated tagging is essential in financial reporting, converting structured content such as tables and text into machine-readable data by linking each number to its meaning.
-Globally, over 2 million companies publish financial reports disclosing earnings, expenses, and liabilities, which provide essential information for investors, regulators, and analysts.
-However, inconsistent terminology poses challenges for reliable interpretation.
-To address this, the eXtensible Business Reporting Language (XBRL) was introduced in 1999 as a global standard for machine-readable financial data (Richards et al., 2006).
-By applying structured tags, as shown in Figure 1, XBRL transforms fragmented disclosures into a consistent format, enabling seamless data exchange, cross-entity comparisons, and large-scale analysis.
-However, accurately tagging over 2,000 numerical facts per report to more than 10,000 standardized concepts remains labor-intensive and error-prone.
-In 2023 alone, over 6,500 tagging errors were identified across 33,000 SEC filings [^3] https://xbrl.us/wp-content/uploads/2023/03/DQC-SECMeetingNotes-20240314.pdf.
-Figure 1.
-An example of Financial Tagging in Realistic, where the orange numbers mark facts to be tagged, black dashed arrows show the structured outputs after tagging, and blue dashed arrows denote the US-GAAP taxonomy referenced during tagging.
-Existing approaches to automated XBRL tagging rely on task-specific models and pretrained language models (PLMs) (Loukas et al., 2022; Sharma et al., 2023), but they require costly annotations, struggle to generalize to new filings, and perform poorly as taxonomies scale.
-In contrast, recent large language models (LLMs) show strong zero- and few-shot reasoning without task-specific supervision (Wei et al., 2022; Cao et al., 2025; Wang et al., 2025; Peng et al., 2025), yet their application to XBRL tagging remains underexplored due to two major limitations (Table 1).
-First, task framing is oversimplified: prior benchmarks like FiNER (Loukas et al., 2022) and FNXL (Sharma et al., 2023) treat tagging as extreme multi-class classification over flat label sets, offering little context for disambiguating fine-grained financial facts.
-These datasets typically cover only 1k+ concepts, leaving most of the 10k+ US-GAAP taxonomy untested.
-As shown in Table 8, SOTA LLMs such as DeepSeek-V3 (Liu et al., 2024) and GPT-4o (Hurst et al., 2024) achieve 0.0 precision, recall, and F1 under this setting.
-Second, structured data is ignored: existing datasets exclude tables, despite their central role in financial reporting.
-Many key facts appear in structured formats, as shown in recent financial QA benchmarks (Chen et al., 2021; Zhu et al., 2021; Mathew et al., 2021), making this omission a key gap between benchmarks and real-world tagging needs.
-Table 1.
-Detailed comparison of financial NLP benchmarks across task types, sources, and structural capabilities.
+Abstract
+Accurate interpretation of numerical data in financial reports is critical for markets and regulators.
+Although XBRL (eXtensible Business Reporting Language) provides a standard for tagging financial figures, mapping thousands of facts to over 10k US-GAAP concepts remains costly and error-prone.
+Existing benchmarks oversimplify this task as flat, single-step classification over small subsets of concepts, ignoring the hierarchical semantics of the taxonomy and the structured nature of financial documents.
+Consequently, these benchmarks fail to evaluate Large Language Models (LLMs) under realistic reporting conditions.
+To bridge this gap, we introduce 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first comprehensive benchmark for structure-aware and full-scope XBRL tagging.
+We decompose the complex tagging process into two subtasks: (1) FinNI (Financial Numeric Identification), which extracts entities and types from heterogeneous contexts (text and tables); and (2) FinCL (Financial Concept Linking), which maps extracted entities to the full US-GAAP taxonomy.
+This two-stage formulation enables a fair assessment of LLMs’ capabilities in numerical reasoning and taxonomy alignment.
+Evaluating diverse LLMs in zero-shot settings reveals that while models generalize well in extraction, they struggle significantly with fine-grained concept linking, highlighting critical limitations in domain-specific structure-aware reasoning.
+Code is available at GitHub  [^1] https://github.com/The-FinAI/FinTagging and datasets are available at Hugging Face  [^2] https://huggingface.co/collections/TheFinAI/fintagging.
+0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:: Benchmarking LLMs for Extracting and Structuring Financial Information
+Yan Wang1, Lingfei Qian1, Xueqing Peng1, Yang Ren1, Yi Han2, Keyi Wang3, Dongji Feng4, Fengran Mo5, Shengyuan Lin6, Qinchuan Zhang7, Kaiwen He3, Chenri Luo3, Jianxing Chen3, Junwei Wu3, Chen Xu1, Ziyang Xu1, Jimin Huang1thanks:  Corresponding author, Guojun Xiong8, Xiao-Yang Liu3, Qianqian Xie1, Jian-Yun Nie5 1The Fin AI, 2Georgia Institute of Technology, 3Columbia University, 4California State University 5University of Montreal, 6Carnegie Mellon University, 7Rensselaer Polytechnic Institute, 8Harvard University Correspondence: jimin.huang@thefin.ai
+1 Introduction
+Although LLMs excel at question answering qian2025fino1, information extraction balasubramanian2025leveraging, and long-document summarization yuan2025strucsum, their ability to perform financial tagging remains underexplored.
+Unlike open-ended generation, tagging requires models to commit to a precise semantic interpretation for each reported number, grounding it in dispersed evidence across tables and surrounding text, and aligning it to a standardized concept that downstream systems can interpret consistently.
+Each year, over 2 million companies publish financial reports disclosing earnings, expenses, and liabilities, providing essential information for investors, regulators, and analysts.
+However, inconsistent terminology complicates reliable interpretation.
+XBRL, introduced in 1999, defines the global standard for machine-readable financial reporting richards2006introduction and is now adopted in 65 countries [^3] https://www.xbrl.org/the-standard/what/what-is-xbrl/.
+In practice, XBRL tagging requires grounding thousands of reported numbers per filing in mixed table and text evidence and aligning them to over 17,000 standardized taxonomy concepts(Figure 1), a workflow that remains largely manual and error-prone [^4] https://xbrl.us/wp-content/uploads/2023/03/DQC-SECMeetingNotes-20240314.pdf.
+Figure 1: An example of Financial Tagging in realistic scenarios, where the blue numbers mark facts to be tagged.
+Nevertheless, current XBRL tagging benchmarks often emphasize what is easiest to annotate rather than what is hardest to understand.
+Many reduce tagging to classification over a truncated taxonomy and remove table structure that carries substantial semantic evidence in real filings (Table 1).
+As a result, existing datasets typically cover only a few thousand common concepts instead of the full 17,000+ taxonomy, which limits evaluation on rare but important concepts and makes it difficult to assess generalization under long-tailed label distributions.
+In addition, prior work commonly formulates tagging as closed-set extreme classification loukas2022finer; sharma2023financial.
+This framing places heavy weight on recalling label identifiers and becomes brittle when the full taxonomy cannot be represented in-context, increasing the likelihood of invalid or hallucinated tags.
+Consistent with this concern, our ablations show near-zero performance even for strong LLMs such as DeepSeek-V3 liu2024deepseek and GPT-4o hurst2024gpt under full-taxonomy settings.
+Finally, existing benchmarks focus primarily on narrative text and omit tables, despite tables being the main carrier of numerical disclosures.
+Consequently, they do not evaluate the joint reasoning over multi-level tabular structure and surrounding text that has been studied in financial QA chen2021finqa; zhu2021tat and is required for realistic end-to-end tagging.
+Table 1: Detailed comparison of financial NLP benchmarks across task types, sources, and structural capabilities.
 “Num.Reasoning” indicates the numerical reasoning.
 “Struct.IE” denotes the structured information extraction.
-Benchmark Scenario Data Source Task Modality
-#Entity
-Label
-#Taxonomy
-Label
-Num.
-Reasoning?
-Struct.
-IE?
-Concept
-Linking?
-FinQA (Chen et al., 2021) Decision making SEC 10-K QA text/table 0 0
-ConvFinQA (Chen et al., 2022) Decision making SEC 10-K QA text/table 0 0
-TAT-QA (Zhu et al., 2021) Financial analysis Chinese financial reports QA text/table 0 0
-DocVQA (Mathew et al., 2021) Enterprise automation Financial document images QA text/image 0 0
-FiNER-ORD (Shah et al., 2023) Financial tagging Chinese financial disclosures Classification text 49 0
-FinRED (Sharma et al., 2022) Knowledge graph construction English financial news Classification text 38 0
-FiNER (Loukas et al., 2022) XBRL tagging SEC 10-K Extreme classification text 0 139 / 17,388
-FNXL (Sharma et al., 2023) XBRL tagging SEC 10-K Extreme classification text 0 2,800 / 17,388
-FinTagging (ours) Financial & XBRL tagging SEC 10-K IE + Alignment text/table 5 17,388 / 17,388
-To address these issues, we present 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first LLM-ready benchmark for full-scope, structure-aware XBRL tagging, which challenges models to perform end-to-end fact extraction and taxonomy alignment on corporate filings.
-As shown in Table 1, unlike prior work (Shah et al., 2023; Loukas et al., 2022; Sharma et al., 2023) that frames tagging as flat multi-class classification and focuses solely on narrative text, 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: requires models to jointly extract financial facts and align them with the US-GAAP taxonomy across both unstructured text and structured tables.
-This is the first LLM-oriented formulation that scales to the full set of 10k+ taxonomy labels while avoiding the intractability of single-step classification, enabling more fine-grained evaluation of model performance.
-In collaboration with financial reporting experts, we decompose the task into two components: the first, structured numerical information extraction, assesses a model’s ability to identify key financial facts from both text and tables; the second, fine-grained concept linking, evaluates its ability to map each fact to the correct taxonomy concept among all semantically similar candidates.
-Unlike prior benchmarks limited to 1,000 frequent concepts, our setting covers the entire taxonomy, exposing the limitations of frequency-biased classification and demanding precise semantic reasoning (Bovee et al., 2005).
-We construct two evaluation sets: FinNI-eval [^4] https://huggingface.co/datasets/TheFinAI/FinNI-eval-v1 for numerical fact extraction and FinCL-eval [^5] https://huggingface.co/datasets/TheFinAI/FinCL-eval-v1 for concept linking, both derived from real XBRL submissions and annotated with gold-standard mappings.
-We also introduce a unified evaluation framework that jointly measures extraction accuracy and concept alignment, providing a rigorous zero-shot assessment of LLMs’ capabilities for practical, table-aware XBRL tagging.
-We evaluate 13 state-of-the-art LLMs under a zero-shot setting on three fronts: (1) end-to-end macro-F1 over the full 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark, (2) subtask-specific performance on FinNI and FinCL, and (3) ablation of our unified extraction-and-alignment evaluation framework.
-DeepSeek-V3 (Liu et al., 2024) and GPT-4o (Hurst et al., 2024) achieve the highest macro-F1 scores, indicating that our benchmark design enables strong handling of both frequent and rare financial tags, an improvement over traditional PLM baselines evaluated under token-classification settings.
-In subtask analyses, these LLMs excel at numerical information extraction (FinNI) but continue to struggle with precise concept linking (FinCL), underscoring the remaining challenge of fine-grained semantic alignment.
-Crucially, our ablation study shows that, without the joint evaluation framework, even top LLMs produce invalid tagging outputs, highlighting the framework’s essential role in realistic assessment of XBRL tagging readiness.
-We conclude our main contributions as follows: (1) We introduce 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first LLM-ready benchmark for full-scope, table-aware XBRL tagging with a two-stage pipeline, so that modern LLMs can be probed in a true zero-shot setting.
-For each stage, we release a new, professionally annotated evaluation set: FinNI-eval and FinCL-eval.
-(2) We conduct an extensive evaluation of state-of-the-art LLMs in zero-shot settings, systematically analyzing their performance across information extraction and semantic alignment subtasks, as well as overall tagging accuracy.
-(3) Our experimental results reveal a substantial performance gap between LLMs and the task requirements, particularly in fine-grained semantic alignment.
-This highlights the limitations of current LLMs in complex financial applications and underscores the need for continued advancements in task-specific adaptation.
-2. Related Work
-XBRL Tagging Benchmarks: Previous benchmarks have framed XBRL tagging as a large-scale classification problem, primarily focusing on entity recognition.
-FiNER (Loukas et al., 2022) introduced a dataset of 1.1 million sentences from SEC filings, annotated with 139 XBRL entity types, addressing challenges like context-sensitive numeric entities and financial domain-specific expressions.
-Building on this, Sharma et al.
-(Sharma et al., 2023) proposed the Financial Numeric Extreme Labelling (FNXL) dataset, expanding the label space to approximately 2,800 XBRL tags and employing extreme classification techniques to handle the increased scale.
-While these works emphasize the complexity of financial taxonomies, their flat extreme classification approach overlooks the reasoning and alignment needed for realistic XBRL tagging, limiting their suitability for evaluating large language models.
-XBRL Tagging Methods: To improve XBRL tagging accuracy, prior work has incorporated structured knowledge and domain-specific modeling.
-Saini et al.
-(Saini et al., 2021) proposed GalaXC, a graph neural network leveraging label hierarchies for improved classification.
-Ma et al.
-(Ma et al., 2022) enriched transformer models with taxonomy definitions to disambiguate semantically similar tags.
-Loukas et al.
-(Loukas et al., 2022) introduced SECBERT and numeral-aware input transformations to enhance robustness.
-For extreme multi-label tagging, Sharma et al.
-(Sharma et al., 2023) applied AttentionXML to focus on context-relevant segments.
-Wang (Wang, 2023) aligned custom tags with standard taxonomies via semantic similarity using TF-IDF, Word2Vec, and FinBERT.
-More recently, Han et al.
-(Han et al., 2024) developed XBRL-Agent, an LLM-based system for extracting insights from filings.
-These efforts highlight the shift toward taxonomy-aware and LLM-enhanced solutions for XBRL tagging.
-Financial Evaluation Benchmarks: Broader financial NLP benchmarks address information extraction, reasoning, and document understanding tasks.
-FiNER-ORD (Shah et al., 2023) and FinRED (Sharma et al., 2022) focus on entity recognition and relation extraction, respectively.
-BizBench (Koncel-Kedziorski et al., 2023) assesses the quantitative-reasoning ability of LLMs for both business and finance.
-Pixiu (Xie et al., 2023) evaluates LLMs across classification, QA, and summarization, while FinQA (Chen et al., 2021) and ConvFinQA (Chen et al., 2022) focus on numerical reasoning over text and tables.
-TAT-QA (Zhu et al., 2021) and DocVQA (Mathew et al., 2021) target table-text joint understanding.
-These benchmarks highlight the importance of integrating structured and unstructured information, yet they overlook the taxonomy-driven fact alignment and do not support the structured output required for XBRL tagging.
-3.
-0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:
-3.1. Task Formulation
+Benchmark Scenario Data Source Task Modality #Entity Label #Taxonomy Label Num. Reasoning? Struct. IE? Concept Linking?
+FinQA chen2021finqa Decision making SEC 10-K QA text/table 0 0
+ConvFinQA chen2022convfinqa Decision making SEC 10-K QA text/table 0 0
+TAT-QA zhu2021tat Financial analysis Chinese financial reports QA text/table 0 0
+DocVQA mathew2021docvqa Enterprise automation Financial document images QA text/image 0 0
+FiNER-ORD shah2023finer Financial tagging Chinese financial disclosures Classification text 49 0
+FinRED sharma2022finred Knowledge graph construction English financial news Classification text 38 0
+FiNER loukas2022finer XBRL tagging SEC 10-K Extreme classification text 0 139 / 17,688
+FNXL sharma2023financial XBRL tagging SEC 10-K Extreme classification text 0 2,800 / 17,688
+FINTAGGING Financial & XBRL tagging SEC 10-K IE + Alignment text/table 5 17,688 / 17,688
+To address these issues, we introduce 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first LLM-oriented benchmark that evaluates whether models can recover the meaning of financial facts by extracting numerical values from mixed table and text inputs and aligning them to the full US-GAAP taxonomy.
+Unlike prior work loukas2022finer; sharma2023financial that formulates tagging as flat classification over a truncated label set, 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: requires end-to-end reasoning over both narrative disclosures and tables, and scales evaluation to the complete 17k+ concept space.
+0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: adopts a two-stage task formulation that separates decisions typically conflated in extreme classification.
+Financial Numerical Identification (FinNI) asks models to identify numerical facts together with their supporting evidence, while Financial Concept Linking (FinCL) maps each fact to the correct US-GAAP concept by semantic comparison with taxonomy definitions, reframing tagging as alignment rather than label recall.
+We release two expert-verified evaluation sets, FinNI-eval and FinCL-eval, derived from real-world XBRL filings, along with a unified protocol that evaluates both extraction accuracy and concept-alignment performance.
+Across 13 state-of-the-art LLMs, our results show a consistent extraction–linking gap.
+While top models reliably recover numerical facts from mixed table and text inputs, they struggle to assign the correct US-GAAP concepts, indicating that representations sufficient for extraction do not consistently encode the latent semantic conditions that distinguish closely related concepts, such as scope, aggregation level, and qualifying context.
+As a result, FinNI performance is strong, whereas FinCL remains below 0.19 even after candidate retrieval, with most errors arising from near-neighbor concepts that share similar surface cues but differ in these underlying conditions.
+Pipeline controls further clarify this failure mode: accuracy improves when gold entities are provided and improves again when the gold concept is included among candidates, yet it does not converge, showing that errors persist even when evidence is localized and the correct concept is available.
+The restricted cash example illustrates this pattern, where the model selects a narrower concept triggered by a salient local phrase, despite broader contextual evidence supporting an aggregate concept.
+We conclude our main contributions as follows: (1) A New Formulation: We introduce 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first benchmark that reframes XBRL tagging into a two-stage “Extract-and-Align” paradigm, enabling full-scope evaluation against 17k+ US-GAAP concepts, a scale previously intractable for zero-shot LLMs.
+(2) High-Fidelity Resources: We release two expert-verified datasets, FinNI-eval and FinCL-eval, which incorporate complex tabular structures to assess joint reasoning across structured and unstructured financial modalities.
+(3) Diagnostic Insights: By benchmarking 13 SOTA LLMs, we identify a critical “knowledge-alignment gap”, demonstrating that while LLMs excel at numeric extraction, they require our alignment framework to prevent performance collapse during large-scale taxonomy mapping.
+2 Related Work
+From XBRL Classification to Alignment.
+Traditional XBRL tagging has been framed as extreme multi-class classification.
+Benchmarks such as FiNER loukas2022finer and FNXL sharma2023financial evaluate models on truncated label sets (139–2,800 tags), and methods like GalaXC saini2021galaxc, SECBERT loukas2022finer, and AttentionXML sharma2023financial focus on coping with label sparsity through hierarchical modeling.
+However, these closed-set formulations overlook the semantic reasoning required to navigate the full 17k+ US-GAAP taxonomy and are poorly matched to LLM capabilities.
+Recent LLM-based tools such as XBRL-Agent han2024xbrlagent explore filing analysis but do not evaluate end-to-end fact-to-taxonomy alignment.
+0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: addresses this gap by reframing tagging as a two-stage extract-and-align task.
+Financial Document Understanding.
+In parallel, financial NLP has progressed from entity and relation extraction (e.g., FiNER-ORD shah2023finer, FinRED sharma2022finred) to numerical reasoning over documents.
+Benchmarks such as FinQA chen2021finqa, TAT-QA zhu2021tat, and DocVQA mathew2021docvqa require joint reasoning over text and tables.
+Yet these tasks are designed for QA and do not constrain outputs to standardized accounting concepts.
+Our benchmark instead requires both multi-modal fact extraction and precise alignment to official US-GAAP definitions, focusing on the “last mile” of structured financial reporting.
+3 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:
+3.1 Task Formulation
 Formally, given a financial document $D$ consisting of multiple textual contents and structured tables, a set of predefined value data types $L$ specified by the XBRL specification, and a taxonomy database $\mathcal{T}$ containing a set of financial semantic concepts, the XBRL tagging task is to identify all relevant numerical values in $D$ and annotate each with a structured triplet $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$.
 Here, Tag denotes the linked concept in $\mathcal{T}$ (e.g.
-“us-gaap:CashAndDueFrom Banks”, etc.), Fact represents the extracted numerical value as it appears in context (e.g., “2.5”, “10,000”, “two”, etc.), and Type specifies the corresponding data type in $L$ (including monetaryItemType, percentItemType, sharesItemType, perShareItem, and integerItemType).
+“us-gaap:CashAndDueFrom Banks”, etc.), Fact represents the extracted numerical value as it appears in context (e.g., “2.5”, “10,000”, “two”, etc.), and Type specifies the corresponding data type in $L$ (including monetaryItemType, percentItemType, sharesItemType, perShareItemType, and integerItemType).
 Each numerical fact is assigned a single, most specific tag from the taxonomy, without overlapping.
@@
 We formalize the task as a mapping:
-(1) $f:(D,L,\mathcal{T})\mapsto\left\{(\texttt{Fact}_{i},\texttt{Type}_{i},\texttt{Tag}_{i})\right\}_{i=1}^{n}$
+$f:(D,L,\mathcal{T})\mapsto\left\{(\texttt{Fact}_{i},\texttt{Type}_{i},\texttt{Tag}_{i})\right\}_{i=1}^{n}$ (1)
 where each triplet corresponds to a financial value mention extracted from $D$ and semantically grounded in $\mathcal{T}$.
-Inspired by information extraction and alignment works (Wu et al., 2019; Wadden et al., 2019; Mukherjee et al., 2022; Wang et al., 2022) and after discussed with financial reporting specialists, we formulated 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: into two sub-tasks: Financial Numeric Identification (FinNI), a multi-modal numerical information extraction task, detects numerical value in $D$ and classifies each with its appropriate Type.
+Inspired by information extraction and alignment works wu2019scalable; wadden2019entity; mukherjee2022ectsum; wang2022conditional and after discussing with financial reporting specialists, we formulated 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: into two sub-tasks: Financial Numeric Identification (FinNI), a multi-modal numerical information extraction task, detects numerical value in $D$ and classifies each with its appropriate Type.
 Financial Concept Linking (FinCL), a numerical entity normalization task, then associates each identified value with its most appropriate Tag in $\mathcal{T}$ based on contextual and structural cues.
@@
 We formalize this subtask as a mapping:
-(2) $f_{\text{FinNI}}:(D=(S,T),\ L)\mapsto\left\{(e_{i},l_{i})\right\}_{i=1}^{k}$
+$f_{\text{FinNI}}:(D=(S,T),\ L)\mapsto\left\{(e_{i},l_{i})\right\}_{i=1}^{k}$ (2)
 where $S$ and $T$ represent the textual and tabular components of the document, respectively, and $L$ is the set of predefined value data types.
@@
 Formally, we define the mapping as:
-(3) $f_{\text{FinCL}}:(e,l,C_{e},\mathcal{T})\mapsto\hat{c}$
-where $e$ is a numerical entity identified in the document $D=(S,T)$, $l\in L$ is its predicted data type from FinNI, $C_{e}$ denotes the contextual information surrounding $e$ in $D$, and $\mathcal{T}=\{c_{1},c_{2},\dots,c_{n}\}$ is a financial taxonomy containing $n$ uniquely defined and semantically grounded concepts.
-The model is required to assign a concept $\hat{c}\in\mathcal{T}$ that best reflects the meaning of $e$ in its structural context.
-3.2. Raw Data Collection
-We compiled 142 annual 10-K reports filed in 2023 and 2024 by publicly listed companies from the SEC [^6] https://www.sec.gov/, as shown in Table 2.
-To ensure diversity, the collection covers all 11 major industry sectors, with about 95% of companies distributed across 31 U.S.
-states and the remaining 5% based outside the United States.
-Using BeautifulSoup, we parsed these reports and extracted 319,893 narrative sentences (roughly 76 million characters) along with 21,576 financial tables.
-Company-level details are provided in Appendix D.2.
-Each report follows the SEC’s XBRL filing standard, meaning that the underlying instance, schema, and linkbase documents already contain machine-readable tags (ground-truth tags) that link textual and tabular values to US-GAAP taxonomy concepts.
-We included all sections that may reference or embed XBRL-tagged content to ensure comprehensive coverage.
-For the justification of our benchmark granularity, please see the Appendix E.
-Table 2. Statistics of collected financial reports.
+$f_{\text{FinCL}}:(e,l,C_{e},\mathcal{T})\mapsto\hat{c}$ (3)
+where $e$ is a numerical entity extracted from $D=(S,T)$, and $l\in L$ is its associated data type provided by the FinNI stage.
+$C_{e}$ represents the structural context of $e$, which encompasses linguistic dependencies for narrative mentions and hierarchical row-column coordinates for tabular mentions.
+$\mathcal{T}=\{c_{1},c_{2},\dots,c_{n}\}$ denotes the full-scope US-GAAP taxonomy containing $n$ uniquely defined and semantically grounded concepts.
+The objective of FinCL is to identify the optimal concept $\hat{c}\in\mathcal{T}$ that maximizes the semantic alignment between the fact $e$ and its specific context $C_{e}$ within the high-dimensional taxonomy space.
+3.2 Raw Data Collection
+We compiled 142 annual 10-K filings from publicly listed U.S.
+companies submitted in 2023–2024 (Table 2).
+To ensure coverage diversity, the collection spans all 11 major industry sectors, with about 95% of companies located across 31 U.S.
+states and the remainder based internationally.
+Using BeautifulSoup, we parsed the reports and extracted 319,893 narrative sentences and 21,576 financial tables.
+Company-level statistics are provided in Appendix C.1.
+All filings follow the SEC Inline XBRL (iXBRL) format, which embeds machine-readable tags inside HTML disclosures and links textual and tabular values to US-GAAP taxonomy concepts through schema and linkbase definitions.
+However, these regulatory tags are designed for compliance rather than model evaluation.
+They provide little coverage of negative cases and contain substantial structural noise.
+We therefore construct a high-fidelity evaluation suite that repurposes this raw regulatory metadata into a standardized benchmark for financial fact extraction and concept linking.
+Additional discussion of our benchmark granularity is provided in Appendix C.2.
+Table 2: Statistics of collected financial reports.
 Item Value
@@
 #Table 21,576
-3.3. Data Curation
-To construct the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark and ensure both efficiency and fairness in evaluation, we designed a two-stage filtering and annotation pipeline.
-This pipeline systematically processes the raw 10-K filings to retain representative, high-quality instances that collectively cover all 142 companies, five valid entity types, and every numerical US-GAAP concept observed in the corpus.
-The overall workflow is illustrated in Algorithms 1 and 2.
-Stage 1: Filtering Raw Instances (Algorithm 1)
-The first stage aims to remove noisy or redundant samples while ensuring comprehensive coverage of both entity types and US-GAAP concepts.
-Each narrative sentence or table instance is automatically examined by a rule-based parser that checks if it contains numeric entities tagged with valid US-GAAP concepts.
-The parser performs schema-level consistency checks based on the taxonomy definitions to ensure tagging correctness.
-Instances without valid entities are classified as NEG (negative).
-Those containing valid entities are treated as positive and further divided into two categories: (1) POS_SELECT, which introduces at least one new US-GAAP concept not yet covered; and (2) POS_CANDIDATE, which contains only previously covered concepts.
-This design ensures that each concept appears at least once in the dataset while minimizing redundancy across companies and reports.
-Figure 2. Statistics of numerical entity types.
-To determine which entity types should be considered valid, we first profiled all XBRL numeric item types that appeared in the raw corpus and summarized the results in Figure 2.
-For each type, we calculated both its cross-company coverage (the number of companies in which it occurs) and its overall frequency (the proportion of all annotated numeric entities belonging to that type).
-The figure shows a clear long-tailed distribution: a total of eleven numeric item types were identified, but only a few dominate in both coverage and frequency.
-The five types that consistently appear across nearly all companies and collectively account for the majority of entities are monetaryItemType, percentItemType, sharesItemType, perShareItemType, and integerItemType.
-These five types form the label space for all subsequent stages.
-Under this definition, Algorithm 1 treats an instance as a potential positive only if it contains at least one US-GAAP–tagged entity belonging to this label space; otherwise it is assigned to NEG.
-To maintain structural fidelity, additional validity rules are applied.
-For narrative texts, we retain only segments longer than 20 characters that contain numerical tokens.
-For tables, we preserve the <table>, <tr>, <td>, and <th> tags to ensure layout integrity, while tables with headers only (i.e., without numeric values) are excluded.
-Input: 10-K reports; US-GAAP taxonomy; 5 valid entity types
-Output: POS_SELECT, POS_CANDIDATE, NEG
-Initialize: an empty set of covered concepts; empty lists for POS_SELECT, POS_CANDIDATE, and NEG;
-foreach instance (texts or tables) in 10-K reports do
-  Check if the instance contains numeric entities that are both: (i) tagged with a US-GAAP concept, and (ii) of a valid entity type;
- if no such entities are found then
-     Add the instance to NEG;
- else
-     Identify all US-GAAP concepts in this instance;
-     if the instance introduces at least one new concept not yet covered then
-        Add the instance to POS_SELECT and update the covered concepts set;
-    else
-        Add the instance to POS_CANDIDATE;
-return {POS_SELECT, POS_CANDIDATE, NEG};
-Algorithm 1 Screening of Text and Table Instances
-Stage 2: Company-Capped Sampling and Minimal Cover (Algorithm 2)
-The second stage refines the positive pool by enforcing both per-company balance and concept coverage completeness.
-For each modality (text or table), we first perform company-capped seeding: from each company’s POS_SELECT, up to $K=10$ positive instances are retained as the seed set, while the remaining positives are moved into a candidate pool.
-We then mark all US-GAAP concepts covered by the seed and identify those still uncovered.
-Finally, a greedy minimal-cover procedure is executed: candidates are iteratively selected according to the number of uncovered concepts they can cover, with ties broken by shorter text length, fewer entities, or earlier order.
-Each selected candidate is added to the final set $P^{*}$, and the process continues until all concepts are covered or the candidate pool is exhausted.
-The resulting positive set $P^{*}$ thus maintains full US-GAAP concept coverage, balanced company representation, and reduced redundancy, forming a compact yet coverage-complete dataset.
-Input: POS_SELECT grouped by company (run separately for text and table) from algorithm 1;
-   All concepts observed in the POS_SELECT;
-   Cap $K=10$ (maximum positives per company for seeding)
-Output: Final positive instances $P^{*}$ and leftover instances
-Seed selection. For each company, keep up to $K$ positives as the seed set. Put the rest into a new candidate pool.
-Check coverage. Mark all concepts already covered by the seed. Identify which concepts are still uncovered.
-Greedy covering. While uncovered concepts remain:
-  1. (1)
-    For each candidate, count how many uncovered concepts it can cover.
-  2. (2)
-    Pick the candidate with the highest count (ties broken by shorter text, fewer entities, or earlier order).
-  3. (3)
-    Add the candidate to the final positives, update the covered concepts, and remove it from the pool.
-Finalize. The final positives $P^{*}$ are the union of the seed and the greedy additions. The leftover candidates form the candidate set.
-Algorithm 2 Company-Capped Seeding + Minimal Cover (for Text and Table)
-Final Dataset and Validation
-After two-stage filtering, we obtained 3,084 textual sentences and 3,067 table sequences from financial filings (Table 3).
-Among them, 2,054 text and 2,045 table instances are positives containing valid XBRL tagging information, while 1,030 text and 1,022 table instances are negatives.
-Textual inputs contain on average $96.25$ tokens, whereas table sequences are much longer ($1,204.72$ tokens on average).
-Overall, the dataset includes 62,668 annotated entities linked to 3,953 unique US-GAAP concepts across the five valid entity types.
-Based on this foundation, we further derive two subtask-specific datasets: FinNI-eval, targeting numerical entity identification, and FinCL-eval, targeting concept linking.
+3.3 Large-Scale Automated Annotation
+Our automated pipeline produces a large, high-fidelity corpus reflecting real financial disclosures.
+As summarized in Table 3, the dataset contains 15,986 narrative sentences and 12,801 tables, yielding 261,457 numerical entities linked to 3,953 unique US-GAAP concepts.
+Including both valid facts and naturally occurring noise, it supports a realistic evaluation of LLMs under regulatory-style conditions.
+Table 3: Statistical information for the original dataset in our benchmark. Tokens are calculated using the “cl100k_base” tokenizer ($\pm$ standard deviation).
+Item Sentence Table
+Positive instances 7,768 8,709
+Negative instances 8,218 4,092
+Avg. Tokens/S $80.91\pm 63.62$ $1212.42\pm 1421.76$
+Avg. Entities/S $1.24\pm 1.82$ $18.87\pm 37.16$
+Avg. Concepts/S $1.24\pm 1.82$ $18.87\pm 37.16$
+Total Entities 261,457
+Entity Types 5
+Unique Concepts 3,953
+Automated Annotation Engine.
+We build a rule-based parser (Algorithm 1, Appendix C.3) that reconstructs semantic links between human-readable disclosures and Inline XBRL (iXBRL) metadata.
+For each instance $i$, the parser locates ix:nonfraction elements and extracts the value $v$, concept $c$, and data type $\tau$, while restricting extraction to atomic leaf elements $\epsilon$ to avoid duplicates from nested tags.
+Valid Entity Types.
+To define the label space, we profile all numeric XBRL item types.
+As shown in Figure 2, the distribution is long-tailed: although eleven types appear, five dominate.
+We therefore restrict $\mathcal{L}$ to monetaryItemType, percentItemType, sharesItemType, perShareItemType, and integerItemType.
+An instance is labeled POSITIVE ($\mathcal{P}$) only if it contains at least one concept $c\in\mathcal{T}$ from $\mathcal{L}$; otherwise it is labeled NEGATIVE ($\mathcal{N}$).
+Structural Refinement and Filtering.
+For reliable evaluation, we keep only text segments longer than 20 characters and deduplicate content globally using $\mathcal{S}_{text}$.
+For tables, we retain structural tags (<table>, <tr>, <td>, <th>) but discard tables without numerical facts.
+This filtering preserves semantic accuracy while maintaining realistic layout structure.
+Figure 2: Statistics of numerical entity types.
+3.4 Data Quality Control
+Validating all 261,457 entities is infeasible.
+To obtain a rigorous quality upper bound, we adopt a strategic sampling scheme based on a greedy minimal-cover algorithm (Algorithm 2, Appendix D.1).
+Rather than random sampling, we construct a diversity-maximized subset $\mathcal{P}^{*}$ that satisfies two goals: (1) Company Balance, limiting bias from individual reporting styles, and (2) Taxonomy Completeness, ensuring that each US-GAAP concept observed in the corpus appears at least once.
+We seed $\mathcal{P}^{*}$ with capped samples per company, then iteratively add instances that cover the largest number of uncovered concepts, producing a compact but representative audit set.
+The final subset contains 928 positive sentence instances and 1,021 positive table instances, serving as a taxonomy-wide stress test reviewed by two junior analysts and one senior adjudicator (the demography of these annotators is shown in Appendix D.2).
+Review Protocol.
+The audit followed two stages.
+A pilot phase aligned interpretations of the Verification Guideline (Appendix D.3), after which analysts independently verified whether extracted triplets $(e,\tau,c)$ were fully correct.
+The senior expert adjudicated disagreements to create the final gold set.
+Agreement Results.
+On the independent round, annotators achieved 96% raw agreement and Cohen’s $\kappa=0.81$, indicating substantial agreement landis1977measurement, the equation is shown in Appendix D.4.
+This high-verified accuracy on the most semantically diverse subset provides strong evidence for the reliability of our automated annotation pipeline.
+3.5 Task-Specific Dataset Construction
+Following validation, we integrated the corrected $\mathcal{P}^{*}$ subset back into the full corpus to form a high-fidelity hybrid dataset.
+This refined pool ensures that every US-GAAP concept in the taxonomy has at least one expert-verified representation.
+From this foundation, we derive two subtask-specific datasets: FinNI-eval, targeting numerical entity identification, and FinCL-eval, targeting concept linking.
 Both are the first datasets in the financial domain to emphasize structured information extraction and concept-level alignment.
 Their detailed statistics are reported in Table 4.
-To verify annotation reliability, two domain experts independently reviewed 100 randomly sampled entries.
-The results indicate a 96% raw agreement and a Cohen’s Kappa of 81.13, confirming the robustness and reproducibility of our semi-automated annotation pipeline.
-The validation guideline is provided in Appendix F, and a complete unfiltered version of the benchmark is included in Appendix D.1.
-Table 3. Statistical information for the original dataset in our benchmark. Tokens are calculated using the “cl100k_base” tokenizer ($\pm$ standard deviation).
-Item Sentence Table
-Positive instances 2,054 2,045
-Negative instances 1,030 1,022
-Avg. Tokens/S $96.25\pm 72.75$ $1,204.72\pm 1,162.41$
-Avg. Entities/S $1.88\pm 2.11$ $18.54\pm 29.14$
-Avg. Concepts/S $1.88\pm 2.11$ $18.54\pm 29.14$
-Total Entities 62,668
-Entity Types 5
-Unique Concepts 3,953
-Table 4. Statistics of evaluation datasets for FinNI-eval and FinCL-eval. Tokens are calculated using the “cl100k_base” tokenizer ($\pm$ standard deviation).
+Table 4: Statistics of evaluation datasets for FinNI-eval and FinCL-eval. Tokens are calculated using the “cl100k_base” tokenizer ($\pm$ standard deviation).
 Item FinNI-eval FinCL-eval
-#Instance 6,151 62,668
+#Instance 28,787 261,457
 Avg. Input Tokens $1101.50\pm 992.00$ $60.43\pm 48.25$
@@
 Max Output Tokens 7,813 41
-3.4. FinNI-eval Dataset Construction
-Based on the annotation procedure described in Section 3.3, we construct the FinNI-eval dataset, comprising 6,151 instances derived from the annotated sentences and tables.
+3.5.1 FinNI-eval Dataset Construction
+Based on the annotation procedure described in Section 3.3, we construct the FinNI-eval dataset, comprising 28,787 instances derived from the annotated sentences and tables.
 Specifically, the dataset comprises two components: the input block and the answer block.
@@
 The answer is a JSON list of all identified entity values and types or an empty JSON list.
-The specific prompt template is shown in Figure 9 in the Appendix.
+The specific prompt template is shown in Figure 9.
 FinNI-eval Dataset
@@
 Input: <Sentences or Table>
-Answer: {"results": [{"value": <numeric entity>, "type": <entity type>}]}
-        or {"results": []}
-3.5. FinCL-eval Dataset Construction
-Following the formulation of the FinCL subtask in Section 3.1, we further construct a FinCL-eval dataset that includes 62,668 query-answer pairs for numerical entity normalization.
+Answer: {"results": [{"Fact": <numeric entity>,
+        "Type": <entity type>}]} or {"results": []}
+3.5.2 FinCL-eval Dataset Construction
+Following the formulation of the FinCL subtask in Section 3.1, we further construct a FinCL-eval dataset that includes 261,457 query-answer pairs for numerical entity normalization.
 In addition, we built a US-GAAP taxonomy database containing 17,688 unique financial concepts.
@@
 Answer: <US-GAAP Concept>
-3.6. Evaluation
-3.6.1. Evaluation Framework
-We propose a unified evaluation framework for the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark to assess LLM performance.
-As shown in Figure 3, given a financial report $D$, the framework reformulates tagging into two subtasks, FinNI and FinCL, to generate structured triplets $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$.
-This design jointly evaluates an LLM’s zero-shot ability in fact extraction and concept alignment.
-For fair comparison with token-classification baselines, we report macro and micro Precision, Recall, and F1 (Sokolova and Lapalme, 2009).
-FinNI is evaluated using pair-level metrics, while FinCL uses accuracy.
-Full metric definitions are provided in Appendix G.
-Figure 3.
-A unified evaluation framework on 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark.
-Note: The Sentence is drawn from financial report $D$, and the Query provides an example that incorporates the entity 4.9.
-The black arrows show the overall framework flow, green arrows denote the FinNI process, red arrows indicate the FinCL process, red bidirectional arrows represent retrieval interactions with the taxonomy, and red dashed arrows mark the recomposed input between the query and retrieved candidates.
-FinNI Evaluation: We use pair-level metrics to evaluate the extraction performance on FinNI.
-The objective is to evaluate the LLM’s ability to distinguish between structured and unstructured contexts and to extract financial values based on its semantic understanding and domain-specific knowledge.
-FinCL Evaluation: To assess whether LLMs can perform fine-grained XBRL entity normalization, we reformulate the task as a retrieval-reranking problem, where LLMs are used to disambiguate and select the most appropriate taxonomy concept (reranking stage) from a reduced candidate set (retrieval stage), avoiding the impracticality of direct multi-thousand-way classification (prompt template used for reranking in the FinCL subtask is shown in Figure 10 in Appendix).
-To this end, we first generate embeddings for each taxonomy concept using text-embedding-3-small [^7] https://platform.openai.com/docs/models/text-embedding-3-small, and retrieve the top-$h$ candidate tags from the US-GAAP taxonomy $\mathcal{T}$ based on semantic similarity.
-Therefore, the LLMs are required to rerank the retrieved candidates by leveraging deeper contextual understanding to select the best-matched tag $\hat{c}\in\mathcal{T}_{h}$, formalized as:
-(4) $f_{\text{rerank}}:(e,l,C_{e},\mathcal{T}_{h})\mapsto\hat{c}$
-3.6.2. Evaluation Models
-Our goal is to evaluate the foundational capabilities of SOTA LLMs on the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark, to better understand their strengths and limitations in financial tagging.
-To this end, we evaluate models across three categories using our proposed evaluation framework, as detailed in Table 13.
-These include: (1) one general-purpose closed-source LLM: GPT-4o model (Hurst et al., 2024); (2) eleven general-purpose open-source LLMs including DeepSeek-V3 (Liu et al., 2024), DeepSeek-R1-Distill-Qwen-32B (DeepSeek-AI, 2025), Qwen3-32B (Team, 2025), Qwen3-14B (Team, 2025), Qwen3-1.7B (Team, 2025), Qwen3-0.6B (Team, 2025), Llama-4-Scout-17B-16E-Instruct (Meta, 2025), Llama-3.3-70B-Instruct (Grattafiori et al., 2024), Llama-3.1-8B-Instruct (Grattafiori et al., 2024), Llama-3.2-3B-Instruct (Grattafiori et al., 2024), and gemma-2-27b-it (Team, 2024) models; and (3) one domain-specific financial LLM: Fino1-8B (Qian et al., 2025) model.
-In addition, we compare these LLMs with strong PLMs, including BERT-large (Devlin et al., 2018), FinBERT (Araci, 2019), and SECBERT (Loukas et al., 2022).
-3.6.3. Evaluation setting
-We use the LM Evaluation Harness (Gao et al., 2024) to build customized benchmark suites.
-Proprietary models are accessed via APIs, while open-source models are evaluated locally on a 4×H100 GPU cluster (80GB each).
-We standardize the input length to 2,048 tokens for FinNI and 4,096 for FinCL, with a generation limit of 1,024 tokens to support reasoning-intensive outputs.
-For fine-tuning strong PTMs such as BERT-large, FinBERT, and SECBERT, we curate 10 additional annual reports as training data (see Appendix I).
-All data follow the FiNER (Loukas et al., 2022) and FNXL (Sharma et al., 2023) formats.
-For retrieval in FinCL, we use ElasticSearch [^8] https://www.elastic.co/ with text-embedding-3-small as the embedding model.
-For evaluation, we spent a total of 500 GPU hours and about \$1,500 for GPT (from OpenAI) and DeepSeek (from TogetherAI) API.
-4. Experiment and Result
-4.1. Overall Results
+3.6 Evaluation
+3.6.1 Evaluation Framework
+We propose a unified evaluation framework for 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:.
+As illustrated in Figure 3, each input is decomposed into two subtasks, FinNI and FinCL, which together produce structured triplets $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$.
+This design jointly evaluates zero-shot fact extraction and concept alignment.
+For comparison with token-classification baselines, we report macro and micro Precision, Recall, and F1 sokolova2009systematic.
+FinNI is evaluated using pair-level metrics, and FinCL with accuracy.
+Full metric definitions are in Appendix E.
+Figure 3: A unified evaluation framework on 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark.
+Note: The Input content is drawn from the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: dataset, and the “Shared LLMs” are used for identification in the FinNI task and reranking in the FinCL task.
+FinNI Evaluation. We adopt pair-level metrics to measure extraction quality, focusing on whether models can locate financial values from mixed text–table contexts based on semantic cues.
+FinCL Evaluation.
+To test fine-grained numeric entity normalization, we formulate FinCL as a retrieval–reranking task: a retrieval stage produces a candidate set of taxonomy concepts, and the LLM then reranks candidates to select the most appropriate tag, avoiding impractical extreme classification (prompt template in Figure 10, Appendix).
+To this end, we embed each taxonomy concept using text-embedding-3-small [^5] https://platform.openai.com/docs/models/text-embedding-3-small , then retrieve the top-$h$ candidate tags from the US-GAAP taxonomy $\mathcal{T}$ by semantic similarity.
+The LLM subsequently reranks these candidates using contextual reasoning to select the best tag $\hat{c}\in\mathcal{T}_{h}$:
+$f_{\text{rerank}}:(e,l,C_{e},\mathcal{T}_{h})\mapsto\hat{c}$ (4)
+3.6.2 Evaluation Models
+Our goal is to assess the foundational capabilities of SOTA LLMs on 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: and understand their strengths and limitations in XBRL tagging.
+We evaluate models across three categories using our unified framework (Table 11).
+(1) Closed-source general LLM: GPT-4o hurst2024gpt.
+(2) Open-source general LLMs: DeepSeek-V3 liu2024deepseek, DeepSeek-R1-Distill-Qwen-32B deepseekai2025deepseekr1incentivizingreasoningcapability, Qwen3-32B / 14B / 1.7B / 0.6B qwen3technicalreport, Llama-4-Scout-17B-16E-Instruct meta2025llama, Llama-3.3-70B-Instruct, Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct grattafiori2024llama, and gemma-2-27b-it gemma_2024.
+(3) Domain-specific financial LLM: Fino1-8B qian2025fino1.
+We also compare against strong PLM baselines: BERT-large DBLP:journals/corr/abs-1810-04805, FinBERT araci2019finbert, and SECBERT loukas2022finer.
+3.6.3 Evaluation setting
+[…]
+4 Experiment and Result
+4.1 Overall Results
 Table 5 presents the overall performance on the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark.
 It clearly demonstrates that under our framework, LLMs can effectively handle both frequent and rare financial tags, indicating their ability to mitigate long-tail label challenges and underscoring the advantage of our information extraction and alignment formulation over traditional token-level classification approaches.
-Table 5. Overall Performance. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
+Table 5: Overall Performance. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
 Category Models Macro Micro
@@
 SECBERT 0.0231 0.0295 0.0203 0.1870 0.1697 0.1779
-From the macro-level perspective, which emphasizes balanced performance across both frequent and rare tags, DeepSeek-V3, GPT-4o, and Llama-4-Scout-17B-16E-Instruct achieve the top 3 macro-F1 scores (0.0659, 0.0569, and 0.0400), outperforming all fine-tuned PLMs.
-This highlights the strong generalization of large LLMs and the effectiveness of our task design.
-Qwen3-32B also achieves a solid macro-F1 (0.0324), suggesting that good architecture and pretraining can help smaller models perform well in zero-shot settings.
-From the micro-level perspective, which favors frequent labels, DeepSeek-V3 again performs strongly with a micro-F1 of 0.1162, ranking third overall despite no fine-tuning.
-GPT-4o also performs competitively with a score of 0.0863, outperforming most open-source and domain-specific models.
-4.2. Subtask Results
-4.2.1. FinNI subtask
-Table 6 presents the results of different models on the FinNI subtask.
-Overall, larger models demonstrate stronger capabilities in identifying numerical entities and producing structured outputs, even without financial domain adaptation.
-Among them, DeepSeek-V3 achieves the best performance with both the highest recall (0.8430) and F1 (0.6932), while also maintaining competitive precision.
-In comparison, GPT-4o, the closed-source baseline, performs relatively well with balanced scores, securing the second-best F1 (0.5397).
-Interestingly, Qwen3-32B and Qwen3-14B yield very high precision (0.6991 and 0.6912, respectively), but their recall is extremely low, leading to weak F1 scores, which highlights that precision alone is insufficient in this task.
-Table 6.
-Performance comparison of different models on the FinNI subtask.
+From a macro perspective, which emphasizes balanced performance across frequent and rare tags, large general-purpose LLMs clearly dominate.
+DeepSeek-V3, GPT-4o, and Llama-4-Scout achieve the strongest macro-F1 scores, surpassing all fine-tuned PLMs and indicating better generalization to long-tail concepts.
+Qwen3-32B also performs competitively, suggesting that model architecture and pretraining can compensate for a smaller scale.
+From a micro perspective, which reflects performance on frequent labels, the trend is similar: top LLMs remain competitive even without fine-tuning, while fine-tuned PLMs such as SECBERT and BERT-large excel mainly because they benefit from label frequency rather than broader coverage (Table 5).
+4.2 Subtask Results
+4.2.1 FinNI subtask
+Table 6 summarizes results on FinNI.
+Overall, larger LLMs show markedly stronger ability to identify numerical entities and produce structured outputs, even without financial adaptation.
+DeepSeek-V3 achieves the best overall performance, while GPT-4o performs competitively and ranks among the top models.
+Table 6: Performance comparison of different models on the FinNI subtask.
 Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
@@
 Financial LLM Fino1-8B 0.3431 0.1293 0.1878
-Smaller open-source models such as Qwen3-1.7B and Qwen3-0.6B show severe limitations, with F1 below 0.1, underscoring the difficulty of the FinNI subtask in zero-shot settings.
-In contrast, models like gemma-2-27B-it achieve relatively balanced results, ranking third in both recall and F1.
-Finally, the domain-specific Fino1-8B, despite being fine-tuned on financial reasoning QA, underperforms compared to general-purpose LLMs.
-This suggests that domain pretraining on financial text alone provides limited benefit unless the training task aligns closely with FinNI’s requirements.
-4.2.2. FinCL subtask
-Table 7.
-Performance comparison of different models on the FinCL subtask.
+A notable pattern is that some models (e.g., Qwen3 variants) obtain very high precision but extremely low recall, indicating that they extract only a small fraction of target facts.
+This confirms that precision alone is insufficient for FinNI, where missing facts are often more damaging than producing a few spurious ones.
+Smaller open-source models struggle severely, reflecting the difficulty of this zero-shot extraction setting.
+In contrast, mid-size models such as gemma-2-27B achieve more balanced behavior.
+Interestingly, the financial LLM Fino1-8B still lags behind general-purpose models, suggesting that domain text pretraining alone does not help unless the training task aligns closely with FinNI.
+4.2.2 FinCL subtask
+Table 7 reports FinCL performance under the reranking setup, using candidates retrieved from the best top-200 configuration (Table 14).
+Although retrieval provides essential coverage, low Acc@k values, especially for table entities, substantially cap downstream accuracy.
+Table 7: Performance comparison of different models on the FinCL subtask.
 Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
@@
 Financial LLM Fino1-8B 0.0807
-Table 7 reports the accuracy of different models on the FinCL subtask, which is performed on candidate sets obtained from the best top-200 retrieval results, as shown in Table 16 (Appendix J).
-Although the retrieval stage provides a foundation, the low Acc@k values, especially for table-based entities, directly constrain the FinCL performance.
-Within this reranking formulation, DeepSeek-V3 achieves the best accuracy at 0.1889, closely followed by GPT-4o at 0.1829 and Llama-4-Scout-17B-16E-Instruct at 0.1649.
-Other large open-source models perform notably worse, with Llama-3.3-70B at 0.1318.
-The medium-scale models, such as Qwen3-32B (0.1277), DeepSeek-R1-Distill-Qwen-32B (0.1141), gemma-2-27B-it (0.1099), and Qwen3-14B (0.1144), show similar levels.
-While smaller models, including Llama-3.2-3B and Qwen3-0.6B, perform near random with 0.0415 and 0.0414.
-The domain-specific Fino1-8B also lags at 0.0807, suggesting that domain pretraining without explicit task alignment offers little benefit.
-As a result, overall accuracy remains low, reflecting both the inherent difficulty of handling complex taxonomies and subtle financial semantics, and the cascading effect of retrieval errors that further lowers FinCL performance.
-4.3. Ablation analysis
+Within this setting, large LLMs such as DeepSeek-V3, GPT-4o, and Llama-4-Scout perform best, but overall accuracy remains low, reflecting the difficulty of aligning evidence with fine-grained taxonomy concepts.
+Most other models cluster well below this tier, and smaller models approach random behavior.
+The financial model Fino1-8B also lags behind general LLMs, suggesting that domain pretraining alone is insufficient without task-aligned supervision.
+Taken together, FinCL exposes two compounding challenges: taxonomy disambiguation requires reasoning over subtle semantic cues, and retrieval errors propagate directly into reranking, creating a ceiling effect on end-to-end alignment.
+4.3 Ablation analysis
 We further compare our benchmark against extreme multi-class classification settings.
-The prompt template is shown in Figure 11 in the Appendix, and we select the best-performing model from each category (closed-source, open-source, and financial LLMs).
+The prompt template is shown in Figure 11 in the Appendix, and we select the best-performing model from each category.
 The evaluation uses triplet-level ($\{\texttt{Tag},\texttt{Fact},\texttt{Type}\}$) Precision, Recall, and F1.
-Table 8.
-Performance comparison between with/without our evaluation framework on the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark dataset.
+Table 8: Performance comparison between with/without our evaluation framework on the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark dataset.
 Evaluation Mode Model Precision Recall F1
@@
 Fino1-8B 0 0 0
-Table 8 shows that extreme classification is not a valid way to evaluate LLMs on XBRL tagging.
-This single-step formulation requires presenting the full space of thousands of US-GAAP concepts and directly assigning tags to raw text, without first identifying which spans are true numerical entities.
-In practice, this is infeasible: the model cannot realistically process the entire label space and is forced to rely only on its internal knowledge, which leads to meaningless zero-shot predictions.
-As the results indicate, all tested models, including GPT-4o and DeepSeek-V3, collapse to zero precision, recall, and F1 under this setting.
-In contrast, our 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark reformulates the task into a two-stage process that aligns with how tagging is actually performed.
-The FinNI first detects and types coarse-grained numerical entities from text and tables, ensuring that only valid values enter the linking step.
-The FinCL then narrows the taxonomy search space by retrieving a candidate set of relevant tags and requiring the LLM to select the best match using contextual reasoning.
-With this design, models achieve non-trivial performance: for example, DeepSeek-V3 reaches an F1 of 0.1356 and GPT-4o 0.1026.
-These results demonstrate that FinTagging provides a more faithful evaluation of LLMs’ zero-shot tagging capabilities, surfacing meaningful signals where extreme classification fails.
-4.4.
-Error propagation in 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: pipeline
-Table 9 breaks down model accuracy across three evaluation settings: full pipeline, FinCL, and reranking-only in FinCL with guaranteed gold concepts in the top 200 candidates.
-The results reveal clear evidence of cascading errors.
-End-to-end pipeline accuracy is consistently the lowest, since errors from entity extraction (FinNI) and retrieval both propagate into concept linking.
-When gold entities are provided, accuracy improves substantially (e.g., GPT-4o from 0.1052 to 0.1829; DeepSeek-V3 from 0.1074 to 0.1889), confirming that entity identification is a key bottleneck in this context.
-With gold candidates guaranteed in the pool (rerank-only), performance further improves (e.g., GPT-4o 0.2023; DeepSeek-V3 0.2144), showing that narrowing the label space alleviates but does not eliminate difficulty, as semantic ambiguity among fine-grained US-GAAP concepts remains challenging.
-Table 9. Accuracy comparison across the full pipeline, FinCL subtask, and rerank-only settings.
+As shown in Table 8, treating XBRL tagging as single-step extreme classification fails in practice.
+The model must choose directly from thousands of taxonomy concepts without first identifying valid numerical entities, forcing it to rely on internal priors rather than grounded evidence.
+All models collapse to zero, illustrating that this formulation is not a meaningful way to evaluate LLMs on XBRL tagging.
+In contrast, our two-stage framework mirrors how tagging is actually performed.
+FinNI first detects and types candidate numerical facts, and FinCL then retrieves a small set of plausible tags and requires the model to select among them using context.
+Under this formulation, models achieve non-trivial performance, revealing real differences in capability that are completely hidden under extreme classification.
+4.4 Error propagation in 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: pipeline
+Table 9 compares accuracy across three settings: the full pipeline, FinCL with gold entities, and rerank-only with guaranteed gold candidates.
+The pattern reveals clear cascading errors.
+End-to-end performance is consistently the lowest because mistakes from extraction and retrieval accumulate in the linking stage.
+When gold entities are supplied, accuracy improves substantially, showing that entity identification is a major bottleneck.
+Performance increases again in the rerank-only setting, indicating that narrowing the label space reduces difficulty.
+However, accuracy still remains far from solved, suggesting that fine-grained US-GAAP disambiguation is intrinsically challenging even when retrieval is perfect.
+In addition, we provide representative error cases in Appendix I, illustrating how these challenges arise in practice.
+Table 9: Accuracy comparison across the full pipeline, FinCL subtask, and rerank-only settings.
 Model Pipeline FinCL Rerank-only
@@
 Fino1-8B 0.0419 0.0807 0.0908
-4.5. Error cases
-To better illustrate the semantic ambiguity challenge, we present a representative error case from our best-performing model, DeepSeek-V3. Consider the following input paragraph:
-“Cash equivalents include term deposits with banks, money market funds, and all highly liquid investments with original maturities of three months or less…… At December 28, 2024, we had restricted cash of \$31 million recorded in other current assets and restricted cash of \$121 million recorded in other non-current assets.”
-This paragraph contains 2 “monetary entities”: 31 and 121.
-The gold concepts and predicted concepts that are assigned to both entities are shown in Table 10.
-For case 1, the model’s choice was likely influenced by the phrase “recorded in current assets”.
-Yet the paragraph describes the overall composition of restricted cash, and the correct concept reflects this aggregate view rather than a single classification.
-For case 2, although both concepts denote non-current restricted cash, the model selected a narrower variant, showing the difficulty of distinguishing between highly similar concepts with subtle differences in scope.
-Table 10. Error cases from DeepSeek-V3 rerank.
-Case 1: \$31m
-Gold Concept us-gaap:RestrictedCashAndCashEquivalentsAtCarryingValue
-Gold Concept Rank 18 / 200 (not in Top-5)
-Model Prediction us-gaap:RestrictedCashAndCashEquivalentsCurrent
-Case 2: \$121m
-Gold Concept us-gaap:RestrictedCashAndCashEquivalentsNoncurrent
-Gold Concept Rank 2 / 200 (in Top-5)
-Model Prediction us-gaap:CashCashEquivalentsRestrictedCashAndRestrictedCashEquivalents
-5. Conclusion
-This paper presents 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, a benchmark for evaluating large language models on XBRL tagging of real-world financial reports.
-The task is divided into two subtasks, financial numeric identification (FinNI) and concept linking (FinCL), to enable fine-grained evaluation of both information extraction and taxonomy alignment.
-Results show that while LLMs generalize well to long-tail entities and perform competitively in zero-shot settings, they struggle with accurate alignment to GAAP concepts.
-This reveals limitations in structure-aware reasoning and highlights the need for better semantic understanding.
-0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: offers a foundation for advancing research in XBRL tagging and regulatory reporting.
-The limitations of our work and directions for future research are discussed in Appendix A.
-Appendix A Limitations
+5 Conclusion
+This paper introduces 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, a benchmark for evaluating LLMs on XBRL tagging of real financial reports.
+The task is structured into two subtasks, Financial Numerical Identification and Financial Concept Linking, to separately assess extraction and concept alignment.
+Our results show that LLMs perform reasonably well in zero-shot settings and generalize to long-tail entities, but still struggle to align facts with precise US-GAAP concepts.
+These findings point to gaps in structure-aware and semantic reasoning.
+0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: provides a foundation for future work on reliable XBRL tagging and regulatory reporting.
+Limitations
 In this work, we propose 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: and conduct a comprehensive analysis of LLM performance on the task.
 However, several limitations remain.
-First, we have not yet evaluated the latest models, such as GPT-5, gpt-oss, and Qwen3-235B.
+First, we have not yet evaluated the latest models, such as GPT-5.2, gpt-oss, and Qwen3-235B.
 We are currently conducting these experiments and will report the results in future work.
 Second, we collected only a small dataset for fine-tuning PLMs and did not construct task-specific training data for LLMs.
-Future work may explore building LLM-based financial tagging agents through targeted fine-tuning to further enhance performance.
-Appendix B Significance analysis
+Future work may explore building LLM-based XBRL tagging agents through targeted fine-tuning to enhance performance further.
+Potential Risks
+Potential Positive Societal Impacts.
+0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: provides the first structure-aware benchmark for evaluating large language models on realistic XBRL tagging.
+By reframing tagging as a transparent, two-stage Extract–Align process, this benchmark:
+  • Enables more rigorous and reproducible evaluation of model performance on financial fact extraction and taxonomy alignment.
+  • Supports research on safer automation of financial workflows by exposing failure modes that arise in long-tailed concept spaces and complex tabular contexts.
+  • Lowers the barrier for auditing, interpreting, and improving financial AI systems by releasing open datasets, evaluation protocols, and diagnostic analyses.
+By emphasizing open evaluation rather than proprietary pipelines, 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: contributes to more transparent and accountable financial AI research.
+Potential Negative Societal Impacts.
+Despite these benefits, several risks remain:
+  • Misuse and Overreliance.
+High benchmark performance may be mistakenly interpreted as readiness for real-world XBRL tagging, regulatory reporting, or financial decision-making, even though the benchmark does not replace human review or auditing.
+  • Bias and Coverage Limitations.
+The dataset is drawn from U.S.
+SEC filings and US-GAAP concepts.
+As a result, reporting practices, sector distributions, and taxonomy frequency patterns may introduce biases and may not generalize to other jurisdictions or accounting standards.
+  • Benchmark Narrowness.
+0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: evaluates extraction and alignment only.
+Downstream tasks such as aggregation, reconciliation, audit validation, and materiality assessment remain outside the benchmark and may still fail in deployed systems.
+Ongoing iteration, external feedback, and careful documentation are necessary to mitigate these concerns.
+Ethical Considerations and Usage Disclaimer
+The authors assume responsibility for the design and dissemination of 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:.
+All data used in the benchmark are derived from publicly available SEC filings and contain no intentionally included personal or sensitive information.
+Dataset construction followed community ethical guidelines for research involving public financial disclosures.
+Unless otherwise specified, the datasets and accompanying evaluation resources are released for research and educational purposes.
+They should not be used as a substitute for professional accounting judgment, financial reporting systems, or regulatory compliance procedures.
+The benchmark is designed to evaluate model behavior under controlled conditions.
+Performance on 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: does not imply suitability for production deployment, financial auditing, or automated investment or compliance tasks.
+Any real-world use must include human oversight, documentation of limitations, and appropriate legal and regulatory review.
+We used AI-assisted writing tools to help refine the wording and improve the readability of the manuscript.
+All scientific claims, dataset design choices, and experimental analyses were developed and verified by the authors.
+While reasonable care has been taken to ensure accuracy, the authors provide no warranty regarding completeness or suitability for any particular purpose and disclaim liability for any damages arising from use of the benchmark or associated materials.
+Licensing and Terms of Use
+We release all benchmark data, annotations, and accompanying resources under the CC-BY 4.0 license, which permits redistribution and adaptation with attribution.
+All source documents originate from publicly available SEC filings, and we strictly comply with the original terms governing their public distribution.
+The US-GAAP taxonomy is used in accordance with its publicly documented terms.
+Our released benchmark does not contain proprietary material beyond what is already publicly accessible, and any derivative use must continue to acknowledge both the original providers and our benchmark.
+Appendix A Significance analysis
 To further examine the comparative performance across models, we conducted a pairwise significance analysis, with the results summarized in Figure 4.
@@
 (b) Per-model count of significant pairwise differences.
-Figure 4.
-Visualization of significance analysis across models.
+Figure 4: Visualization of significance analysis across models.
 (a) shows the pairwise-bootstrap significance matrix, while (b) summarizes how many significant differences each model has relative to others.
@@
 The pairwise significance matrix, therefore, provides an interpretable view of model robustness and performance distinctiveness beyond average accuracy metrics.
-Appendix C Literature Review
+Appendix B Literature Review
 The XBRL provides a comprehensive taxonomy for financial reporting, encompassing thousands of detailed tags corresponding to concepts within financial statements.
 Applying NER to assign XBRL tags is an emerging yet challenging area.
-C.1. XBRL Tagging benchmark
-FiNER systematically benchmarked several neural architectures on the finer-139 dataset to address numeric-heavy XBRL tagging (Loukas et al., 2022).
+B.1 XBRL Tagging benchmark
+FiNER systematically benchmarked several neural architectures on the finer-139 dataset to address numeric-heavy XBRL tagging loukas2022finer.
 The initial experiments showed that standard BERT underperforms due to subword fragmentation, then the authors introduced pseudo-token strategies replacing numerals with [NUM] or [SHAPE] tokens to stabilize label assignment across fragmented numeric spans.
@@
 Their experiments demonstrated that subword-aware models with numeric-aware pseudo-tokens outperform both word-level BiLSTMs and vanilla BERT, particularly in numeric-heavy contexts, and avoid nonsensical label sequences.
-FNXL extended this benchmarking paradigm to a much larger label space of 2,794 US-GAAP tags, reframing the task as an extreme classification problem (Sharma et al., 2023).
+FNXL extended this benchmarking paradigm to a much larger label space of 2,794 US-GAAP tags, reframing the task as an extreme classification problem sharma2023financial.
 They compared the FiNER sequence labeling approach with a two-step pipeline that first identifies numeric spans and then assigns labels using AttentionXML.
@@
 Together, these benchmarks reveal the need for context-aware reasoning and label-ranking mechanisms in realistic XBRL tagging scenarios.
-C.2. XBRL Tagging Methods
-The previous studies also explored the approaches address scalability, semantic ambiguity, and reasoning gaps in XBRL tagging to improve the performance.
+B.2 XBRL Tagging Methods
+The previous studies also explored approaches to address scalability, semantic ambiguity, and reasoning gaps in XBRL tagging to improve performance.
 Saini et al.
-(Saini et al., 2021) proposed GalaXC is a graph-based extreme classification framework that jointly learns over document-label graphs with per-label attention across multi-hop neighborhoods.
+saini2021galaxc proposed GalaXC is a graph-based extreme classification framework that jointly learns over document-label graphs with per-label attention across multi-hop neighborhoods.
 By integrating label metadata and transitive label correlations, GalaXC outperformed leading deep classifiers by up to 18% in micro-F1 on standard benchmarks and achieved 25% gains in warm-start scenarios where partial labels are available.
 Moreover, Wang et al.
-(Wang, 2023) addressed the practical challenge of custom tag standardization through a semantic similarity pipeline that leverages TF-IDF, Word2Vec, and FinBERT embeddings.
-Although unsupervised, the method was tested across nearly 200,000 custom tags from SEC filings between 2009 and 2022, and showed strong alignment performance, with vector-based mappings identifying viable standard tag candidates for a substantial proportion of non-compliant elements—offering a low-cost, interpretable solution for downstream financial analysis.
-Shifting focus from classification to comprehension, XBRL-Agent evaluated the capabilities of large language models to reason over full XBRL reports (Han et al., 2024).
-The authors introduced two task types—domain taxonomy understanding and numeric reasoning and found that base LLMs often hallucinated or misinterpreted financial content.
+wang2023standardizing addressed the practical challenge of custom tag standardization through a semantic similarity pipeline that leverages TF-IDF, Word2Vec, and FinBERT embeddings.
+Although unsupervised, the method was tested across nearly 200,000 custom tags from SEC filings between 2009 and 2022, and showed strong alignment performance, with vector-based mappings identifying viable standard tag candidates for a substantial proportion of non-compliant elements, offering a low-cost and interpretable solution for downstream financial analysis.
+Shifting focus from classification to comprehension, XBRL-Agent evaluated the capabilities of large language models to reason over full XBRL reports han2024xbrlagent.
+The authors introduced two task types, domain taxonomy understanding and numeric reasoning, and found that base LLMs often hallucinated or misinterpreted financial content.
 To overcome these issues, XBRL-Agent incorporated retrieval-augmented generation (RAG) and symbolic calculators within an LLM-agent framework.
@@
 Collectively, these works broaden the methodological landscape of XBRL tagging from graph-based label propagation and embedding-based normalization to LLM-driven report analysis and point to a hybrid future where structured priors and reasoning tools jointly support accurate, scalable financial information extraction.
-C.3. Financial Evaluation Benchmarks
+B.3 Financial Evaluation Benchmarks
 In parallel to XBRL-specific advances, the financial NLP community has developed comprehensive benchmarks to assess broader capabilities in information extraction, numerical reasoning, and document understanding.
-FiNER-ORD (Shah et al., 2023) introduced a high-quality, domain-specific NER dataset annotated over financial news, emphasizing general entity types like persons, organizations, and locations.
+FiNER-ORD shah2023finer introduced a high-quality, domain-specific NER dataset annotated over financial news, emphasizing general entity types like persons, organizations, and locations.
 While not numerically focused like FiNER-139, it highlights the lexical diversity of financial discourse and establishes a strong baseline for testing pretrained and zero-shot LLMs in real-world financial NER scenarios.
-FinQA (Chen et al., 2021) pushed toward explainable QA by pairing expert-written questions with annotated multi-step reasoning programs derived from earnings reports.
-ConvFinQA (Chen et al., 2022) extended this challenge to conversational contexts, simulating real-world question flows over sequential financial queries.
-TAT-QA (Zhu et al., 2021) focused on hybrid tabular-text reasoning and required models to align cell values and document narratives, often involving aggregation, comparison, and unit-scale interpretation.
-Pixiu (Xie et al., 2023) introduced a broader evaluation framework by releasing FinMA, a financial LLM instruction-tuned across five tasks, and assessing it on a new benchmark covering sentiment classification, QA, summarization, NER, and stock prediction.
-BizBench (Koncel-Kedziorski et al., 2023) framed financial QA as program synthesis over realistic, multi-modal contexts, integrating reasoning, code generation, and domain knowledge into a single evaluation pyramid.
+FinQA chen2021finqa pushed toward explainable QA by pairing expert-written questions with annotated multi-step reasoning programs derived from earnings reports.
+ConvFinQA chen2022convfinqa extended this challenge to conversational contexts, simulating real-world question flows over sequential financial queries.
+TAT-QA zhu2021tat focused on hybrid tabular-text reasoning and required models to align cell values and document narratives, often involving aggregation, comparison, and unit-scale interpretation.
+Pixiu xie2023pixiu introduced a broader evaluation framework by releasing FinMA, a financial LLM instruction-tuned across five tasks, and assessing it on a new benchmark covering sentiment classification, QA, summarization, NER, and stock prediction.
+BizBench koncel2023bizbench framed financial QA as program synthesis over realistic, multi-modal contexts, integrating reasoning, code generation, and domain knowledge into a single evaluation pyramid.
 While these benchmarks highlight the growing ability of models to integrate structured and unstructured financial data, they overlook taxonomy-driven fact alignment and do not support the structured output formats required for XBRL tagging.
-Appendix D Data and Ticker Information
-D.1. Overall Data Annotation
-Different from the filtered dataset described in Section 3.3, we construct an overall annotation dataset based directly on the unfiltered raw data.
-As summarized in Table 11, the dataset consists of both sentence-level and table-level instances drawn from the narrative and tabular content of financial reports.
-In total, we annotated 15,986 sentence instances and 12,801 table instances, including 7,768 positive and 8,218 negative sentences, as well as 8,709 positive and 4,092 negative tables.
-On average, textual sentences contain 80.91 tokens, while table sequences are considerably longer, averaging 1,212.42 tokens.
-Across all data, we identified 261,457 numerical entities linked to 3,953 unique US-GAAP concepts.
-Table 11. Statistical information for the original dataset in our benchmark. Tokens are calculated using the “cl100k_base” tokenizer ($\pm$ standard deviation).
-Item Sentence Table
-Positive instances 7,768 8,709
-Negative instances 8,218 4,092
-Avg. Tokens/S $80.91\pm 63.62$ $1212.42\pm 1421.76$
-Avg. Entities/S $1.24\pm 1.82$ $18.87\pm 37.16$
-Avg. Concepts/S $1.24\pm 1.82$ $18.87\pm 37.16$
-Total Entities 261,457
-Unique Concepts 3,953
-D.2. The Statistics of the Tickers
+Appendix C Data and Ticker Information
+C.1 The Statistics of the Tickers
 Figure 5, Figure 6, and Figure 7 illustrate the distribution of 142 tickers across industry sectors, market capitalization categories, and geographic regions, highlighting the diversity of our collected financial reports.
 Considering that practical XBRL tagging practices often vary by industry, company size, and legal jurisdiction, we curated a diverse set of companies covering all 11 major industry sectors.
-The sample maintains a balanced distribution across firm sizes, following market capitalization categories: micro-cap (¡\$300M), small-cap (\$300M–\$2B), mid-cap (\$2B–\$10B), large-cap (\$10B–\$200B), and mega-cap (¿\$200B).
+The sample maintains a balanced distribution across firm sizes, following market capitalization categories: micro-cap (<\$300M), small-cap (\$300M–\$2B), mid-cap (\$2B–\$10B), large-cap (\$10B–\$200B), and mega-cap (>\$200B).
 In addition, we incorporated firms from over 30 states and international jurisdictions to capture regional differences in reporting and tagging conventions.
 This diversity ensures that the benchmark reflects realistic heterogeneity observed in financial disclosures across sectors, scales, and regulatory environments.
-Figure 5. Distribution of Tickers by Industry Sector.
-Figure 6. Distribution of Tickers by Market Cap Size.
-Figure 7. Distribution of Tickers by Geography.
-Appendix E Benchmark Granularity Justification
+Figure 5: Distribution of Tickers by Industry Sector.
+Figure 6: Distribution of Tickers by Market Cap Size.
+Figure 7: Distribution of Tickers by Geography.
+C.2 Data Granularity Justification
 Each 10-K report follows the SEC’s XBRL filing standard, where all narrative and tabular disclosures are linked through machine-readable tags defined in the US-GAAP taxonomy.
@@
 Overall, this segmentation strategy allows our benchmark to remain consistent with how XBRL tagging operates in practice (on localized factual statements), while maintaining scalability and interpretability across large collections of filings.
-Appendix F Validation Guideline
-F.1. Task Definition
+C.3 Automated Data Annotation
+Algorithm 1 converts Inline XBRL (iXBRL) filings into structured instances for our benchmark.
+The procedure operates over both narrative sentences and tables, and assigns each instance a positive or negative label based on whether it contains at least one validated financial entity.
+We begin by iterating through all disclosure instances in the corpus and extracting their plain-text content.
+Very short fragments (length $\leq 20$) and duplicated passages are discarded using a global text set, ensuring that the benchmark contains unique and meaningful contexts.
+For each remaining instance, the algorithm traverses all atomic iXBRL elements.
+From each element, we retrieve the numerical value $v$, its associated concept $c$ from the US-GAAP taxonomy, and the declared data type $\tau$.
+We retain only elements whose concept is in $\mathcal{T}$ and whose type belongs to the predefined valid numeric set $\mathcal{L}$.
+Restricting extraction to leaf-level elements avoids double counting values that appear in nested XML structures.
+A local deduplication set further prevents repeated annotation of identical $(v,c)$ pairs within the same instance.
+If at least one valid entity is identified, the instance is recorded as a positive example together with its entity list; otherwise, it is stored as a negative example.
+Negative instances are intentionally preserved, as they reflect realistic disclosure text that models must learn to ignore.
+Overall, this process yields a corpus that is faithful to regulatory metadata while remaining structurally clean and suitable for downstream evaluation.
+Input: Raw 10-K reports (iXBRL) $\mathcal{D}$; US-GAAP taxonomy $\mathcal{T}$; Set of valid data types $\mathcal{L}$
+Output: Positive instances $\mathcal{P}$; Negative instances $\mathcal{N}$
+Initialize: $\mathcal{P}\leftarrow\emptyset$, $\mathcal{N}\leftarrow\emptyset$, Global unique text set $\mathcal{S}_{text}\leftarrow\emptyset$;
+foreach instance (sentence or table) $i\in\mathcal{D}$ do
+ $text\leftarrow$ Extract plain text content from instance $i$;
+ if length of $text\leq 20$ or $text\in\mathcal{S}_{text}$ then
+    skip to the next instance;
+ $E\leftarrow\emptyset$ // List of validated entities in $i$;
+ $S_{fact}\leftarrow\emptyset$ // Local fact deduplication set;
+ foreach atomic iXBRL element $\epsilon$ within instance $i$ do
+      Retrieve value $v$, financial concept $c$, and data type $\tau$ from $\epsilon$;
+    if $c\in\mathcal{T}$ and $\tau\in\mathcal{L}$ and $v$ is not empty then
+       if $(v,c)\notin S_{fact}$ then
+            Add $(v,\tau,c)$ to the entity list $E$;
+            Add $(v,c)$ to the local deduplication set $S_{fact}$;
+ if $E$ is not empty then
+      Add $(text,E)$ to the positive set $\mathcal{P}$;
+ else
+      Add $(text,\emptyset)$ to the negative set $\mathcal{N}$;
+  Add texttext to the global set $\mathcal{S}_{text}$;
+return $\mathcal{P},\mathcal{N}$;
+Algorithm 1 Automated Instance Annotation
+Appendix D Annotation Data Quality Control
+D.1 Validation Data Filtering
+To establish a reliable upper bound on annotation quality, we construct a strategic audit subset rather than validating the full corpus.
+Algorithm 2 formalizes the sampling process as a greedy minimal-cover problem over observed US-GAAP concepts.
+The algorithm takes as input the set of positive instances $\mathcal{P}$, together with the set of all taxonomy concepts observed in the corpus, $\mathcal{C}_{all}$.
+The goal is to produce a compact subset $\mathcal{P}^{*}$ that is both diverse and concept-complete.
+Seeding for company diversity.
+We first iterate over companies and select up to $K$ instances per company.
+This prevents the audit set from being dominated by firms with longer filings or more frequent disclosures.
+After each selection, we update the set of uncovered concepts $\mathcal{C}_{rem}$ by removing concepts already represented among the seeds.
+Greedy minimal cover.
+We then repeatedly add instances that contribute maximal marginal concept coverage.
+For each candidate instance, we compute the number of uncovered concepts it contains and select the instance with the largest score.
+When ties occur, we prefer shorter instances to reduce annotation effort while retaining conceptual value.
+After adding each instance, we update $\mathcal{C}_{rem}$ accordingly and continue until either all concepts are covered or no additional candidates remain.
+Resulting audit subset.
+This procedure yields a compact subset that balances two competing objectives: coverage of the observed taxonomy and representation across companies.
+The resulting $\mathcal{P}^{*}$ therefore serves as a stress-test sample that is both manageable for human review and representative of the broader benchmark space.
+Input: Positive instances $\mathcal{P}$ from Algorithm 1; Observed taxonomy concepts $\mathcal{C}_{all}$; Company cap $K=10$
+Output: Strategic audit subset $\mathcal{P}^{*}$
+Initialize: $\mathcal{P}^{*}\leftarrow\emptyset$;
+Initialize: $\mathcal{C}_{rem}\leftarrow\mathcal{C}_{all}$ // Set of concepts yet to be covered;
+Step 1: Diversity-Preserving Seeding
+foreach unique company in $\mathcal{P}$ do
+   Select up to $K$ instances belonging to this company and add to $\mathcal{P}^{*}$;
+   Update $\mathcal{C}_{rem}$ by removing concepts covered by these seed instances;
+Step 2: Greedy Minimal Cover
+while $\mathcal{C}_{rem}$ is not empty and candidates remain in $\mathcal{P}$ do
+ foreach candidate instance $i\in\mathcal{P}\setminus\mathcal{P}^{*}$ do
+      Calculate $Score(i)=|\text{Concepts in }i\cap\mathcal{C}_{rem}|$;
+  Select instance $i^{*}$ with the highest $Score(i)$;
+ if multiple instances have the same $Score(i^{*})$ then
+      Break ties by selecting the instance with the shortest text length;
+  Add $i^{*}$ to $\mathcal{P}^{*}$ and update $\mathcal{C}_{rem}$ by removing newly covered concepts;
+return $\mathcal{P}^{*}$;
+Algorithm 2 Strategic Audit Set Sampling
+D.2 Annotator Demography
+To ensure reliable verification of financial entities and taxonomy mappings, all audits were conducted by annotators with prior exposure to XBRL-based reporting and financial statement analysis.
+The review team consisted of two junior analysts and one senior adjudicator.
+The two junior analysts are graduate-level researchers with training in accounting and financial data analytics.
+Both have received formal instruction on US-GAAP concepts, XBRL tagging conventions, and common disclosure practices in SEC filings.
+They have prior experience working with financial datasets in academic or industry projects, including tasks involving concept normalization, document review, and regulatory text interpretation.
+Before participating in the audit, each analyst completed a guided calibration session using our Verification Guideline to ensure consistent treatment of numerical entities, metadata attributes, and taxonomy semantics.
+The senior adjudicator has over a decade of experience in financial reporting analysis and structured disclosure systems.
+Their professional background includes work with XBRL instances, taxonomy structure, and validation procedures used in regulatory environments.
+In addition to overseeing the audit workflow, the senior adjudicator resolved disagreements, provided clarification on ambiguous tagging scenarios, and ensured alignment with US-GAAP definitions and XBRL best practices.
+Together, this composition provides both annotation robustness and domain oversight: junior analysts contribute detailed review capacity, while the senior adjudicator ensures conceptual correctness and methodological consistency across the audit.
+All annotators were members of the research team and were invited to participate as part of their normal research training and collaborative project responsibilities.
+Participation was voluntary, and no separate monetary compensation was provided.
+Because the work involved publicly available financial disclosures and did not require annotators to process personal or sensitive data, the review posed minimal risk to participants.
+D.3 Validation Guideline
+D.3.1 Task Definition
 Given a context and an extracted triplet consisting of:
@@
   • 0 for incorrect.
-F.2. Validation Procedure
+D.3.2 Validation Procedure
 Follow the rules below, the definitions of entity type are shown in Table 10:
-  1. (1)
+  1. 1.
     Entity Check Determine whether the extracted entity is a numerical value.
@@
     • If no, set is_correct = 0.
-  2. (2)
+  2. 2.
     Entity Type Validation Verify whether the identified entity type is correct based on the context and the definitions in Table 10.
@@
     • If no, set is_correct = 0.
-  3. (3)
+  3. 3.
     US-GAAP Concept Validation Assess whether the assigned US-GAAP concept is appropriate based on the taxonomy tag definitions.
@@
     • If no, set is_correct = 0.
-Table 12. Entity type definitions used for validation.
+Table 10: Entity type definitions used for validation.
 Entity Type Definition
@@
 percentItemType Ratios or percentages, such as tax rates, growth rates, or discount rates, usually expressed with a percentage symbol (%).
-F.3. Important Instructions
-  1. (1)
+D.3.3 Important Instructions
+  1. 1.
     Non-Arabic Formats Financial numerical entities may appear in word form (e.g., ten million) and must be correctly identified and converted into standard numerical format.
-  2. (2)
+  2. 2.
     Magnitude Terms If a number is followed by a magnitude term (e.g., hundred, million, billion), do not expand it into the full numerical value:
@@
     • 10.6 million $\rightarrow$ extract only 10.6, not 10,600,000.
-  3. (3)
+  3. 3.
     Standardization of Format Remove formatting symbols while preserving the numerical value:
@@
     • Remove commas (e.g., 1,000 $\rightarrow$ 1000).
-Appendix G Evaluation Metrics
-To provide a fair evaluation of overall benchmark performance, we adopt a set of metrics, focusing primarily on macro-level and micro-level evaluation strategies inspired by the previous work (Sharma et al., 2023).
+D.4 Agreement Computation
+We report both raw agreement and Cohen’s $\kappa$ to quantify inter-annotator reliability.
+Let $N$ denote the total number of instances, and let $n_{yy}$ be the number of cases on which both annotators assign the same label $y$. The raw agreement is
+$P_{o}=\frac{1}{N}\sum_{y}n_{yy}.$ (5)
+To adjust for agreement occurring by chance, we compute Cohen’s $\kappa$.
+Let $p_{1y}$ and $p_{2y}$ denote the marginal probabilities that annotator 1 and annotator 2 assign label $y$, respectively.
+The expected chance agreement is
+$P_{e}=\sum_{y}p_{1y}\,p_{2y},$ (6)
+and Cohen’s $\kappa$ is
+$\kappa=\frac{P_{o}-P_{e}}{1-P_{e}}.$ (7)
+Following landis1977measurement, our $\kappa=0.81$ indicates substantial agreement.
+Appendix E Evaluation Metrics
+To provide a fair evaluation of overall benchmark performance, we adopt a set of metrics, focusing primarily on macro-level and micro-level evaluation strategies inspired by the previous work sharma2023financial.
 Macro-level evaluation computes precision, recall, and F1 scores independently for each BIO-concept label derived from the US-GAAP taxonomy, and then averages them without weighting.
@@
 We evaluate the performance based on the following metrics:
-(5) $\text{Precision}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{P}|}$
-(6) $\text{Recall}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{G}|}$
-(7) $\text{F1}=\frac{2\times\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$
+$\text{Precision}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{P}|}$ (8)
+$\text{Recall}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{G}|}$ (9)
+$\text{F1}=\frac{2\times\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (10)
 where $(e_{i},l_{i})=(e^{\prime}_{j},l^{\prime}_{j})$ if and only if both the entity span $e$ and its assigned type $l$ exactly match.
 For the FinCL subtask, Given a set of queries $\mathcal{Q}=\{q_{1},q_{2},\dots,q_{N}\}$, where each $q_{i}$ is associated with a ground-truth concept $c_{i}^{*}$ and a predicted concept $\hat{c}_{i}$, the accuracy is defined as:
-(8) $\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\delta(\hat{c}_{i},c_{i}^{*})$
+$\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\delta(\hat{c}_{i},c_{i}^{*})$ (11)
 where $\delta(\hat{c}_{i},c_{i}^{*})=1$ if $\hat{c}_{i}=c_{i}^{*}$, and 0 otherwise.
-Appendix H Evaluation Models Details
+Appendix F Evaluation Models Details
 Table 11 provides an overview of the models evaluated in this study, categorized by openness, domain specialization, and architectural foundation. The evaluation covers a diverse range of models:
-  • Closed-source LLMs: We include GPT-4o (Hurst et al., 2024), accessed via OpenAI’s API, as a representative of cutting-edge proprietary models with demonstrated performance across a variety of NLP tasks.
+  • Closed-source LLMs: We include GPT-4o hurst2024gpt, accessed via OpenAI’s API, as a representative of cutting-edge proprietary models with demonstrated performance across a variety of NLP tasks.
 Although model size details are undisclosed, GPT-4o serves as an upper-bound reference in our benchmark.
-  • Open-source LLMs: This group encompasses recent, high-performing open models such as DeepSeek-V3 (685B) (Liu et al., 2024), DeepSeek-R1-Distill-Qwen (32B) (DeepSeek-AI, 2025), and multiple variants of Qwen3  (Team, 2025)(ranging from 0.6B to 32B).
-We also include Llama-3.3, Llama-3.2, and 3.1 variants  (Grattafiori et al., 2024) (70B, 3B, and 8B), Llama-4-Scout-17B-16E-Instruct (Meta, 2025)(109B), as well as Google’s Gemma-2-27B  (Team, 2024), to ensure architectural diversity and scalability comparison.
+  • Open-source LLMs: This group encompasses recent, high-performing open models such as DeepSeek-V3 (685B) liu2024deepseek, DeepSeek-R1-Distill-Qwen (32B) deepseekai2025deepseekr1incentivizingreasoningcapability, and multiple variants of Qwen3  qwen3technicalreport(ranging from 0.6B to 32B).
+We also include Llama-3.3, Llama-3.2, and 3.1 variants  grattafiori2024llama (70B, 3B, and 8B), Llama-4-Scout-17B-16E-Instruct meta2025llama(109B), as well as Google’s gemma-2-27B  gemma_2024, to ensure architectural diversity and scalability comparison.
 These models are primarily instruction-tuned and optimized for general-purpose NLP tasks.
-  • Financial-specific LLMs: We evaluate Fino1-8B  (Qian et al., 2025), a domain-specialized model trained on financial corpora, designed to better capture the terminology and structure unique to financial disclosures.
+  • Financial-specific LLMs: We evaluate Fino1-8B  qian2025fino1, a domain-specialized model trained on financial corpora, designed to better capture the terminology and structure unique to financial disclosures.
 This category allows us to assess the benefits of domain adaptation in complex tagging and reasoning tasks.
-  • Pretrained Language Models (PLMs): To establish strong baselines, we include non-generative encoder models: BERT-large  (Devlin et al., 2018), FinBERT  (Araci, 2019), and SECBERT  (Loukas et al., 2022).
+  • Pretrained Language Models (PLMs): To establish strong baselines, we include non-generative encoder models: BERT-large  DBLP:journals/corr/abs-1810-04805, FinBERT  araci2019finbert, and SECBERT  loukas2022finer.
 These models have been widely used in prior financial NLP tasks and allow for a comparative analysis between generative LLMs and traditional pretrained models in terms of domain understanding and structured output capability.
 Together, these models offer a comprehensive evaluation spectrum, from general-purpose to domain-specific, encoder-based to decoder-based, and open to closed source, facilitating an in-depth assessment of their performance across our proposed benchmark tasks.
-Table 13. Model categories and corresponding repositories.
+Table 11: Model categories and corresponding repositories.
 Model Size Source
@@
 SECBERT $\sim$110M nlpaueb/sec-bert-base
-Appendix I The details for the fine-tuning PTMs
-I.1. Training data collection and processing
+Appendix G The details for the fine-tuning PTMs
+G.1 Training data collection and processing
 Similar to the collection process for the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark data, we gathered an additional 10 annual 10-K financial reports filed with the SEC for the period from February 13, 2024, to February 13, 2025, as summarized in Table 12.
@@
 The companies included in this dataset follow the XBRL standard, ensuring comprehensive coverage for training PTMs.
-Table 14. Financial report statistics summary for raw training data
+Table 12: Financial report statistics summary for raw training data
 Item Information
@@
 Overall, the annotated dataset includes 25,199 entities, covering 1,435 unique US-GAAP concepts.
-Table 15. Statistics of training data (tokens calculated with “cl100k_base” tokenizer, $\pm$ standard deviation).
+Table 13: Statistics of training data (tokens calculated with “cl100k_base” tokenizer, $\pm$ standard deviation).
 Structure Pos/Neg #Instance Avg. Tokens/S Avg. Entities/S Avg. Concepts/S Total Entities Unique Concepts
@@
 Specifically, B denotes the beginning of an entity phrase, I marks the continuation (inside) of an entity phrase, and O indicates tokens outside of any entity.
-As shown in Figure 8, ”4.9” and ”4.5” are single-token numerical entities labeled only with a B prefix (e.g., “B-us-gaap:AccountsReceivableNetNoncurrent”).
-To comprehensively cover all US-GAAP tags, we combine the entire set of 17,388 tags from the US-GAAP 2024 taxonomy with the BIO labeling scheme to construct an extreme classification label space, resulting in 34,777 unique entity labels ($2\times 17388+1$).
-Figure 8. An example for training set annotation with BIO scheme.
+As shown in Figure 8, “4.9” and “4.5” are single-token numerical entities labeled only with a B prefix (e.g., “B-us-gaap:AccountsReceivableNetNoncurrent”).
+To comprehensively cover all US-GAAP tags, we combine the entire set of 17,688 tags from the US-GAAP 2023 and 2024 taxonomy with the BIO labeling scheme to construct an extreme classification label space, resulting in 34,777 unique entity labels ($2\times 17388+1$).
+Figure 8: An example for training set annotation with BIO scheme.
 After constructing the training set, we reconstruct the testing set from the original benchmark dataset. The training settings are detailed below.
-I.2. Training settings
-We fine-tune three pretrained models, BERT-large (Devlin et al., 2018), FinBERT (Araci, 2019), and SECBERT (Loukas et al., 2022), on our training set using the HuggingFace Transformers library.
+G.2 Training settings
+We fine-tune three pretrained models, BERT-large DBLP:journals/corr/abs-1810-04805, FinBERT araci2019finbert, and SECBERT loukas2022finer, on our training set using the HuggingFace Transformers library.
 All models are trained with a batch size of 4, a learning rate of 3e-5, and for 20 epochs.
@@
 All experiments are run under a fixed random seed to ensure reproducibility.
-Appendix J Retrieval Results in FinCL subtask
+Appendix H Retrieval Results in FinCL subtask
 We investigate the impact of different context construction strategies for the queried entity at the retrieval stage.
@@
 In contrast, the SAC strategy builds context based on the structural location of the entity: if the entity appears in a sentence, the entire sentence is used; if the entity appears in a table, we linearize the entire row into a Markdown-style key-value format to serve as the context.
-Table 16. Acc@k retrieval performance on the FinCL task. FWC: fixed window-based; SAC: structure-aware.
+Table 14: Acc@k retrieval performance on the FinCL task. FWC: fixed window-based; SAC: structure-aware.
 Strategy Structure Acc@1 Acc@10 Acc@20 Acc@50 Acc@100 Acc@150 Acc@200
@@
 These findings underscore the importance of structure-aware context construction, especially in scenarios where inputs span multiple formats such as sentences and tables.
+Appendix I Error cases
+To better illustrate the semantic ambiguity challenge, we present a representative error case from our best-performing model, DeepSeek-V3. Consider the following input paragraph:
+“Cash equivalents include term deposits with banks, money market funds, and all highly liquid investments with original maturities of three months or less…… At December 28, 2024, we had restricted cash of \$31 million recorded in other current assets and restricted cash of \$121 million recorded in other non-current assets.”
+This paragraph contains 2 “monetary entities”: 31 and 121.
+The gold concepts and predicted concepts that are assigned to both entities are shown in Table 15.
+For case 1, the model’s choice was likely influenced by the phrase “recorded in current assets”.
+Yet the paragraph describes the overall composition of restricted cash, and the correct concept reflects this aggregate view rather than a single classification.
+For case 2, although both concepts denote non-current restricted cash, the model selected a narrower variant, showing the difficulty of distinguishing between highly similar concepts with subtle differences in scope.
+Table 15: Error cases from DeepSeek-V3 rerank.
+Case 1: \$31m
+Gold Concept us-gaap:RestrictedCashAndCashEquivalentsAtCarryingValue
+Gold Concept Rank 18 / 200 (not in Top-5)
+Model Prediction us-gaap:RestrictedCashAndCashEquivalentsCurrent
+Case 2: \$121m
+Gold Concept us-gaap:RestrictedCashAndCashEquivalentsNoncurrent
+Gold Concept Rank 2 / 200 (in Top-5)
+Model Prediction us-gaap:CashCashEquivalentsRestrictedCashAndRestrictedCashEquivalents
 Prompt Template for FinNI Subtask You are a financial information extraction expert specializing in identifying financial numerical entities in XBRL reports.
@@
 (5) Output the extracted financial entities in JSON list format without explanations, structured as follows: {"result":[{"Fact": <Extracted Numerical Entity>, "Type": <Identified Entity Type>}]} Input: {text/table} Output:
-Figure 9. Prompt template used for the FinNI subtask.
+Figure 9: Prompt template used for the FinNI subtask.
 Prompt Template for FinCL Subtask (Reranking) You are a financial tagging assistant trained in US-GAAP taxonomy.
@@
 Input Query: <entity> + <entity type> + <context> Candidate Tags: {Top 200 US-GAAP tags} Answer:
-Figure 10. Prompt template used for the Reranking statge in the FinCL subtask.
+Figure 10: Prompt template used for the Reranking stage in the FinCL subtask.
 Prompt Template for Ablation You are an XBRL tagging expert specializing in annotating financial numerical facts in XBRL reports.
@@
 (6) Output the extracted financial entities in JSON list format without explanations, structured as follows: {"result":[{"Fact": <Extracted Numerical Entity>, "Type": <Identified Entity Type>, "Tag": <Assigned US-GAAP tag>}]} Input: {text/table} Output:
-Figure 11. Prompt template used for ablation study.
+Figure 11: Prompt template used for ablation study.

### v4

@@
 
-Abstract
+Abstract.
 Accurate interpretation of numerical data in financial reports is critical for markets and regulators.
@@
 Code is available at GitHub  [^1] https://github.com/The-FinAI/FinTagging and datasets are available at Hugging Face  [^2] https://huggingface.co/collections/TheFinAI/fintagging.
-0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:: Benchmarking LLMs for Extracting and Structuring Financial Information
-Yan Wang1, Lingfei Qian1, Xueqing Peng1, Yang Ren1, Yi Han2, Keyi Wang3, Dongji Feng4, Fengran Mo5, Shengyuan Lin6, Qinchuan Zhang7, Kaiwen He3, Chenri Luo3, Jianxing Chen3, Junwei Wu3, Chen Xu1, Ziyang Xu1, Jimin Huang1thanks:  Corresponding author, Guojun Xiong8, Xiao-Yang Liu3, Qianqian Xie1, Jian-Yun Nie5 1The Fin AI, 2Georgia Institute of Technology, 3Columbia University, 4California State University 5University of Montreal, 6Carnegie Mellon University, 7Rensselaer Polytechnic Institute, 8Harvard University Correspondence: jimin.huang@thefin.ai
-1 Introduction
-Although LLMs excel at question answering qian2025fino1, information extraction balasubramanian2025leveraging, and long-document summarization yuan2025strucsum, their ability to perform financial tagging remains underexplored.
-Unlike open-ended generation, tagging requires models to commit to a precise semantic interpretation for each reported number, grounding it in dispersed evidence across tables and surrounding text, and aligning it to a standardized concept that downstream systems can interpret consistently.
-Each year, over 2 million companies publish financial reports disclosing earnings, expenses, and liabilities, providing essential information for investors, regulators, and analysts.
+XBRL tagging, Benchmark, Large language model, Information extraction, Information retrieval, Reranking
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Natural language processingccs: Computing methodologies Information extractionccs: Information systems Information extractionccs: Information systems Top-k retrieval in databases
+1. Introduction
+Although LLMs excel at question answering (Qian et al., 2025), information extraction (Balasubramanian et al., 2025), and long-document summarization (Yuan et al., 2025), their ability to perform financial tagging remains underexplored.
+Unlike open-ended generation, tagging requires committing each reported number to a precise semantic interpretation, grounding it in mixed table and text evidence, and aligning it to standardized concepts interpretable by downstream systems.
+Each year, over 2 million companies publish financial reports disclosing earnings, expenses, and liabilities.
 However, inconsistent terminology complicates reliable interpretation.
-XBRL, introduced in 1999, defines the global standard for machine-readable financial reporting richards2006introduction and is now adopted in 65 countries [^3] https://www.xbrl.org/the-standard/what/what-is-xbrl/.
-In practice, XBRL tagging requires grounding thousands of reported numbers per filing in mixed table and text evidence and aligning them to over 17,000 standardized taxonomy concepts(Figure 1), a workflow that remains largely manual and error-prone [^4] https://xbrl.us/wp-content/uploads/2023/03/DQC-SECMeetingNotes-20240314.pdf.
-Figure 1: An example of Financial Tagging in realistic scenarios, where the blue numbers mark facts to be tagged.
-Nevertheless, current XBRL tagging benchmarks often emphasize what is easiest to annotate rather than what is hardest to understand.
-Many reduce tagging to classification over a truncated taxonomy and remove table structure that carries substantial semantic evidence in real filings (Table 1).
-As a result, existing datasets typically cover only a few thousand common concepts instead of the full 17,000+ taxonomy, which limits evaluation on rare but important concepts and makes it difficult to assess generalization under long-tailed label distributions.
-In addition, prior work commonly formulates tagging as closed-set extreme classification loukas2022finer; sharma2023financial.
-This framing places heavy weight on recalling label identifiers and becomes brittle when the full taxonomy cannot be represented in-context, increasing the likelihood of invalid or hallucinated tags.
-Consistent with this concern, our ablations show near-zero performance even for strong LLMs such as DeepSeek-V3 liu2024deepseek and GPT-4o hurst2024gpt under full-taxonomy settings.
-Finally, existing benchmarks focus primarily on narrative text and omit tables, despite tables being the main carrier of numerical disclosures.
-Consequently, they do not evaluate the joint reasoning over multi-level tabular structure and surrounding text that has been studied in financial QA chen2021finqa; zhu2021tat and is required for realistic end-to-end tagging.
-Table 1: Detailed comparison of financial NLP benchmarks across task types, sources, and structural capabilities.
+XBRL, introduced in 1999, defines the global standard for machine-readable financial reporting (Richards et al., 2006) and is now adopted in 65 countries [^3] https://www.xbrl.org/the-standard/what/what-is-xbrl/ .
+In practice, XBRL tagging requires grounding thousands of reported numbers per filing in mixed evidence and aligning them to over 17,000 standardized taxonomy concepts (Figure 1), a process that remains largely manual and error-prone [^4] https://xbrl.us/wp-content/uploads/2023/03/DQC-SECMeetingNotes-20240314.pdf.
+Figure 1. An example of Financial Tagging in realistic scenarios, where the blue numbers mark facts to be tagged.
+Nevertheless, existing XBRL tagging benchmarks emphasize annotation convenience over semantic difficulty.
+Many reduce tagging to classification over truncated taxonomies and remove tabular structure that carries critical evidence in real filings (Table 1).
+As a result, evaluation is limited to a small subset of common concepts, obscuring long-tailed generalization.
+Moreover, prior work typically formulates tagging as closed-set extreme classification (Loukas et al., 2022; Sharma et al., 2023), which becomes brittle when the full taxonomy cannot be represented in-context and increases the risk of invalid or hallucinated tags.
+Our ablations confirm this limitation: even strong LLMs show near-zero performance under full-taxonomy settings.
+In addition, most benchmarks omit tables entirely, despite tables being the primary carrier of numerical disclosures, and thus fail to evaluate the joint reasoning over structured tables and narrative context required in realistic tagging (Chen et al., 2021; Zhu et al., 2021).
+Table 1.
+Detailed comparison of financial NLP benchmarks across task types, sources, and structural capabilities.
 “Num.Reasoning” indicates the numerical reasoning.
@@
 Benchmark Scenario Data Source Task Modality #Entity Label #Taxonomy Label Num. Reasoning? Struct. IE? Concept Linking?
-FinQA chen2021finqa Decision making SEC 10-K QA text/table 0 0
-ConvFinQA chen2022convfinqa Decision making SEC 10-K QA text/table 0 0
-TAT-QA zhu2021tat Financial analysis Chinese financial reports QA text/table 0 0
-DocVQA mathew2021docvqa Enterprise automation Financial document images QA text/image 0 0
-FiNER-ORD shah2023finer Financial tagging Chinese financial disclosures Classification text 49 0
-FinRED sharma2022finred Knowledge graph construction English financial news Classification text 38 0
-FiNER loukas2022finer XBRL tagging SEC 10-K Extreme classification text 0 139 / 17,688
-FNXL sharma2023financial XBRL tagging SEC 10-K Extreme classification text 0 2,800 / 17,688
+FinQA (Chen et al., 2021) Decision making SEC 10-K QA text/table 0 0
+ConvFinQA (Chen et al., 2022) Decision making SEC 10-K QA text/table 0 0
+TAT-QA (Zhu et al., 2021) Financial analysis Chinese financial reports QA text/table 0 0
+DocVQA (Mathew et al., 2021) Enterprise automation Financial document images QA text/image 0 0
+FiNER-ORD (Shah et al., 2023) Financial tagging Chinese financial disclosures Classification text 49 0
+FinRED (Sharma et al., 2022) Knowledge graph construction English financial news Classification text 38 0
+FiNER (Loukas et al., 2022) XBRL tagging SEC 10-K Extreme classification text 0 139 / 17,688
+FNXL (Sharma et al., 2023) XBRL tagging SEC 10-K Extreme classification text 0 2,800 / 17,688
 FINTAGGING Financial & XBRL tagging SEC 10-K IE + Alignment text/table 5 17,688 / 17,688
-To address these issues, we introduce 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first LLM-oriented benchmark that evaluates whether models can recover the meaning of financial facts by extracting numerical values from mixed table and text inputs and aligning them to the full US-GAAP taxonomy.
-Unlike prior work loukas2022finer; sharma2023financial that formulates tagging as flat classification over a truncated label set, 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: requires end-to-end reasoning over both narrative disclosures and tables, and scales evaluation to the complete 17k+ concept space.
-0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: adopts a two-stage task formulation that separates decisions typically conflated in extreme classification.
-Financial Numerical Identification (FinNI) asks models to identify numerical facts together with their supporting evidence, while Financial Concept Linking (FinCL) maps each fact to the correct US-GAAP concept by semantic comparison with taxonomy definitions, reframing tagging as alignment rather than label recall.
-We release two expert-verified evaluation sets, FinNI-eval and FinCL-eval, derived from real-world XBRL filings, along with a unified protocol that evaluates both extraction accuracy and concept-alignment performance.
-Across 13 state-of-the-art LLMs, our results show a consistent extraction–linking gap.
-While top models reliably recover numerical facts from mixed table and text inputs, they struggle to assign the correct US-GAAP concepts, indicating that representations sufficient for extraction do not consistently encode the latent semantic conditions that distinguish closely related concepts, such as scope, aggregation level, and qualifying context.
-As a result, FinNI performance is strong, whereas FinCL remains below 0.19 even after candidate retrieval, with most errors arising from near-neighbor concepts that share similar surface cues but differ in these underlying conditions.
-Pipeline controls further clarify this failure mode: accuracy improves when gold entities are provided and improves again when the gold concept is included among candidates, yet it does not converge, showing that errors persist even when evidence is localized and the correct concept is available.
-The restricted cash example illustrates this pattern, where the model selects a narrower concept triggered by a salient local phrase, despite broader contextual evidence supporting an aggregate concept.
+To address these issues, we introduce 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first LLM-oriented benchmark that evaluates end-to-end financial tagging over mixed table and text inputs with alignment to the full US-GAAP taxonomy.
+Unlike prior work (Loukas et al., 2022; Sharma et al., 2023), 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: reframes tagging as an extract-and-align process rather than flat classification, enabling evaluation across all 17k+ concepts.
+0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: adopts a two-stage formulation.
+Financial Numerical Identification (FinNI) focuses on extracting numerical facts with supporting evidence, while Financial Concept Linking (FinCL) maps each fact to the correct US-GAAP concept via semantic alignment with taxonomy definitions.
+We release two expert-verified evaluation sets, FinNI-eval and FinCL-eval, derived from real-world XBRL.
 We conclude our main contributions as follows: (1) A New Formulation: We introduce 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, the first benchmark that reframes XBRL tagging into a two-stage “Extract-and-Align” paradigm, enabling full-scope evaluation against 17k+ US-GAAP concepts, a scale previously intractable for zero-shot LLMs.
 (2) High-Fidelity Resources: We release two expert-verified datasets, FinNI-eval and FinCL-eval, which incorporate complex tabular structures to assess joint reasoning across structured and unstructured financial modalities.
-(3) Diagnostic Insights: By benchmarking 13 SOTA LLMs, we identify a critical “knowledge-alignment gap”, demonstrating that while LLMs excel at numeric extraction, they require our alignment framework to prevent performance collapse during large-scale taxonomy mapping.
-2 Related Work
-From XBRL Classification to Alignment.
-Traditional XBRL tagging has been framed as extreme multi-class classification.
-Benchmarks such as FiNER loukas2022finer and FNXL sharma2023financial evaluate models on truncated label sets (139–2,800 tags), and methods like GalaXC saini2021galaxc, SECBERT loukas2022finer, and AttentionXML sharma2023financial focus on coping with label sparsity through hierarchical modeling.
-However, these closed-set formulations overlook the semantic reasoning required to navigate the full 17k+ US-GAAP taxonomy and are poorly matched to LLM capabilities.
-Recent LLM-based tools such as XBRL-Agent han2024xbrlagent explore filing analysis but do not evaluate end-to-end fact-to-taxonomy alignment.
-0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: addresses this gap by reframing tagging as a two-stage extract-and-align task.
-Financial Document Understanding.
-In parallel, financial NLP has progressed from entity and relation extraction (e.g., FiNER-ORD shah2023finer, FinRED sharma2022finred) to numerical reasoning over documents.
-Benchmarks such as FinQA chen2021finqa, TAT-QA zhu2021tat, and DocVQA mathew2021docvqa require joint reasoning over text and tables.
-Yet these tasks are designed for QA and do not constrain outputs to standardized accounting concepts.
-Our benchmark instead requires both multi-modal fact extraction and precise alignment to official US-GAAP definitions, focusing on the “last mile” of structured financial reporting.
-3 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:
-3.1 Task Formulation
-Formally, given a financial document $D$ consisting of multiple textual contents and structured tables, a set of predefined value data types $L$ specified by the XBRL specification, and a taxonomy database $\mathcal{T}$ containing a set of financial semantic concepts, the XBRL tagging task is to identify all relevant numerical values in $D$ and annotate each with a structured triplet $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$.
-Here, Tag denotes the linked concept in $\mathcal{T}$ (e.g.
-“us-gaap:CashAndDueFrom Banks”, etc.), Fact represents the extracted numerical value as it appears in context (e.g., “2.5”, “10,000”, “two”, etc.), and Type specifies the corresponding data type in $L$ (including monetaryItemType, percentItemType, sharesItemType, perShareItemType, and integerItemType).
-Each numerical fact is assigned a single, most specific tag from the taxonomy, without overlapping.
-Specifically, monetaryItemType refers to a financial amount, such as revenue or expenses, typically reported in currency units (e.g., USD).
-percentItemType represents a rate or ratio, expressed as a decimal (e.g., 25%).
-sharesItemType indicates the number of shares, such as stocks held or issued.
-perShareItemType captures values reported on a per-share basis, like earnings per share.
-Lastly, integerItemType is used for whole number counts, such as the number of employees or transactions.
+(3) Diagnostic Insights: By benchmarking 13 state-of-the-art LLMs, we identify a critical “knowledge-alignment gap”, demonstrating that while LLMs excel at numeric extraction, they require our alignment framework to prevent performance collapse during large-scale taxonomy mapping.
+2. Related Work
+Prior XBRL tagging research primarily formulates the task as extreme multi-class classification over truncated label sets, as in FiNER and FNXL (Loukas et al., 2022; Sharma et al., 2023), with methods such as GalaXC, SECBERT, and AttentionXML (Saini et al., 2021; Loukas et al., 2022; Sharma et al., 2023) addressing label sparsity via hierarchical modeling.
+However, these closed-set formulations do not reflect the semantic reasoning required to navigate the full 17k+ US-GAAP taxonomy and are poorly aligned with LLM capabilities.
+In parallel, financial document understanding has advanced from entity and relation extraction (Shah et al., 2023; Sharma et al., 2022) to joint text–table numerical reasoning in QA benchmarks such as FinQA, TAT-QA, and DocVQA (Chen et al., 2021; Zhu et al., 2021; Mathew et al., 2021), yet these tasks do not require alignment to standardized accounting concepts.
+Recent LLM-based tools such as XBRL-Agent (Han et al., 2024) explore filing analysis but stop short of end-to-end fact-to-taxonomy evaluation.
+In contrast, 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: reframes XBRL tagging as a two-stage extract-and-align task, explicitly targeting multi-modal fact extraction and fine-grained alignment to official US-GAAP definitions.
+3.
+0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:
+3.1. Task Formulation
+Formally, given a financial document $D$ containing text and structured tables, a set of predefined XBRL value types $L$, and a taxonomy $\mathcal{T}$ of financial concepts, the XBRL tagging task is to identify all numerical values in $D$ and annotate each with a structured triplet $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$.
+Here, Fact denotes the extracted numerical value as expressed in context, Type specifies its XBRL data type in $L$, and Tag refers to the linked concept in $\mathcal{T}$ (e.g., us-gaap:CashAndDueFromBanks).
+Each numerical fact is assigned a single, most specific taxonomy concept.
+The value type set $L$ includes monetaryItemType, percentItemType, sharesItemType, perShareItemType, and integerItemType, corresponding to monetary amounts, ratios, share counts, per-share values, and integer quantities, respectively.
 We formalize the task as a mapping:
-$f:(D,L,\mathcal{T})\mapsto\left\{(\texttt{Fact}_{i},\texttt{Type}_{i},\texttt{Tag}_{i})\right\}_{i=1}^{n}$ (1)
+(1) $f:(D,L,\mathcal{T})\mapsto\left\{(\texttt{Fact}_{i},\texttt{Type}_{i},\texttt{Tag}_{i})\right\}_{i=1}^{n}$
 where each triplet corresponds to a financial value mention extracted from $D$ and semantically grounded in $\mathcal{T}$.
-Inspired by information extraction and alignment works wu2019scalable; wadden2019entity; mukherjee2022ectsum; wang2022conditional and after discussing with financial reporting specialists, we formulated 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: into two sub-tasks: Financial Numeric Identification (FinNI), a multi-modal numerical information extraction task, detects numerical value in $D$ and classifies each with its appropriate Type.
+Inspired by information extraction and alignment works (Wu et al., 2019; Wadden et al., 2019; Mukherjee et al., 2022; Wang et al., 2022) and after discussing with financial reporting specialists, we formulated 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: into two sub-tasks: Financial Numeric Identification (FinNI), a multi-modal numerical information extraction task, detects numerical value in $D$ and classifies each with its appropriate Type.
 Financial Concept Linking (FinCL), a numerical entity normalization task, then associates each identified value with its most appropriate Tag in $\mathcal{T}$ based on contextual and structural cues.
-Financial Numeric Identification (FinNI)
+3.1.1. Financial Numeric Identification (FinNI)
 The first subtask of 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: focuses on identifying numerical values in a financial document and assigning each a coarse-grained value data type.
@@
 We formalize this subtask as a mapping:
-$f_{\text{FinNI}}:(D=(S,T),\ L)\mapsto\left\{(e_{i},l_{i})\right\}_{i=1}^{k}$ (2)
+(2) $f_{\text{FinNI}}:(D=(S,T),\ L)\mapsto\left\{(e_{i},l_{i})\right\}_{i=1}^{k}$
 where $S$ and $T$ represent the textual and tabular components of the document, respectively, and $L$ is the set of predefined value data types.
 Each $e_{i}$ is a numerical entity extracted from either $S$ or $T$, and $l_{i}\in L$ denotes its assigned data type.
-Financial Concept Linking (FinCL)
+3.1.2. Financial Concept Linking (FinCL)
 Building on the output of the FinNI subtask, the goal of FinCL is to semantically ground each identified numerical entity $e$ by linking it to a concept $\hat{c}$ in a predefined financial taxonomy, which is the Tag component of each triplet $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$ defined in the overall task.
 Formally, we define the mapping as:
-$f_{\text{FinCL}}:(e,l,C_{e},\mathcal{T})\mapsto\hat{c}$ (3)
+(3) $f_{\text{FinCL}}:(e,l,C_{e},\mathcal{T})\mapsto\hat{c}$
 where $e$ is a numerical entity extracted from $D=(S,T)$, and $l\in L$ is its associated data type provided by the FinNI stage.
@@
 The objective of FinCL is to identify the optimal concept $\hat{c}\in\mathcal{T}$ that maximizes the semantic alignment between the fact $e$ and its specific context $C_{e}$ within the high-dimensional taxonomy space.
-3.2 Raw Data Collection
+3.2. Raw Data Collection
 We compiled 142 annual 10-K filings from publicly listed U.S.
-companies submitted in 2023–2024 (Table 2).
-To ensure coverage diversity, the collection spans all 11 major industry sectors, with about 95% of companies located across 31 U.S.
-states and the remainder based internationally.
-Using BeautifulSoup, we parsed the reports and extracted 319,893 narrative sentences and 21,576 financial tables.
-Company-level statistics are provided in Appendix C.1.
-All filings follow the SEC Inline XBRL (iXBRL) format, which embeds machine-readable tags inside HTML disclosures and links textual and tabular values to US-GAAP taxonomy concepts through schema and linkbase definitions.
-However, these regulatory tags are designed for compliance rather than model evaluation.
-They provide little coverage of negative cases and contain substantial structural noise.
-We therefore construct a high-fidelity evaluation suite that repurposes this raw regulatory metadata into a standardized benchmark for financial fact extraction and concept linking.
-Additional discussion of our benchmark granularity is provided in Appendix C.2.
-Table 2: Statistics of collected financial reports.
+companies submitted to the U.S.
+Securities and Exchange Commission (SEC) [^5] https://www.sec.gov/ in 2023–2024 (Table 2).
+The collection covers all 11 major industry sectors, with companies primarily based in the United States and a small portion internationally.
+Parsing the filings with BeautifulSoup yields 319,893 narrative sentences and 21,576 financial tables.
+All filings follow the SEC Inline XBRL (iXBRL) format, which embeds machine-readable tags within HTML disclosures and links textual and tabular values to US-GAAP taxonomy concepts.
+However, these regulatory tags are designed for compliance rather than model evaluation, providing limited negative coverage and substantial structural noise.
+We therefore repurpose the raw iXBRL metadata to construct a standardized, high-fidelity evaluation suite for financial fact extraction and concept linking.
+Table 2. Statistics of collected financial reports.
 Item Value
@@
 #Table 21,576
-3.3 Large-Scale Automated Annotation
+3.3. Large-Scale Automated Annotation
 Our automated pipeline produces a large, high-fidelity corpus reflecting real financial disclosures.
@@
 Including both valid facts and naturally occurring noise, it supports a realistic evaluation of LLMs under regulatory-style conditions.
-Table 3: Statistical information for the original dataset in our benchmark. Tokens are calculated using the “cl100k_base” tokenizer ($\pm$ standard deviation).
+Table 3. Statistical information for the original dataset in our benchmark. Tokens are calculated using the “cl100k_base” tokenizer ($\pm$ standard deviation).
 Item Sentence Table
@@
 Automated Annotation Engine.
-We build a rule-based parser (Algorithm 1, Appendix C.3) that reconstructs semantic links between human-readable disclosures and Inline XBRL (iXBRL) metadata.
-For each instance $i$, the parser locates ix:nonfraction elements and extracts the value $v$, concept $c$, and data type $\tau$, while restricting extraction to atomic leaf elements $\epsilon$ to avoid duplicates from nested tags.
+We build a rule-based parser that reconstructs semantic links between human-readable disclosures and Inline XBRL (iXBRL) metadata.
+For each instance $i$, the parser extracts iXBRL fact elements (e.g., ix:nonfraction) and retrieves the value $v$, concept $c$, and data type $\tau$, implemented as ParseIXBRL in Algorithm 3.
+To avoid duplicates from nested tags, we restrict extraction to atomic leaf elements $\epsilon$.
+Input: Raw iXBRL reports $\mathcal{D}$; taxonomy $\mathcal{T}$; valid types $\mathcal{L}$
+Output: Positive instances $\mathcal{P}$; Negative instances $\mathcal{N}$
+Initialize $\mathcal{P}\!\leftarrow\!\emptyset$, $\mathcal{N}\!\leftarrow\!\emptyset$, seen text set $\mathcal{S}\!\leftarrow\!\emptyset$;
+foreach instance $i\in\mathcal{D}$ (sentence/table) do
+  $text\leftarrow\textsc{ExtractText}(i)$;
+  if $|text|\leq 20$ or $text\in\mathcal{S}$ then
+    continue
+ $E\leftarrow\{(v,\tau,c)\mid(v,c,\tau)\!\leftarrow\!\textsc{ParseIXBRL}(i),\ c\!\in\!\mathcal{T},\ \tau\!\in\!\mathcal{L},\ v\neq\emptyset\}$;
+  $E\leftarrow\textsc{DedupBy}(E,\ (v,c))$;
+ if $E\neq\emptyset$ then
+     $\mathcal{P}\leftarrow\mathcal{P}\cup\{(text,E)\}$;
+ else
+     $\mathcal{N}\leftarrow\mathcal{N}\cup\{(text,\emptyset)\}$;
+ $\mathcal{S}\leftarrow\mathcal{S}\cup\{text\}$;
+return $\mathcal{P},\mathcal{N}$;
+Algorithm 1 Automated Instance Annotation
 Valid Entity Types.
@@
 Structural Refinement and Filtering.
-For reliable evaluation, we keep only text segments longer than 20 characters and deduplicate content globally using $\mathcal{S}_{text}$.
+For reliable evaluation, we keep only text segments longer than 20 characters and deduplicate content globally using texttext.
 For tables, we retain structural tags (<table>, <tr>, <td>, <th>) but discard tables without numerical facts.
 This filtering preserves semantic accuracy while maintaining realistic layout structure.
-Figure 2: Statistics of numerical entity types.
-3.4 Data Quality Control
+Figure 2. Statistics of numerical entity types.
+3.4. Data Quality Control
 Validating all 261,457 entities is infeasible.
-To obtain a rigorous quality upper bound, we adopt a strategic sampling scheme based on a greedy minimal-cover algorithm (Algorithm 2, Appendix D.1).
-Rather than random sampling, we construct a diversity-maximized subset $\mathcal{P}^{*}$ that satisfies two goals: (1) Company Balance, limiting bias from individual reporting styles, and (2) Taxonomy Completeness, ensuring that each US-GAAP concept observed in the corpus appears at least once.
-We seed $\mathcal{P}^{*}$ with capped samples per company, then iteratively add instances that cover the largest number of uncovered concepts, producing a compact but representative audit set.
-The final subset contains 928 positive sentence instances and 1,021 positive table instances, serving as a taxonomy-wide stress test reviewed by two junior analysts and one senior adjudicator (the demography of these annotators is shown in Appendix D.2).
+We therefore adopt a strategic sampling scheme based on a greedy minimal-cover method (algorithm 4) to construct a compact audit set $\mathcal{P}^{*}$.
+The procedure balances two objectives: Company Balance, by capping the number of sampled instances per company, and Taxonomy Completeness, by ensuring coverage of all observed US-GAAP concepts.
+Specifically, we first seed $\mathcal{P}^{*}$ with up to $K$ instances per company, then iteratively add instances that cover the largest number of remaining uncovered concepts.
+This process yields a diversity-maximized subset consisting of 928 sentence instances and 1,021 table instances, which is reviewed by two junior analysts and one senior adjudicator.
+Input: Positive instances $\mathcal{P}$; observed concepts $\mathcal{C}_{all}$; cap $K=10$
+Output: Strategic subset $\mathcal{P}^{*}$
+$\mathcal{P}^{*}\!\leftarrow\!\emptyset$; $\mathcal{C}_{rem}\!\leftarrow\!\mathcal{C}_{all}$;
+foreach company $g$ do
+  Add up to $K$ instances from $g$ to $\mathcal{P}^{*}$;
+  $\mathcal{C}_{rem}\leftarrow\mathcal{C}_{rem}\setminus\textsc{Concepts}(\mathcal{P}^{*})$;
+while $\mathcal{C}_{rem}\neq\emptyset$ and $\mathcal{P}\setminus\mathcal{P}^{*}\neq\emptyset$ do
+  $i^{*}\leftarrow\arg\max_{i\in\mathcal{P}\setminus\mathcal{P}^{*}}|\textsc{Concepts}(i)\cap\mathcal{C}_{rem}|$;
+  if ties then
+    select $i^{*}$ with the shortest text length;
+ $\mathcal{P}^{*}\leftarrow\mathcal{P}^{*}\cup\{i^{*}\}$; $\mathcal{C}_{rem}\leftarrow\mathcal{C}_{rem}\setminus\textsc{Concepts}(i^{*})$;
+return $\mathcal{P}^{*}$;
+Algorithm 2 Strategic Audit Set Sampling
 Review Protocol.
 The audit followed two stages.
-A pilot phase aligned interpretations of the Verification Guideline (Appendix D.3), after which analysts independently verified whether extracted triplets $(e,\tau,c)$ were fully correct.
+A pilot phase aligned interpretations of the Verification Guideline [^6] https://docs.google.com/document/d/1TgTxz-fozBImeRs8Y9kLViHnr8Bjs1qZbsB8u5NjfmA/edit?usp=sharing, after which analysts independently verified whether extracted triplets $(v,\tau,c)$ were fully correct.
 The senior expert adjudicated disagreements to create the final gold set.
 Agreement Results.
-On the independent round, annotators achieved 96% raw agreement and Cohen’s $\kappa=0.81$, indicating substantial agreement landis1977measurement, the equation is shown in Appendix D.4.
+On the independent round, annotators achieved 96% raw agreement and Cohen’s $\kappa=0.81$, indicating substantial agreement (Landis and Koch, 1977).
 This high-verified accuracy on the most semantically diverse subset provides strong evidence for the reliability of our automated annotation pipeline.
-3.5 Task-Specific Dataset Construction
+3.5. Task-Specific Dataset Construction
 Following validation, we integrated the corrected $\mathcal{P}^{*}$ subset back into the full corpus to form a high-fidelity hybrid dataset.
@@
 Their detailed statistics are reported in Table 4.
-Table 4: Statistics of evaluation datasets for FinNI-eval and FinCL-eval. Tokens are calculated using the “cl100k_base” tokenizer ($\pm$ standard deviation).
+Table 4. Statistics of evaluation datasets for FinNI-eval and FinCL-eval. Tokens are calculated using the “cl100k_base” tokenizer ($\pm$ standard deviation).
 Item FinNI-eval FinCL-eval
@@
 Max Output Tokens 7,813 41
-3.5.1 FinNI-eval Dataset Construction
-Based on the annotation procedure described in Section 3.3, we construct the FinNI-eval dataset, comprising 28,787 instances derived from the annotated sentences and tables.
-Specifically, the dataset comprises two components: the input block and the answer block.
-The input block includes a task instruction and the input content.
-The instruction defines the FinNI task by specifying the responsibilities of the LLMs, the definitions of entity types, and the rules to be followed.
-The answer block is formatted as a list aligned with the input block.
-The answer is a JSON list of all identified entity values and types or an empty JSON list.
-The specific prompt template is shown in Figure 9.
+3.5.1. FinNI-eval Dataset Construction
+Following the annotation procedure in Section 3.3, we construct the FinNI-eval dataset with 28,787 instances derived from annotated sentences and tables.
+Each instance consists of an input block and an answer block.
+The input block includes a task instruction and the corresponding content, while the answer block is a JSON list of identified numerical values with their types (or an empty list).
 FinNI-eval Dataset
@@
         "Type": <entity type>}]} or {"results": []}
-3.5.2 FinCL-eval Dataset Construction
-Following the formulation of the FinCL subtask in Section 3.1, we further construct a FinCL-eval dataset that includes 261,457 query-answer pairs for numerical entity normalization.
-In addition, we built a US-GAAP taxonomy database containing 17,688 unique financial concepts.
-To create the FinCL-eval dataset, we utilize all positive instances identified in Section 3.3.
-Each query consists of a numerical entity, its corresponding entity type, and its surrounding context, while the answer is the associated US-GAAP concept.
+3.5.2. FinCL-eval Dataset Construction
+Following the FinCL formulation in Section 3.1.2, we construct the FinCL-eval dataset with 261,457 query–answer pairs for numerical entity normalization, along with a US-GAAP taxonomy containing 17,688 concepts.
+Each query includes a numerical entity, its type, and surrounding context, while the answer is the corresponding US-GAAP concept.
 FinCL-eval Dataset
@@
 Answer: <US-GAAP Concept>
-3.6 Evaluation
-3.6.1 Evaluation Framework
+3.6. Evaluation
 We propose a unified evaluation framework for 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:.
-As illustrated in Figure 3, each input is decomposed into two subtasks, FinNI and FinCL, which together produce structured triplets $\{\texttt{Fact},\texttt{Type},\texttt{Tag}\}$.
-This design jointly evaluates zero-shot fact extraction and concept alignment.
-For comparison with token-classification baselines, we report macro and micro Precision, Recall, and F1 sokolova2009systematic.
-FinNI is evaluated using pair-level metrics, and FinCL with accuracy.
-Full metric definitions are in Appendix E.
-Figure 3: A unified evaluation framework on 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark.
-Note: The Input content is drawn from the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: dataset, and the “Shared LLMs” are used for identification in the FinNI task and reranking in the FinCL task.
-FinNI Evaluation. We adopt pair-level metrics to measure extraction quality, focusing on whether models can locate financial values from mixed text–table contexts based on semantic cues.
+As shown in Figure 3, each input is decomposed into two subtasks, which together produce structured triplets.
+This design jointly evaluates zero-shot numerical extraction and concept alignment.
+For comparison with token-classification baselines, we report macro and micro Precision, Recall, and F1 (Sokolova and Lapalme, 2009).
+FinNI is evaluated using pair-level metrics, while FinCL is evaluated with accuracy.
+FinNI Evaluation. FinNI uses pair-level metrics to assess whether models correctly extract numerical values from text-table inputs.
 FinCL Evaluation.
-To test fine-grained numeric entity normalization, we formulate FinCL as a retrieval–reranking task: a retrieval stage produces a candidate set of taxonomy concepts, and the LLM then reranks candidates to select the most appropriate tag, avoiding impractical extreme classification (prompt template in Figure 10, Appendix).
-To this end, we embed each taxonomy concept using text-embedding-3-small [^5] https://platform.openai.com/docs/models/text-embedding-3-small , then retrieve the top-$h$ candidate tags from the US-GAAP taxonomy $\mathcal{T}$ by semantic similarity.
-The LLM subsequently reranks these candidates using contextual reasoning to select the best tag $\hat{c}\in\mathcal{T}_{h}$:
-$f_{\text{rerank}}:(e,l,C_{e},\mathcal{T}_{h})\mapsto\hat{c}$ (4)
-3.6.2 Evaluation Models
-Our goal is to assess the foundational capabilities of SOTA LLMs on 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: and understand their strengths and limitations in XBRL tagging.
-We evaluate models across three categories using our unified framework (Table 11).
-(1) Closed-source general LLM: GPT-4o hurst2024gpt.
-(2) Open-source general LLMs: DeepSeek-V3 liu2024deepseek, DeepSeek-R1-Distill-Qwen-32B deepseekai2025deepseekr1incentivizingreasoningcapability, Qwen3-32B / 14B / 1.7B / 0.6B qwen3technicalreport, Llama-4-Scout-17B-16E-Instruct meta2025llama, Llama-3.3-70B-Instruct, Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct grattafiori2024llama, and gemma-2-27b-it gemma_2024.
-(3) Domain-specific financial LLM: Fino1-8B qian2025fino1.
-We also compare against strong PLM baselines: BERT-large DBLP:journals/corr/abs-1810-04805, FinBERT araci2019finbert, and SECBERT loukas2022finer.
-3.6.3 Evaluation setting
+FinCL is formulated as a retrieval–reranking task to avoid impractical extreme classification.
+We embed each taxonomy concept using text-embedding-3-small [^7] https://platform.openai.com/docs/models/text-embedding-3-small and retrieve the top-$h$ candidates from the US-GAAP taxonomy by semantic similarity.
+The LLM then reranks candidates using contextual reasoning to select the final tag.
+Figure 3.
+A unified evaluation framework on 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark.
+Note: The Input content is drawn from the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: dataset, and the ”Shared LLMs” are used for identification in the FinNI task and reranking in the FinCL task.
+3.6.1. Evaluation Models
 […]
-4 Experiment and Result
-4.1 Overall Results
+4. Experiment and Result
+4.1. Overall Results
 Table 5 presents the overall performance on the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark.
 It clearly demonstrates that under our framework, LLMs can effectively handle both frequent and rare financial tags, indicating their ability to mitigate long-tail label challenges and underscoring the advantage of our information extraction and alignment formulation over traditional token-level classification approaches.
-Table 5: Overall Performance. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
-Category Models Macro Micro
+Table 5. Overall Performance. Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
+Models Macro Micro
 P R F1 P R F1
-Closed-source LLM GPT-4o 0.0865 0.0626 0.0569 0.1052 0.0731 0.0863
-Open-source LLMs DeepSeek-V3 0.0949 0.0778 0.0659 0.1074 0.1264 0.1162
+GPT-4o 0.0865 0.0626 0.0569 0.1052 0.0731 0.0863
+DeepSeek-V3 0.0949 0.0778 0.0659 0.1074 0.1264 0.1162
 Llama-4-Scout-17B-16E-Instruct 0.0683 0.0414 0.0400 0.1045 0.0526 0.0699
@@
 Qwen3-0.6B 0.0026 0.0008 0.0010 0.0562 0.0002 0.0005
-Financial LLM Fino1-8B 0.0344 0.0143 0.0151 0.0419 0.0128 0.0197
-Fine-tuned PLMs BERT-large 0.0252 0.0266 0.0205 0.1518 0.1283 0.1391
+Fino1-8B 0.0344 0.0143 0.0151 0.0419 0.0128 0.0197
+BERT-large 0.0252 0.0266 0.0205 0.1518 0.1283 0.1391
 FinBERT 0.0046 0.0064 0.0042 0.0872 0.0526 0.0656
@@
 Qwen3-32B also performs competitively, suggesting that model architecture and pretraining can compensate for a smaller scale.
-From a micro perspective, which reflects performance on frequent labels, the trend is similar: top LLMs remain competitive even without fine-tuning, while fine-tuned PLMs such as SECBERT and BERT-large excel mainly because they benefit from label frequency rather than broader coverage (Table 5).
-4.2 Subtask Results
-4.2.1 FinNI subtask
-Table 6 summarizes results on FinNI.
-Overall, larger LLMs show markedly stronger ability to identify numerical entities and produce structured outputs, even without financial adaptation.
-DeepSeek-V3 achieves the best overall performance, while GPT-4o performs competitively and ranks among the top models.
-Table 6: Performance comparison of different models on the FinNI subtask.
+From a micro perspective, which reflects performance on frequent labels, the trend is similar: top LLMs remain competitive even without fine-tuning, while fine-tuned PLMs such as SECBERT and BERT-large excel mainly because they benefit from label frequency rather than broader coverage.
+4.2. Subtask Results
+Table 6 summarizes performance on the FinNI and FinCL subtasks.
+Overall, large LLMs outperform smaller models across both tasks in a zero-shot setting.
+On FinNI, DeepSeek-V3 achieves the best F1 score, driven by substantially higher recall, while GPT-4o performs competitively.
+In contrast, several models (e.g., Qwen3 variants) exhibit high precision but extremely low recall, indicating that they extract only a small fraction of target facts, which leads to poor F1 performance.
+Mid-sized models such as gemma-2-27B show more balanced behavior.
+In contrast, smaller models and the financial LLM Fino1-8B lag, suggesting that domain pretraining alone does not yield robust gains for numerical extraction.
+Table 6.
+Performance comparison of different models on the FinNI and FinCL subtasks.
 Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
-Category Model Precision Recall F1
-Closed-source LLM GPT-4o 0.5893 0.4977 0.5397
-Open-source LLMs DeepSeek-V3 0.5886 0.8430 0.6932
-Llama-4-Scout-17B-16E-Instruct 0.4668 0.3164 0.3771
-Llama-3.3-70B-Instruct 0.4826 0.3301 0.3920
-DeepSeek-R1-Distill-Qwen-32B 0.5676 0.1942 0.2894
-Qwen3-32B 0.6991 0.1804 0.2868
-gemma-2-27b-it 0.5060 0.4526 0.4778
-Qwen3-14B 0.6912 0.1487 0.2448
-Llama-3.1-8B-Instruct 0.3874 0.1761 0.2421
-Llama-3.2-3B-Instruct 0.1856 0.1203 0.1460
-Qwen3-1.7B 0.7362 0.0281 0.0541
-Qwen3-0.6B 0.2803 0.0019 0.0038
-Financial LLM Fino1-8B 0.3431 0.1293 0.1878
-A notable pattern is that some models (e.g., Qwen3 variants) obtain very high precision but extremely low recall, indicating that they extract only a small fraction of target facts.
-This confirms that precision alone is insufficient for FinNI, where missing facts are often more damaging than producing a few spurious ones.
-Smaller open-source models struggle severely, reflecting the difficulty of this zero-shot extraction setting.
-In contrast, mid-size models such as gemma-2-27B achieve more balanced behavior.
-Interestingly, the financial LLM Fino1-8B still lags behind general-purpose models, suggesting that domain text pretraining alone does not help unless the training task aligns closely with FinNI.
-4.2.2 FinCL subtask
-Table 7 reports FinCL performance under the reranking setup, using candidates retrieved from the best top-200 configuration (Table 14).
-Although retrieval provides essential coverage, low Acc@k values, especially for table entities, substantially cap downstream accuracy.
-Table 7: Performance comparison of different models on the FinCL subtask.
-Bolded values indicate the best performance, underlined values represent the second-best, and italicized values denote the third-best performance.
-Category Model Accuracy
-Closed-source LLM GPT-4o 0.1829
-Open-source LLMs DeepSeek-V3 0.1889
-Llama-4-Scout-17B-16E-Instruct 0.1649
-Llama-3.3-70B-Instruct 0.1318
-DeepSeek-R1-Distill-Qwen-32B 0.1141
-Qwen3-32B 0.1277
-gemma-2-27b-it 0.1099
-Qwen3-14B 0.1144
-Llama-3.1-8B-Instruct 0.0913
-Llama-3.2-3B-Instruct 0.0415
-Qwen3-1.7B 0.0735
-Qwen3-0.6B 0.0414
-Financial LLM Fino1-8B 0.0807
-Within this setting, large LLMs such as DeepSeek-V3, GPT-4o, and Llama-4-Scout perform best, but overall accuracy remains low, reflecting the difficulty of aligning evidence with fine-grained taxonomy concepts.
-Most other models cluster well below this tier, and smaller models approach random behavior.
-The financial model Fino1-8B also lags behind general LLMs, suggesting that domain pretraining alone is insufficient without task-aligned supervision.
-Taken together, FinCL exposes two compounding challenges: taxonomy disambiguation requires reasoning over subtle semantic cues, and retrieval errors propagate directly into reranking, creating a ceiling effect on end-to-end alignment.
-4.3 Ablation analysis
+Model FinNI FinCL
+Precision Recall F1 Accuracy
+GPT-4o 0.5893 0.4977 0.5397 0.1829
+DeepSeek-V3 0.5886 0.8430 0.6932 0.1889
+Llama-4-Scout-17B-16E-Instruct 0.4668 0.3164 0.3771 0.1649
+Llama-3.3-70B-Instruct 0.4826 0.3301 0.3920 0.1318
+DeepSeek-R1-Distill-Qwen-32B 0.5676 0.1942 0.2894 0.1141
+Qwen3-32B 0.6991 0.1804 0.2868 0.1277
+gemma-2-27b-it 0.5060 0.4526 0.4778 0.1099
+Qwen3-14B 0.6912 0.1487 0.2448 0.1144
+Llama-3.1-8B-Instruct 0.3874 0.1761 0.2421 0.0913
+Llama-3.2-3B-Instruct 0.1856 0.1203 0.1460 0.0415
+Qwen3-1.7B 0.7362 0.0281 0.0541 0.0735
+Qwen3-0.6B 0.2803 0.0019 0.0038 0.0414
+Fino1-8B 0.3431 0.1293 0.1878 0.0807
+FinCL remains significantly more challenging than FinNI.
+All FinCL results are reported under a retrieval–reranking setup, where models select the final tag from the top-200 retrieved US-GAAP candidates.
+Although retrieval results are not separately tabulated due to space constraints, structure-aware retrieval achieves over 0.65 Acc@200 for sentence-based entities but remains below 0.30 Acc@200 for table-based entities, resulting in an overall Acc@200 of approximately 0.33.
+This indicates that retrieval coverage is already limited, especially for tabular contexts, which imposes a ceiling on downstream reranking.
+Even under this constrained candidate space, the best-performing models achieve accuracy below 0.19, with DeepSeek-V3 and GPT-4o forming a narrow top tier.
+These results highlight a clear extraction–alignment gap: while modern LLMs can recover numerical facts with reasonable accuracy, fine-grained US-GAAP disambiguation remains intrinsically difficult, particularly for table-origin entities.
+4.3. Ablation analysis
 We further compare our benchmark against extreme multi-class classification settings.
-The prompt template is shown in Figure 11 in the Appendix, and we select the best-performing model from each category.
+We select the best-performing model from each category.
 The evaluation uses triplet-level ($\{\texttt{Tag},\texttt{Fact},\texttt{Type}\}$) Precision, Recall, and F1.
-Table 8: Performance comparison between with/without our evaluation framework on the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark dataset.
+Table 7.
+Performance comparison between with/without our evaluation framework on the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark dataset.
 Evaluation Mode Model Precision Recall F1
@@
 Fino1-8B 0 0 0
-As shown in Table 8, treating XBRL tagging as single-step extreme classification fails in practice.
-The model must choose directly from thousands of taxonomy concepts without first identifying valid numerical entities, forcing it to rely on internal priors rather than grounded evidence.
-All models collapse to zero, illustrating that this formulation is not a meaningful way to evaluate LLMs on XBRL tagging.
-In contrast, our two-stage framework mirrors how tagging is actually performed.
-FinNI first detects and types candidate numerical facts, and FinCL then retrieves a small set of plausible tags and requires the model to select among them using context.
-Under this formulation, models achieve non-trivial performance, revealing real differences in capability that are completely hidden under extreme classification.
-4.4 Error propagation in 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: pipeline
-Table 9 compares accuracy across three settings: the full pipeline, FinCL with gold entities, and rerank-only with guaranteed gold candidates.
-The pattern reveals clear cascading errors.
-End-to-end performance is consistently the lowest because mistakes from extraction and retrieval accumulate in the linking stage.
-When gold entities are supplied, accuracy improves substantially, showing that entity identification is a major bottleneck.
-Performance increases again in the rerank-only setting, indicating that narrowing the label space reduces difficulty.
-However, accuracy still remains far from solved, suggesting that fine-grained US-GAAP disambiguation is intrinsically challenging even when retrieval is perfect.
-In addition, we provide representative error cases in Appendix I, illustrating how these challenges arise in practice.
-Table 9: Accuracy comparison across the full pipeline, FinCL subtask, and rerank-only settings.
+Table 7 shows that modeling XBRL tagging as single-step extreme classification fails in practice: models must select from thousands of taxonomy concepts without grounding in identified numerical entities, leading all models to collapse to zero.
+In contrast, our two-stage framework reflects real tagging workflows: FinNI first identifies and types numerical facts, and FinCL then selects from a retrieved candidate set using context.
+This formulation yields non-trivial performance and exposes meaningful differences among models that are entirely obscured under extreme classification.
+4.4.
+Error propagation in 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: pipeline
+Table 8 compares accuracy under three settings: the full pipeline, FinCL with gold entities, and rerank-only with gold candidates.
+Results show clear error cascading: end-to-end performance is lowest due to accumulated extraction and retrieval errors.
+Providing gold entities substantially improves accuracy, identifying entity extraction as a key bottleneck.
+Rerank-only settings further increase performance, yet accuracy remains low, indicating that fine-grained US-GAAP disambiguation is inherently challenging even with perfect retrieval.
+Table 8. Accuracy comparison across the full pipeline, FinCL subtask, and rerank-only settings.
 Model Pipeline FinCL Rerank-only
@@
 Fino1-8B 0.0419 0.0807 0.0908
-5 Conclusion
+5. Conclusion
 This paper introduces 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:, a benchmark for evaluating LLMs on XBRL tagging of real financial reports.
@@
 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: provides a foundation for future work on reliable XBRL tagging and regulatory reporting.
-Limitations
-In this work, we propose 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: and conduct a comprehensive analysis of LLM performance on the task.
-However, several limitations remain.
-First, we have not yet evaluated the latest models, such as GPT-5.2, gpt-oss, and Qwen3-235B.
-We are currently conducting these experiments and will report the results in future work.
-Second, we collected only a small dataset for fine-tuning PLMs and did not construct task-specific training data for LLMs.
-Future work may explore building LLM-based XBRL tagging agents through targeted fine-tuning to enhance performance further.
-Potential Risks
-Potential Positive Societal Impacts.
-0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: provides the first structure-aware benchmark for evaluating large language models on realistic XBRL tagging.
-By reframing tagging as a transparent, two-stage Extract–Align process, this benchmark:
-  • Enables more rigorous and reproducible evaluation of model performance on financial fact extraction and taxonomy alignment.
-  • Supports research on safer automation of financial workflows by exposing failure modes that arise in long-tailed concept spaces and complex tabular contexts.
-  • Lowers the barrier for auditing, interpreting, and improving financial AI systems by releasing open datasets, evaluation protocols, and diagnostic analyses.
-By emphasizing open evaluation rather than proprietary pipelines, 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: contributes to more transparent and accountable financial AI research.
-Potential Negative Societal Impacts.
-Despite these benefits, several risks remain:
-  • Misuse and Overreliance.
-High benchmark performance may be mistakenly interpreted as readiness for real-world XBRL tagging, regulatory reporting, or financial decision-making, even though the benchmark does not replace human review or auditing.
-  • Bias and Coverage Limitations.
-The dataset is drawn from U.S.
-SEC filings and US-GAAP concepts.
-As a result, reporting practices, sector distributions, and taxonomy frequency patterns may introduce biases and may not generalize to other jurisdictions or accounting standards.
-  • Benchmark Narrowness.
-0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: evaluates extraction and alignment only.
-Downstream tasks such as aggregation, reconciliation, audit validation, and materiality assessment remain outside the benchmark and may still fail in deployed systems.
-Ongoing iteration, external feedback, and careful documentation are necessary to mitigate these concerns.
-Ethical Considerations and Usage Disclaimer
-The authors assume responsibility for the design and dissemination of 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:.
-All data used in the benchmark are derived from publicly available SEC filings and contain no intentionally included personal or sensitive information.
-Dataset construction followed community ethical guidelines for research involving public financial disclosures.
-Unless otherwise specified, the datasets and accompanying evaluation resources are released for research and educational purposes.
-They should not be used as a substitute for professional accounting judgment, financial reporting systems, or regulatory compliance procedures.
-The benchmark is designed to evaluate model behavior under controlled conditions.
-Performance on 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863I0.43922 0.45098 0.95294N0.5098 0.43529 0.93725T0.58431 0.42353 0.92157A0.6549 0.40784 0.90588G0.72549 0.39216 0.8902G0.79608 0.37647 0.87451I0.87059 0.36471 0.85882N0.94118 0.34902 0.84314G\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: does not imply suitability for production deployment, financial auditing, or automated investment or compliance tasks.
-Any real-world use must include human oversight, documentation of limitations, and appropriate legal and regulatory review.
-We used AI-assisted writing tools to help refine the wording and improve the readability of the manuscript.
-All scientific claims, dataset design choices, and experimental analyses were developed and verified by the authors.
-While reasonable care has been taken to ensure accuracy, the authors provide no warranty regarding completeness or suitability for any particular purpose and disclaim liability for any damages arising from use of the benchmark or associated materials.
-Licensing and Terms of Use
-We release all benchmark data, annotations, and accompanying resources under the CC-BY 4.0 license, which permits redistribution and adaptation with attribution.
-All source documents originate from publicly available SEC filings, and we strictly comply with the original terms governing their public distribution.
-The US-GAAP taxonomy is used in accordance with its publicly documented terms.
-Our released benchmark does not contain proprietary material beyond what is already publicly accessible, and any derivative use must continue to acknowledge both the original providers and our benchmark.
 Appendix A Significance analysis
@@
 (b) Per-model count of significant pairwise differences.
-Figure 4: Visualization of significance analysis across models.
+Figure 4.
+Visualization of significance analysis across models.
 (a) shows the pairwise-bootstrap significance matrix, while (b) summarizes how many significant differences each model has relative to others.
@@
 Applying NER to assign XBRL tags is an emerging yet challenging area.
-B.1 XBRL Tagging benchmark
-FiNER systematically benchmarked several neural architectures on the finer-139 dataset to address numeric-heavy XBRL tagging loukas2022finer.
+B.1. XBRL Tagging benchmark
+FiNER systematically benchmarked several neural architectures on the finer-139 dataset to address numeric-heavy XBRL tagging (Loukas et al., 2022).
 The initial experiments showed that standard BERT underperforms due to subword fragmentation, then the authors introduced pseudo-token strategies replacing numerals with [NUM] or [SHAPE] tokens to stabilize label assignment across fragmented numeric spans.
@@
 Their experiments demonstrated that subword-aware models with numeric-aware pseudo-tokens outperform both word-level BiLSTMs and vanilla BERT, particularly in numeric-heavy contexts, and avoid nonsensical label sequences.
-FNXL extended this benchmarking paradigm to a much larger label space of 2,794 US-GAAP tags, reframing the task as an extreme classification problem sharma2023financial.
+FNXL extended this benchmarking paradigm to a much larger label space of 2,794 US-GAAP tags, reframing the task as an extreme classification problem (Sharma et al., 2023).
 They compared the FiNER sequence labeling approach with a two-step pipeline that first identifies numeric spans and then assigns labels using AttentionXML.
@@
 Together, these benchmarks reveal the need for context-aware reasoning and label-ranking mechanisms in realistic XBRL tagging scenarios.
-B.2 XBRL Tagging Methods
+B.2. XBRL Tagging Methods
 The previous studies also explored approaches to address scalability, semantic ambiguity, and reasoning gaps in XBRL tagging to improve performance.
 Saini et al.
-saini2021galaxc proposed GalaXC is a graph-based extreme classification framework that jointly learns over document-label graphs with per-label attention across multi-hop neighborhoods.
+(Saini et al., 2021) proposed GalaXC is a graph-based extreme classification framework that jointly learns over document-label graphs with per-label attention across multi-hop neighborhoods.
 By integrating label metadata and transitive label correlations, GalaXC outperformed leading deep classifiers by up to 18% in micro-F1 on standard benchmarks and achieved 25% gains in warm-start scenarios where partial labels are available.
 Moreover, Wang et al.
-wang2023standardizing addressed the practical challenge of custom tag standardization through a semantic similarity pipeline that leverages TF-IDF, Word2Vec, and FinBERT embeddings.
+(Wang, 2023) addressed the practical challenge of custom tag standardization through a semantic similarity pipeline that leverages TF-IDF, Word2Vec, and FinBERT embeddings.
 Although unsupervised, the method was tested across nearly 200,000 custom tags from SEC filings between 2009 and 2022, and showed strong alignment performance, with vector-based mappings identifying viable standard tag candidates for a substantial proportion of non-compliant elements, offering a low-cost and interpretable solution for downstream financial analysis.
-Shifting focus from classification to comprehension, XBRL-Agent evaluated the capabilities of large language models to reason over full XBRL reports han2024xbrlagent.
+Shifting focus from classification to comprehension, XBRL-Agent evaluated the capabilities of large language models to reason over full XBRL reports (Han et al., 2024).
 The authors introduced two task types, domain taxonomy understanding and numeric reasoning, and found that base LLMs often hallucinated or misinterpreted financial content.
@@
 Collectively, these works broaden the methodological landscape of XBRL tagging from graph-based label propagation and embedding-based normalization to LLM-driven report analysis and point to a hybrid future where structured priors and reasoning tools jointly support accurate, scalable financial information extraction.
-B.3 Financial Evaluation Benchmarks
+B.3. Financial Evaluation Benchmarks
 In parallel to XBRL-specific advances, the financial NLP community has developed comprehensive benchmarks to assess broader capabilities in information extraction, numerical reasoning, and document understanding.
-FiNER-ORD shah2023finer introduced a high-quality, domain-specific NER dataset annotated over financial news, emphasizing general entity types like persons, organizations, and locations.
+FiNER-ORD (Shah et al., 2023) introduced a high-quality, domain-specific NER dataset annotated over financial news, emphasizing general entity types like persons, organizations, and locations.
 While not numerically focused like FiNER-139, it highlights the lexical diversity of financial discourse and establishes a strong baseline for testing pretrained and zero-shot LLMs in real-world financial NER scenarios.
-FinQA chen2021finqa pushed toward explainable QA by pairing expert-written questions with annotated multi-step reasoning programs derived from earnings reports.
-ConvFinQA chen2022convfinqa extended this challenge to conversational contexts, simulating real-world question flows over sequential financial queries.
-TAT-QA zhu2021tat focused on hybrid tabular-text reasoning and required models to align cell values and document narratives, often involving aggregation, comparison, and unit-scale interpretation.
-Pixiu xie2023pixiu introduced a broader evaluation framework by releasing FinMA, a financial LLM instruction-tuned across five tasks, and assessing it on a new benchmark covering sentiment classification, QA, summarization, NER, and stock prediction.
-BizBench koncel2023bizbench framed financial QA as program synthesis over realistic, multi-modal contexts, integrating reasoning, code generation, and domain knowledge into a single evaluation pyramid.
+FinQA (Chen et al., 2021) pushed toward explainable QA by pairing expert-written questions with annotated multi-step reasoning programs derived from earnings reports.
+ConvFinQA (Chen et al., 2022) extended this challenge to conversational contexts, simulating real-world question flows over sequential financial queries.
+TAT-QA (Zhu et al., 2021) focused on hybrid tabular-text reasoning and required models to align cell values and document narratives, often involving aggregation, comparison, and unit-scale interpretation.
+Pixiu (Xie et al., 2023) introduced a broader evaluation framework by releasing FinMA, a financial LLM instruction-tuned across five tasks, and assessing it on a new benchmark covering sentiment classification, QA, summarization, NER, and stock prediction.
+BizBench (Koncel-Kedziorski et al., 2023) framed financial QA as program synthesis over realistic, multi-modal contexts, integrating reasoning, code generation, and domain knowledge into a single evaluation pyramid.
 While these benchmarks highlight the growing ability of models to integrate structured and unstructured financial data, they overlook taxonomy-driven fact alignment and do not support the structured output formats required for XBRL tagging.
 Appendix C Data and Ticker Information
-C.1 The Statistics of the Tickers
+C.1. The Statistics of the Tickers
 Figure 5, Figure 6, and Figure 7 illustrate the distribution of 142 tickers across industry sectors, market capitalization categories, and geographic regions, highlighting the diversity of our collected financial reports.
 Considering that practical XBRL tagging practices often vary by industry, company size, and legal jurisdiction, we curated a diverse set of companies covering all 11 major industry sectors.
-The sample maintains a balanced distribution across firm sizes, following market capitalization categories: micro-cap (<\$300M), small-cap (\$300M–\$2B), mid-cap (\$2B–\$10B), large-cap (\$10B–\$200B), and mega-cap (>\$200B).
+The sample maintains a balanced distribution across firm sizes, following market capitalization categories: micro-cap (¡\$300M), small-cap (\$300M–\$2B), mid-cap (\$2B–\$10B), large-cap (\$10B–\$200B), and mega-cap (¿\$200B).
 In addition, we incorporated firms from over 30 states and international jurisdictions to capture regional differences in reporting and tagging conventions.
 This diversity ensures that the benchmark reflects realistic heterogeneity observed in financial disclosures across sectors, scales, and regulatory environments.
-Figure 5: Distribution of Tickers by Industry Sector.
-Figure 6: Distribution of Tickers by Market Cap Size.
-Figure 7: Distribution of Tickers by Geography.
-C.2 Data Granularity Justification
+Figure 5. Distribution of Tickers by Industry Sector.
+Figure 6. Distribution of Tickers by Market Cap Size.
+Figure 7. Distribution of Tickers by Geography.
+C.2. Data Granularity Justification
 Each 10-K report follows the SEC’s XBRL filing standard, where all narrative and tabular disclosures are linked through machine-readable tags defined in the US-GAAP taxonomy.
@@
 Overall, this segmentation strategy allows our benchmark to remain consistent with how XBRL tagging operates in practice (on localized factual statements), while maintaining scalability and interpretability across large collections of filings.
-C.3 Automated Data Annotation
+C.3. Automated Data Annotation
 Algorithm 3 converts Inline XBRL (iXBRL) filings into structured instances for our benchmark.
@@
 Appendix D Annotation Data Quality Control
-D.1 Validation Data Filtering
+D.1. Validation Data Filtering
 To establish a reliable upper bound on annotation quality, we construct a strategic audit subset rather than validating the full corpus.
@@
 Algorithm 4 Strategic Audit Set Sampling
-D.2 Annotator Demography
+D.2. Annotator Demography
 To ensure reliable verification of financial entities and taxonomy mappings, all audits were conducted by annotators with prior exposure to XBRL-based reporting and financial statement analysis.
@@
 Because the work involved publicly available financial disclosures and did not require annotators to process personal or sensitive data, the review posed minimal risk to participants.
-D.3 Validation Guideline
-D.3.1 Task Definition
+D.3. Validation Guideline
+D.3.1. Task Definition
 Given a context and an extracted triplet consisting of:
@@
   • 0 for incorrect.
-D.3.2 Validation Procedure
+D.3.2. Validation Procedure
 Follow the rules below, the definitions of entity type are shown in Table 9:
-  1. 1.
+  1. (1)
     Entity Check Determine whether the extracted entity is a numerical value.
@@
     • If no, set is_correct = 0.
-  2. 2.
+  2. (2)
     Entity Type Validation Verify whether the identified entity type is correct based on the context and the definitions in Table 9.
@@
     • If no, set is_correct = 0.
-  3. 3.
+  3. (3)
     US-GAAP Concept Validation Assess whether the assigned US-GAAP concept is appropriate based on the taxonomy tag definitions.
@@
     • If no, set is_correct = 0.
-Table 10: Entity type definitions used for validation.
+Table 9. Entity type definitions used for validation.
 Entity Type Definition
@@
 percentItemType Ratios or percentages, such as tax rates, growth rates, or discount rates, usually expressed with a percentage symbol (%).
-D.3.3 Important Instructions
-  1. 1.
+D.3.3. Important Instructions
+  1. (1)
     Non-Arabic Formats Financial numerical entities may appear in word form (e.g., ten million) and must be correctly identified and converted into standard numerical format.
-  2. 2.
+  2. (2)
     Magnitude Terms If a number is followed by a magnitude term (e.g., hundred, million, billion), do not expand it into the full numerical value:
@@
     • 10.6 million $\rightarrow$ extract only 10.6, not 10,600,000.
-  3. 3.
+  3. (3)
     Standardization of Format Remove formatting symbols while preserving the numerical value:
@@
     • Remove commas (e.g., 1,000 $\rightarrow$ 1000).
-D.4 Agreement Computation
+D.4. Agreement Computation
 We report both raw agreement and Cohen’s $\kappa$ to quantify inter-annotator reliability.
 Let $N$ denote the total number of instances, and let $n_{yy}$ be the number of cases on which both annotators assign the same label $y$. The raw agreement is
-$P_{o}=\frac{1}{N}\sum_{y}n_{yy}.$ (5)
+(4) $P_{o}=\frac{1}{N}\sum_{y}n_{yy}.$
 To adjust for agreement occurring by chance, we compute Cohen’s $\kappa$.
@@
 The expected chance agreement is
-$P_{e}=\sum_{y}p_{1y}\,p_{2y},$ (6)
+(5) $P_{e}=\sum_{y}p_{1y}\,p_{2y},$
 and Cohen’s $\kappa$ is
-$\kappa=\frac{P_{o}-P_{e}}{1-P_{e}}.$ (7)
-Following landis1977measurement, our $\kappa=0.81$ indicates substantial agreement.
+(6) $\kappa=\frac{P_{o}-P_{e}}{1-P_{e}}.$
+Following Landis and Koch (1977), our $\kappa=0.81$ indicates substantial agreement.
 Appendix E Evaluation Metrics
-To provide a fair evaluation of overall benchmark performance, we adopt a set of metrics, focusing primarily on macro-level and micro-level evaluation strategies inspired by the previous work sharma2023financial.
+To provide a fair evaluation of overall benchmark performance, we adopt a set of metrics, focusing primarily on macro-level and micro-level evaluation strategies inspired by the previous work (Sharma et al., 2023).
 Macro-level evaluation computes precision, recall, and F1 scores independently for each BIO-concept label derived from the US-GAAP taxonomy, and then averages them without weighting.
@@
 We evaluate the performance based on the following metrics:
-$\text{Precision}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{P}|}$ (8)
-$\text{Recall}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{G}|}$ (9)
-$\text{F1}=\frac{2\times\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (10)
+(7) $\text{Precision}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{P}|}$
+(8) $\text{Recall}=\frac{|\mathcal{G}\cap\mathcal{P}|}{|\mathcal{G}|}$
+(9) $\text{F1}=\frac{2\times\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$
 where $(e_{i},l_{i})=(e^{\prime}_{j},l^{\prime}_{j})$ if and only if both the entity span $e$ and its assigned type $l$ exactly match.
 For the FinCL subtask, Given a set of queries $\mathcal{Q}=\{q_{1},q_{2},\dots,q_{N}\}$, where each $q_{i}$ is associated with a ground-truth concept $c_{i}^{*}$ and a predicted concept $\hat{c}_{i}$, the accuracy is defined as:
-$\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\delta(\hat{c}_{i},c_{i}^{*})$ (11)
+(10) $\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\delta(\hat{c}_{i},c_{i}^{*})$
 where $\delta(\hat{c}_{i},c_{i}^{*})=1$ if $\hat{c}_{i}=c_{i}^{*}$, and 0 otherwise.
@@
 Table 10 provides an overview of the models evaluated in this study, categorized by openness, domain specialization, and architectural foundation. The evaluation covers a diverse range of models:
-  • Closed-source LLMs: We include GPT-4o hurst2024gpt, accessed via OpenAI’s API, as a representative of cutting-edge proprietary models with demonstrated performance across a variety of NLP tasks.
+  • Closed-source LLMs: We include GPT-4o (Hurst et al., 2024), accessed via OpenAI’s API, as a representative of cutting-edge proprietary models with demonstrated performance across a variety of NLP tasks.
 Although model size details are undisclosed, GPT-4o serves as an upper-bound reference in our benchmark.
-  • Open-source LLMs: This group encompasses recent, high-performing open models such as DeepSeek-V3 (685B) liu2024deepseek, DeepSeek-R1-Distill-Qwen (32B) deepseekai2025deepseekr1incentivizingreasoningcapability, and multiple variants of Qwen3  qwen3technicalreport(ranging from 0.6B to 32B).
-We also include Llama-3.3, Llama-3.2, and 3.1 variants  grattafiori2024llama (70B, 3B, and 8B), Llama-4-Scout-17B-16E-Instruct meta2025llama(109B), as well as Google’s gemma-2-27B  gemma_2024, to ensure architectural diversity and scalability comparison.
+  • Open-source LLMs: This group encompasses recent, high-performing open models such as DeepSeek-V3 (685B) (Liu et al., 2024), DeepSeek-R1-Distill-Qwen (32B) (DeepSeek-AI, 2025), and multiple variants of Qwen3  (Team, 2025)(ranging from 0.6B to 32B).
+We also include Llama-3.3, Llama-3.2, and 3.1 variants  (Grattafiori et al., 2024) (70B, 3B, and 8B), Llama-4-Scout-17B-16E-Instruct (Meta, 2025)(109B), as well as Google’s gemma-2-27B  (Team, 2024), to ensure architectural diversity and scalability comparison.
 These models are primarily instruction-tuned and optimized for general-purpose NLP tasks.
-  • Financial-specific LLMs: We evaluate Fino1-8B  qian2025fino1, a domain-specialized model trained on financial corpora, designed to better capture the terminology and structure unique to financial disclosures.
+  • Financial-specific LLMs: We evaluate Fino1-8B  (Qian et al., 2025), a domain-specialized model trained on financial corpora, designed to better capture the terminology and structure unique to financial disclosures.
 This category allows us to assess the benefits of domain adaptation in complex tagging and reasoning tasks.
-  • Pretrained Language Models (PLMs): To establish strong baselines, we include non-generative encoder models: BERT-large  DBLP:journals/corr/abs-1810-04805, FinBERT  araci2019finbert, and SECBERT  loukas2022finer.
+  • Pretrained Language Models (PLMs): To establish strong baselines, we include non-generative encoder models: BERT-large  (Devlin et al., 2018), FinBERT  (Araci, 2019), and SECBERT  (Loukas et al., 2022).
 These models have been widely used in prior financial NLP tasks and allow for a comparative analysis between generative LLMs and traditional pretrained models in terms of domain understanding and structured output capability.
 Together, these models offer a comprehensive evaluation spectrum, from general-purpose to domain-specific, encoder-based to decoder-based, and open to closed source, facilitating an in-depth assessment of their performance across our proposed benchmark tasks.
-Table 11: Model categories and corresponding repositories.
+Table 10. Model categories and corresponding repositories.
 Model Size Source
@@
 Appendix G The details for the fine-tuning PTMs
-G.1 Training data collection and processing
+G.1. Training data collection and processing
 Similar to the collection process for the 0.29412 0.47843 0.98431F0.36863 0.46667 0.96863i0.43922 0.45098 0.95294n0.5098 0.43529 0.93725T0.58431 0.42353 0.92157a0.6549 0.40784 0.90588g0.72549 0.39216 0.8902g0.79608 0.37647 0.87451i0.87059 0.36471 0.85882n0.94118 0.34902 0.84314g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark data, we gathered an additional 10 annual 10-K financial reports filed with the SEC for the period from February 13, 2024, to February 13, 2025, as summarized in Table 11.
@@
 The companies included in this dataset follow the XBRL standard, ensuring comprehensive coverage for training PTMs.
-Table 12: Financial report statistics summary for raw training data
+Table 11. Financial report statistics summary for raw training data
 Item Information
@@
 Overall, the annotated dataset includes 25,199 entities, covering 1,435 unique US-GAAP concepts.
-Table 13: Statistics of training data (tokens calculated with “cl100k_base” tokenizer, $\pm$ standard deviation).
+Table 12. Statistics of training data (tokens calculated with “cl100k_base” tokenizer, $\pm$ standard deviation).
 Structure Pos/Neg #Instance Avg. Tokens/S Avg. Entities/S Avg. Concepts/S Total Entities Unique Concepts
@@
 Specifically, B denotes the beginning of an entity phrase, I marks the continuation (inside) of an entity phrase, and O indicates tokens outside of any entity.
-As shown in Figure 8, “4.9” and “4.5” are single-token numerical entities labeled only with a B prefix (e.g., “B-us-gaap:AccountsReceivableNetNoncurrent”).
+As shown in Figure 8, ”4.9” and ”4.5” are single-token numerical entities labeled only with a B prefix (e.g., “B-us-gaap:AccountsReceivableNetNoncurrent”).
 To comprehensively cover all US-GAAP tags, we combine the entire set of 17,688 tags from the US-GAAP 2023 and 2024 taxonomy with the BIO labeling scheme to construct an extreme classification label space, resulting in 34,777 unique entity labels ($2\times 17388+1$).
-Figure 8: An example for training set annotation with BIO scheme.
+Figure 8. An example for training set annotation with BIO scheme.
 After constructing the training set, we reconstruct the testing set from the original benchmark dataset. The training settings are detailed below.
-G.2 Training settings
-We fine-tune three pretrained models, BERT-large DBLP:journals/corr/abs-1810-04805, FinBERT araci2019finbert, and SECBERT loukas2022finer, on our training set using the HuggingFace Transformers library.
+G.2. Training settings
+We fine-tune three pretrained models, BERT-large (Devlin et al., 2018), FinBERT (Araci, 2019), and SECBERT (Loukas et al., 2022), on our training set using the HuggingFace Transformers library.
 All models are trained with a batch size of 4, a learning rate of 3e-5, and for 20 epochs.
@@
 In contrast, the SAC strategy builds context based on the structural location of the entity: if the entity appears in a sentence, the entire sentence is used; if the entity appears in a table, we linearize the entire row into a Markdown-style key-value format to serve as the context.
-Table 14: Acc@k retrieval performance on the FinCL task. FWC: fixed window-based; SAC: structure-aware.
+Table 13. Acc@k retrieval performance on the FinCL task. FWC: fixed window-based; SAC: structure-aware.
 Strategy Structure Acc@1 Acc@10 Acc@20 Acc@50 Acc@100 Acc@150 Acc@200
@@
 For case 2, although both concepts denote non-current restricted cash, the model selected a narrower variant, showing the difficulty of distinguishing between highly similar concepts with subtle differences in scope.
-Table 15: Error cases from DeepSeek-V3 rerank.
+Table 14. Error cases from DeepSeek-V3 rerank.
 Case 1: \$31m
@@
 (5) Output the extracted financial entities in JSON list format without explanations, structured as follows: {"result":[{"Fact": <Extracted Numerical Entity>, "Type": <Identified Entity Type>}]} Input: {text/table} Output:
-Figure 9: Prompt template used for the FinNI subtask.
+Figure 9. Prompt template used for the FinNI subtask.
 Prompt Template for FinCL Subtask (Reranking) You are a financial tagging assistant trained in US-GAAP taxonomy.
@@
 Input Query: <entity> + <entity type> + <context> Candidate Tags: {Top 200 US-GAAP tags} Answer:
-Figure 10: Prompt template used for the Reranking stage in the FinCL subtask.
+Figure 10. Prompt template used for the Reranking stage in the FinCL subtask.
 Prompt Template for Ablation You are an XBRL tagging expert specializing in annotating financial numerical facts in XBRL reports.
@@
 (6) Output the extracted financial entities in JSON list format without explanations, structured as follows: {"result":[{"Fact": <Extracted Numerical Entity>, "Type": <Identified Entity Type>, "Tag": <Assigned US-GAAP tag>}]} Input: {text/table} Output:
-Figure 11: Prompt template used for ablation study.
+Figure 11. Prompt template used for ablation study.